### v2

# Assessment of Simulation-based Inference Methods for Stochastic Compartmental Models

Abstract
Global pandemics, such as the recent COVID-19 crisis, highlight the need for stochastic epidemic models that can capture the randomness inherent in the spread of disease. Such models must be accompanied by methods for estimating parameters in order to generate fast nowcasts and short-term forecasts that can inform public health decisions. This paper presents a comparison of two advanced Bayesian inference methods: 1) pseudo-marginal particle Markov chain Monte Carlo, short Particle Filters (PF), and 2) Conditional Normalizing Flows (CNF). We investigate their performance on two commonly used compartmental models: a classical Susceptible-Infected-Recovered (SIR) model and a two-variant Susceptible-Exposed-Infected-Recovered (SEIR) model, complemented by an observation model that maps latent trajectories to empirical data. Addressing the challenges of intractable likelihoods for parameter inference in stochastic settings, our analysis highlights how these likelihood-free methods provide accurate and robust inference capabilities. The results of our simulation study further underscore the effectiveness of these approaches in capturing the stochastic dynamics of epidemics, providing prediction capabilities for the control of epidemic outbreaks. Results on an Ethiopian cohort study demonstrate operational robustness under real‑world noise and irregular data sampling. To facilitate reuse and to enable building pipelines that ultimately contribute to better informed decision making in public health, we make code and synthetic datasets publicly available.
1 Introduction
The COVID-19 pandemic underscored the critical importance of epidemic modeling research. Rapid, data-driven approaches have demonstrated their capacity to provide public health decision-makers with essential insights for mitigating emerging infectious disease outbreaks and controlling epidemic dynamics [ethiopian_data, doms2018assessing]. Beyond COVID-19, other outbreaks such as Ebola and MPOX have likewise been designated as “public health emergencies of international concern” by the World Health Organization (WHO) [gpmb2024]. These crises and consequent public health challenges highlight the need for advanced modeling techniques capable of capturing the intricate dynamics of disease spread.
Epidemiological and medical research revealed that the spread of infectious diseases is shaped by a variety of processes, including transmission dynamics, patient responses to treatment, and pathogen mutation. A wide spectrum of mathematical modeling approaches is used to describe the underlying dynamics, spanning from standard ordinary differential equation (ODE) models [McKendrick_Kermack, hethcote2000] to agent-based models [bruch2015agent]. On the one hand, ODE-based population models provide a fast to simulate framework, but fall short to describe the variability of events and the resulting complex, stochastic interactions, due to their deterministic nature. Therefore, stochastic models are essential for realistically representing infectious disease dynamics observed in real-world settings, especially in small populations or during the early stages of an outbreak [britton2010stochastic]. On the other hand, purely stochastic models, such as agent-based and continuous-time Markov chain (CTMC) models [dragatz_book, infect_dis_modelling] provide highly detailed stochastic representations of individual-level processes, but come at the cost of prohibitive computational burden [petzold_2013]. Mathematical models based on stochastic differential equations – so called stochastic (meta-)population models – can solve the trade-off between fast model simulations and keeping the stochasticity in the model [dragatz_paper]. In this study, we consider compartmental models described by stochastic differential equations (SDE). In compartmental models individuals are arranged in a finite number of mutually disjoint compartments and interaction happens through transfer from one compartment to another which is described by transition equations [Brauer2008]. Their description in terms of SDEs can be seen as a mesoscopic view on the model dynamics. The SDEs are derived as a diffusion approximation to the microscopic, individual-based, discrete-state CTMC description [dragatz_paper, expansion_vanKampen, fearnhead2014inference] and in the large population limit resemble the macroscopic, population-based, (deterministic) ODE models [kurtz1970solutions]. For a more detailed discussion on stochastic epidemic modeling see [infect_dis_modelling, dragatz_paper, AnderssonAndBritton] and references contained therein.
A key challenge in mathematical-epidemiological modeling is determining the unknown model parameters that best explain the observed data. This process, often referred to as parameter inference, is commonly approached by maximizing the likelihood function, which measures the probability of the observed data given specific parameter values. For stochastic models, however, likelihood-based inference is often infeasible when fitting to discrete-time data, as the underlying probability distributions can be highly complex or the parameter space very high-dimensional leading to an intractable likelihood function. In such cases, naive approximations quickly become computationally prohibitive.
Many methods have been developed to address the issue of intractable likelihoods and posterior distributions. Early attempts used data augmentation or employed Markov chain Monte Carlo (MCMC) methods [dragatz_book, gibson1998estimating, roberts1999bayesian]. Later approaches sought to reduce computational cost through surrogate models such as Gaussian processes [swallow2022challenges] or linear noise approximations of the underlying Markov jump process [fintzi2022linear]. Hybrid strategies combining data augmentation with MCMC have also been proposed [mckinley2014simulation, el2016bayesian]. Sequential Monte Carlo methods were introduced as a flexible approach for analyzing stochastic state space models and performing Bayesian filtering tasks [doucet_sequential, chopin2013]. These methods, also known as particle filters (PF), have been used to generate parameter proposals of an MCMC algorithm, leading to pseudo-marginal methods for Bayesian inference on model parameters [andrieu_roberts, andrieu_2010, chopin_2020], a popular method for exact Bayesian inference in stochastic models [chopin_2020, wiqvist_2021].
Instead of determining model parameters via sampling based on evaluations of the likelihood or an approximation to it, it is possible to conduct the inference procedure solely based on simulations. In the widely used technique of Approximate Bayesian Computing (ABC), simulated and observed data is compared based on summary statistics [rubin1984, minter_2019]. However, manually selected summary statistics can be uninformative and their selection poses a challenge on its own. Additionally, as these methods require massive numbers of simulations, more recent machine learning approaches have been adapted for simulation-based inference with neural networks [cranmer2020frontier], namely neural posterior estimation methods. These promising techniques try to circumvent summary statistics and directly approximate the distribution of the parameters from samples generated by the simulator. Therein, conditional normalizing flows (CNF) are an advanced method that uses invertible neural networks to generate samples from the posterior distribution of the parameters [radev2020bayesflow].
PF and CNFs show great potential for Bayesian inference on stochastic models and are increasingly recognized within the epidemiological modeling community for providing probabilistic estimates of disease characteristics to effectively monitor pandemic outbreaks [radev2021outbreakflow, storvik2023]. However, their adoption has been limited by the lack of a comprehensive comparison assessing their relative performance. This work offers a praxis-driven comparison between both inference methods, assessing strengths and weaknesses of the particular approaches, as well as their alignment. In the first part, we described the two stochastic compartmental models, the well-known SIR model [McKendrick_Kermack] and a two-variant SEIR model, which was used to analyze the spread of COVID-19 in Ethiopia [ethiopian_data]. Together with the observational model bridging compartment trajectories with empirical data they serve as the foundation for our study. Moreover, we outline Bayesian inference for stochastic differential equation-based models, and explain the chosen techniques of CNF and PF together with the implementation of our comparison workflow. In the second part of the work, we outline the results from comparing the two inference methods on synthetically generated data for the two chosen compartmental models and additionally using real-data for the two-variant SEIR model. Both methods provide robust and reliable inference results on the stochastic version of the SIR model with synthetically generated data, validating their implementation at hand. Analyzing the agreement of the estimated marginal posteriors and model fits based on different synthetic datasets for the two-variant SEIR model, including (partially) missing data, shows that both methods yield good fits to the data at hand. However, ill-conditioning of the model leads to differences in the shape of the marginal posteriors due to difference in parameter space exploration by the inference methods. Applying both methods to an reparametrized version with parameter dimension reduction can overcome these issues and improves posterior alignment. Inference based on the data from [ethiopian_data] shows the applicability of both methods to real-world data. A discussion of the results, their limitations and further research directions follows in Section 6.
2 Stochastic Compartmental Models
Compartmental models are a mathematical framework to describe how individuals in a population move between different states and are fundamental for modeling infectious diseases [Brauer2008]. Such a model is characterized by the vector of compartment sizes $X\in\mathbb{R}^{d_{X}}$ and the transitions of individuals between compartments. A transition from compartment $i$ to compartment $j$ can be described in terms of the stochiometric coefficient $s_{i,j}\in\mathbb{R}^{d_{x}}$, the vector of changes to the compartment sizes, and the rate at which the transition happens
$a_{i,j}^{N}(X)=cX_{i}\sum_{k\in I_{i,j}}\frac{X_{k}}{N},$
where $c$ denotes a positive parameter, $I_{i,j}$ is an index set containing all indices $k\neq i$ of compartments involved in that transition, and $N$ gives the total number of individuals in the population. Let $\mathcal{T}$ denote the set of all tuples $(i,j)$, such that a transition from $X_{i}$ to $X_{j}$ exists. Then the time evolution of the compartment sizes can be modeled using a multivariate stochastic differential equation [dragatz_paper]:
$dX(t)$ $=\mu(X(t))dt+\sigma(X(t))d\mathbf{B}(t)$
$=\sum_{(i,j)\in\mathcal{T}}a_{i,j}(X(t))s_{i,j}dt+\sum_{(i,j)\in\mathcal{T}}\sqrt{a_{i,j}(X(t)))}s_{i,j}s_{i,j}^{T}d\mathbf{B}(t)$
where $\mathbf{B}$ denotes a vector of independent Brownian motions. The representation of compartmental models in terms of multivariate SDEs can capture the randomness of disease spread in their diffusion part $\sigma(X(t))$ and retain a drift component $\mu(X(t))$ for representing the deterministic trend of the population [AnderssonAndBritton, dragatz_book]. This makes them a favorable choice for modeling disease outbreaks and dynamics.
In this work, we consider two examples of SDE-based compartmental models, a simple standard Susceptible-Infected-Recovered (SIR) model [McKendrick_Kermack] and an extended multi-variant Susceptible-Exposed-Infected-Recovered (SEIR) model [hethcote2000] considering two variants of a single pathogen.
Figure 1: Schematic representations of the models and methods. (a) Graph of compartments and possible transitions with corresponding rate parameters for the SIR model. (b) Graph of compartments and possible transitions with corresponding rate parameters for the two‑variant SEIR model. (c) Workflow for the assessment of the Bayesian inference.
2.1 SIR Model
The standard SIR model (Figure 1(a)) considers three compartments: $S$ (Susceptible), $I$ (Infectious) and $R$ (Recovered). The model describes the number of individuals in these compartments given the transition parameters $\beta$ (the transmission rate) and $\gamma$ (the recovery rate). Due to its simplicity, the solution of the SIR model and its properties can be analyzed analytically [harko2014exact] and it often serves as an initial model for the modeling of unknown diseases [daley1999epidemic]. For consistency across models, in the following we compare recovery in terms of the mean infectious period, $\gamma^{-1}$, rather than $\gamma$. The events dynamics are described by three compartments and the transitions between them (Table 1).
Table 1: Description of the SIR model. Table including all events, such as initialization and transition of individuals, with their respective rate and model parameters.
Event Description Rate Parameter
Initialization Set initial number of infected individuals - $\mathrm{I}_{0}$
Infection $\mathrm{S}_{\text{}}\rightarrow\mathrm{I}_{\text{}}$ $\beta\frac{\mathrm{S}(t)\mathrm{I}(t)}{\mathrm{N}(t)}$ $\beta$
Recovery $\mathrm{I}_{\text{}}\rightarrow\mathrm{R}_{\text{}}$ $\gamma\mathrm{I}(t)$ $\gamma^{-1}$
2.2 Two-Variant SEIR Model
Many extensions of the SIR model have been developed to account for the heterogeneity of various infectious diseases [keeling2008], for example the Susceptible-Exposed-Infectious-Recovered model, in which the Exposed compartment covers the latent period between being infected and becoming infectious. Recently, an extension to the ODE-based SEIR model was introduced to capture the early COVID-19 dynamics in Ethiopia [ethiopian_data]. In addition to a classical SEIR model, this model distinguishes between two virus variants, namely the wild-type and a novel variant, leading to a total of 10 compartments (Figure 1(b)).
The novel variant is more potent in the sense that it shows a prolonged infectious period and that immunity from the wild-type variant does not protect against reinfection with the novel variant. However, an infection with the novel variant type confers immunity against subsequent infections with both the wild-type and the novel variant. The dynamics of the compartment sizes are described by initialization, variant entering and transitions between the compartments (Table 2).
The model exhibits non-identifiability issues in the sense that different parameter combinations, lying along certain curves, can generate nearly indistinguishable epidemic trajectories (see Supplementary Information A.3). Therefore, we also consider a reparametrized version of the model with a reduced number of model parameters to be estimated. We fix $\kappa^{-1}$ and use the following redefined parameters
$r_{0}=\beta\cdot\gamma^{-1},\quad e_{0}=\kappa^{-1}+\gamma^{-1},\quad s_{0}=s\cdot\gamma^{-1},$
while keeping $t_{\mathrm{var}}$ and $I_{0}$ as well as the observational and noise models unchanged. This transformation allows a comparisons of the inference methods CNF and PF in either parameter space without altering the underlying system.
Table 2: Description of the two-variant SEIR model. Table including all events, such as initialization and transition of individuals, with their respective rate and model parameters.
Event Description Rate Parameter
Initialization Set initial number of infected individuals - $\mathrm{I}_{0}$
Variant entry 500 individuals infected with variant enter the population - $t_{\mathrm{var}}$
Wild-type infection $\mathrm{S}_{\text{}}\rightarrow\mathrm{E}_{\text{wt}}$ $\beta\frac{\mathrm{S}(t)\mathrm{I_{wt}}(t)}{\mathrm{N}(t)}$ $\beta$
Wild-type incubation $\mathrm{E}_{\text{wt}}\rightarrow\mathrm{I}_{\text{wt}}$ $\kappa\mathrm{E_{wt}}(t)$ $\kappa^{-1}$
Wild-type recovery $\mathrm{I}_{\text{wt}}\rightarrow\mathrm{R}_{\text{wt}}$ $\gamma\mathrm{R_{wt}}(t)$ $\gamma^{-1}$
Variant infection $\mathrm{S}_{\text{}}\rightarrow\mathrm{E}_{\text{var}}$ $\beta\frac{\mathrm{S}(t)(\mathrm{I_{var}}(t)+\mathrm{I}_{\text{wt·var}}(t))}{\mathrm{N}(t)}$ $\beta$
Variant incubation $\mathrm{E}_{\text{var}}\rightarrow\mathrm{I}_{\text{var}}$ $\kappa\mathrm{E_{var}}(t)$ $\kappa^{-1}$
Variant recovery $\mathrm{I}_{\text{var}}\rightarrow\mathrm{R}_{\text{var}}$ $\frac{\gamma}{1.35}\mathrm{I}_{\text{var}}(t)$ $\gamma^{-1}$
Second infection $\mathrm{R}_{\text{wt}}\rightarrow\mathrm{E}_{\text{wt·var}}$ $\beta\frac{\mathrm{R_{wt}}(t)(\mathrm{I_{var}}(t)+\mathrm{I}_{\text{wt·var}}(t))}{\mathrm{N}(t)}$ $\beta$
Second incubation $\mathrm{E}_{\text{wt·var}}\rightarrow\mathrm{I}_{\text{wt·var}}$ $\kappa\mathrm{E}_{\text{wt·var}}(t)$ $\kappa^{-1}$
Second recovery $\mathrm{I}_{\text{wt·var}}\rightarrow\mathrm{R}_{\text{wt·var}}$ $\frac{\gamma}{1.35}\mathrm{I}_{\text{wt·var}}(t)$ $\gamma^{-1}$
2.3 Observation Model
The course of an epidemic is usually only partially observed and data is often only available in aggregated form on population level. For both models introduced above, we consider an observation model reflecting two often measured quantities to monitor infectious diseases. For the SIR model, the observables are
$y(t)=\left(\begin{array}[]{c}\frac{\mathrm{I}_{\text{}}(t)}{\mathrm{N}_{\text{}}(t)}\\[6.99997pt] \frac{\mathrm{N}_{\text{}}(t)-\mathrm{S}_{\text{}}(t)}{\mathrm{N}_{\text{}}(t)}\end{array}\right).$ (3)
For the two-variant SEIR model, they are given by
$y(t)=\left(\begin{array}[]{c}\frac{s(\mathrm{I}_{\text{wt}}(t)+\mathrm{I}_{\text{wt·var}}(t)+\mathrm{I}_{\text{var}}(t))}{\mathrm{N}_{\text{}}(t)}\\[6.99997pt] \frac{\mathrm{R}_{\text{wt}}(t)+\mathrm{E}_{\text{wt·var}}(t)+\mathrm{I}_{\text{wt·var}}(t)+\mathrm{R}_{\text{wt·var}}(t)+\mathrm{R}_{\text{var}}(t)}{\mathrm{N}_{\text{}}(t)}\\ \end{array}\right),$ (6)
where the first entry is the ratio of currently infected people describing the count data from testing, and the second entry describes seroprevalence, the ratio of individuals with antibodies present due to a prior infection. Because the observation process may have several limitations, an additional scaling parameter $s$ is used to capture that only a fraction of cases is reported [ethiopian_data].
To reflect measurement uncertainty, synthetic realizations of the observables are perturbed with noise. In this study, we leverage additive Gaussian noise. That means we have
$\overline{y}_{1}(t_{i})$ $=y_{1}(t_{i})+\varepsilon_{i,1},\qquad\varepsilon_{i,1}\sim\mathcal{N}(0,\sigma_{i,1}^{2}),$
$\overline{y}_{2}(t_{i})$ $=y_{2}(t_{i})+\varepsilon_{i,2},\qquad\varepsilon_{i,2}\sim\mathcal{N}(0,\sigma_{i,2}^{2}),$
where we truncate $\overline{y}_{1}(t_{i})$ and $\overline{y}_{2}(t_{i})$ at a lower bound of 0 to avoid unrealistic negative observations. As in the dataset given by [ethiopian_data], the standard deviations $\sigma_{i,1},\sigma_{i,2}$ were computed directly using a normal approximation to the binomially distributed testing count data. Alternative probability distributions may be more appropriate in other contexts. Further details on the computations of measurement noise are available in Supplementary Information A.2.
From each of the two stochastic compartmental models we generated synthetical data emulating the time course of a disease outbreak. We generated dense datasets with observations every seven or ten days. Additionally, to create realistic, heterogeneous observation schedules, we randomly subsampled the dense time series into sparse datasets reducing the number of time points from 57 to fewer than 15 per observable. We also introduced partial missingness in the sparse datasets, where only one of the two observables was available at certain time points. Details of the data generation and different datasets used are outlined in Supplementary Information A.6.
3 Bayesian Inference for Stochastic Compartmental Models
To infer the unknown model parameters $\mathbf{\theta}$, which best explain the data $\mathcal{D}^{*}=\{\overline{y}_{t_{i}}\mid i\in\{1,...,n\}\}$}, we apply Bayesian inference techniques. For this, we denote the space of unobserved states of the diffusion process by $\mathcal{X}$.
Bayes’ rule allows to combine prior information $p(\theta)$ (e.g., from expert judgment, previous studies, biological constraints, or model restrictions) with the likelihood $p(\mathcal{D}^{*}\mid\theta)$, which measures how probable the observed data are given specific parameter values, yielding the posterior distribution
$p(\theta\mid\mathcal{D}^{*})=\frac{p(\mathcal{D}^{*}\mid\theta)p(\theta)}{p(\mathcal{D}^{*})}\,\propto\,p(\mathcal{D}^{*}\mid\theta)p(\theta).$ (7)
The posterior represents the updated belief about the parameters after taking the observed data into account via the likelihood, effectively integrating both previous knowledge and the new information provided by the data. While the full posterior distribution is typically of primary interest, we additionally consider the mode (called maximum a-posteriori) of the posterior distribution as the best point estimate [bernardo1994bayesian].
One challenge in applying Bayesian inference to stochastic compartmental models is dealing with an intractable posterior distribution, because evaluating the likelihood
$p(\mathcal{D}^{*}\mid\theta)=\int_{\mathcal{X}}p(X,\mathcal{D}^{*}\mid\theta)dX$ (8)
requires marginalizing over stochastic realizations of the unobserved states at measurement time points.
In most cases, evaluating the density of hidden states analytically is impossible [Doucet_SMC], which prohibits evaluating the likelihood. Even in the case of a tractable likelihood, normalization of the posterior is done via the marginal likelihood, which can remain intractable. High‑dimensional parameter spaces further make discretization or numerical integration prohibitively expensive [wilkinson2010, hasenauer2013thesis].
Approximation techniques that estimate the posterior distribution without directly computing the complicated likelihood function enable robust parameter estimation, even in settings where traditional methods may fail due to mathematical or computational complexity. In this section, we will introduce two approaches (particle filters [golightly_particle_2011] and conditional normalizing flows [cranmer2020frontier]) that circumvent the computation of the likelihood function. PF approximate the likelihood with an unbiased estimator obtained by sequential importance sampling, and conditional normalizing flows learn a series of invertible mappings to transform a Gaussian density into the desired density.
3.1 Particle Markov Chain Monte Carlo Methods
Particle Markov chain Monte Carlo (PMCMC) methods are based on the idea of using an unbiased estimator of the likelihood function, obtained by a particle filter, inside a Markov chain Monte Carlo (MCMC) method [golightly_particle_2011].
Particle filters are sequential Monte Carlo (SMC) techniques based on importance sampling with resampling used for handling non-linear and non-Gaussian time-series models. Let $X_{t_{0}:t_{M}}$ denote a discrete time-series of the unobserved states and $y_{t_{0}:t_{M}}$ a time-series of observations at the same time points. The main idea is to iterate over the following three steps. First, $N$ particles, i.e., realizations of the dynamic process, are propagated through the latent space over time. This means that, given the trajectory of the latent process for the $k$-th particle and the observables until time $t_{i-1}$, we sample the state of the latent process at time $t_{i}$, $X_{t_{i}}^{(k)}\sim q_{t_{i}}\left(\cdot\mid X_{t_{0}:t_{i-1}}^{(k)},y_{t_{0}:t_{i-1}},\theta\right)$, based on a chosen proposal distribution $q_{t_{j}}$. Second, at each time point, particles are weighted, based on their agreement with the observed values for that time point. And third, they are resampled using auxiliary variables $A_{t_{j}}^{1:N}$ to discard low-weight particles. The resampled particles are then again propagated forward to the next time point. Hence, for each time $t_{i}$, the particle filter constructs a system of $N$ weighted particles $\{X_{t_{i}}^{(k)},\,w_{t_{i}}^{(k)}\}_{k=1}^{N}$ that approximates the filtering distribution $p(X_{t_{i}}\,\mid\,y_{t_{0}:t_{i}},\theta)$. Mathematically, this is closely linked to the framework of Feynman-Kac models [moral2004feynman, andrieu_2010, chopin_2020].
For the task of estimating the likelihood in stochastic compartmental models, we use a bootstrap filter Algorithm 1. This is a PF algorithm, where particles at the next time point are generated solely based on the transition probabilities $p_{\theta}(X_{t_{j}}\mid X_{t_{j-1}})$ of the latent process and then those new particles are filtered based on their agreement with the data at hand $g_{\theta}(Y_{t_{j}}\mid X_{t_{j}})$, where the function $g_{\theta}$ is determined by the observation and noise model. The bootstrap filter is easy to implement and widely applicable due to its simplicity. It only requires evaluation of the observation model $g_{\theta}(y_{t_{j}}\mid X_{t_{j}})$. However, it is not necessary to compute the transition densities $p_{\theta}(X_{t_{j}}\mid X_{t_{j-1}})$; we solely need the ability to simulate the state of the process at the next time point given the current state, making it practical for any complex or stochastic model.
Algorithm 1 Bootstrap Filter with Adaptive Resampling
1:parameter $\theta$, number of particles NN, observation density function $g_{\theta}$, transition density function $p_{\theta}$, initial distribution $\pi$
[…]
3:Sample $X_{t_{0}}^{(n)}\sim\pi(x_{t_{0}})$
[…]
6:for $i=1$ to $M$ do
7:  if $\mathrm{ESS(W}_{t_{i-1}}^{1:N})<\mathrm{ESS_{min}}$ then
8:   Sample $A^{1:N}_{t_{i}}\sim\text{resample}(W_{t_{i}}^{1:N})$    $\triangleright$ See Supplementary Information A.4.2
9:   $\hat{w}_{t_{i-1}}^{(n)}\leftarrow 1$
10:  else
11:   $A_{t_{i}}^{(n)}\leftarrow n$            $\triangleright$ No resampling
12:   $\hat{w}_{t_{i-1}}^{(n)}\leftarrow w_{t_{i-1}}^{(n)}$
13:  end if
14:  Sample $X_{t_{i}}^{(n)}\sim p_{\theta}(x_{t_{i}}\mid X_{t_{i-1}}^{A_{i_{j}}^{(n)}})$
15:  $w_{t_{i}}^{(n)}\leftarrow\hat{w}_{t_{i-1}}^{(n)}g_{\theta}(X_{t_{i}}^{(n)}\mid X_{t_{i-1}}^{(A_{t_{i}}^{(n)})})$
16:  $W_{t_{i}}^{(n)}\leftarrow\frac{w_{t_{i}}^{(n)}}{\sum_{k=1}^{N}w_{t_{i}}^{(k)}}$
17:end for
Although the name filtering refers to the task of estimating the latent state $X_{t_{i}}$ online given data $y_{t_{0}:t_{i}}$, a PF yields, as a by-product, a scheme to compute an unbiased estimator of the likelihood contributions ${\{p(y_{t_{0}:t_{i}}\,\mid\,\theta)\}}_{i\geq 0}$, crucial for the task of parameter inference [chopin_2020]. This estimator is obtained by computing the normalizing constants
$l_{t_{i}}^{N}(\theta)=\frac{\sum_{k=1}^{N}w_{t_{i}}^{(k)}}{\sum_{k=1}^{N}w_{t_{i-1}}^{(k)}}\quad\text{and}\quad L_{t_{i}}^{N}(\theta)=\prod_{j=0}^{i}l_{t_{j}}^{N},$ (9)
where $L_{t_{i}}^{N}(\theta)$ corresponds to the desired estimate of $p(y_{t_{0}:t_{i}}\mid\theta)$. The estimate $L_{t_{M}}^{N}(\theta)$ can then be used inside a pseudo-marginal MCMC algorithm to sample from the posterior $p(\theta\mid y_{t_{0}:t_{M}})$.
In this work, we use the likelihood estimate obtained by the bootstrap filter within the acceptance probability of a Metropolis-Hasting (MH) algorithm as the outer MCMC scheme, making it a member of the class of Pseudo-Marginal Metropolis-Hasting methods (PMMH) [andrieu_roberts, andrieu_2010, chopin_2020].
Using an unbiased estimator $\hat{p}(y|z,\theta)$ of the $p(y|\theta)$ that is obtained by the use of auxiliary random variables $z\sim r(\cdot)$ within the acceptance of a MH algorithm then corresponds to simulating a Markov chain $\{\theta^{i},z^{i}\}_{i\geq 0}$ targeting the joint density
$\bar{p}(\theta,z)\propto p(\theta)r(z)\hat{p}(y\mid\theta,z).$
For the bootstrap filter, we can, for instance, think of the $A_{t_{j}}^{1:N}$ used in the resampling as being the auxiliary random variables. The joint density admits the correct posterior density of the model $p(\theta\mid y)$ as a marginal and therefore the algorithm samples the parameter vectors $\theta$ from the true posterior [andrieu_roberts, andrieu_2010, doucet_2015]. With respect to an extended distribution, PMMH algorithms are standard MCMC samplers.
One can use any unbiased estimator within any MCMC algorithm to produce a PMCMC algorithm. Different filtering and sampling algorithms impose different practical and implementational challenges and we focus here on the most general applicable version [doucet_2015, chopin_2020]. In the subsequent sections PF method always refers to a PMMH algorithm consisting of a bootstrap filter inside an adaptive Metropolis-Hastings algorithm.
3.2 Neural Posterior Estimation And Conditional Normalizing Flows
Neural posterior estimation (NPE) is a simulation-based inference (SBI) method which frames posterior construction as a (conditional) density-estimation problem. Given a simulator that produces samples $(\mathcal{D},\theta)$ by first drawing $\theta\sim p(\theta)$ and then $\mathcal{D}\sim p(\mathcal{D}\mid\theta)$, NPE fits a conditional neural network $p_{\phi}(\theta\mid\mathcal{D})$ to approximate the true posterior $p(\theta\mid\mathcal{D})$. Typically the data $\mathcal{D}$ are preprocessed by a (learned) summary statistic [radev2020bayesflow], which serves as the input to the conditional density estimator. For clarity, this preprocessing step is omitted from the notation below.
In this work, we focus on Conditional Normalizing Flows (CNFs) [rezende2015variational]. In general, normalizing flows transform a simple base distribution (e.g., a standard Gaussian) into a complex target distribution via a sequence of invertible, differentiable mappings. In our setting, this means that a normalizing flow defines a diffeomorphism $f_{\mathcal{D}^{*}}$ such that (for fixed $\mathcal{D}^{*}$) the posterior $p(\theta\mid\mathcal{D}^{*})$ is the “pushforward” of the base distribution $q$ along $f_{\mathcal{D}^{*}}$, i.e.,
$\theta\sim p(\theta\mid\mathcal{D}^{*})\Leftrightarrow\theta=f_{\mathcal{D}^{*}}^{-1}(z),\quad z\sim q(z),$
This construction enables efficient sampling and exact density evaluation via the change‑of‑variables formula, making normalizing flows a powerful tool for modeling complex probability distributions.
CNFs extend this idea by allowing the transformation to additionally depend on datasets $\mathcal{D}$, generated by first drawing $\theta\sim p(\theta)$ and then $\mathcal{D}\sim p(\mathcal{D}\mid\theta)$. Rather than learning a single transformation $f$, CNFs learn a family of conditional flows $\{f_{\mathcal{D}}\}$, with each $f_{\mathcal{D}}$ mapping the base distribution $q$ to the posterior $p(\theta\mid\mathcal{D})$ [winkler2019conditional]. This conditioning enables CNFs to capture how the posterior changes with different observations, making them particularly well-suited for amortized inference tasks, where a model learns to perform inference efficiently across multiple datasets rather than solving each case from scratch [kobyzev2020normalizing, rezende2015variational].
Instead of hand-designing each $f_{\mathcal{D}}$, we parameterize the family using a neural network $f_{\phi}(\cdot,\,\mathcal{D})$ whose parameters $\phi$ are shared across all $\mathcal{D}$, but whose behavior is modulated by $\mathcal{D}$ itself [ardizzone2019guided]. We train the neural network by minimizing the Kullback-Leibler divergence between the true and learned posteriors for all $\mathcal{D}$ in the training distribution [weilbach2020structured, siahkoohi2020conditional, radev2023bayesflow]. This means, our optimization objective is finding $\hat{\phi}$
$\displaystyle\hat{\phi}=\arg\min_{\phi}\;-\mathbb{E}_{(\theta,\mathcal{D})\sim p(\theta)p(\mathcal{D}\mid\theta)}\big[\log p_{\phi}(\theta\mid\mathcal{D})\big]$
For CNFs this requires computing the change‑of‑variables log‑density
$\log p_{\phi}(\theta\mid\mathcal{D})=\log q(f_{\phi}(\theta,\mathcal{D}))+\log\Big|\det\Big(\frac{\partial f_{\phi}(\theta,\mathcal{D})}{\partial\theta}\Big)\Big|$
for simulator samples and optimizing $\phi$ by stochastic gradient descent. This amortizes the cost of inference: Once trained, the CNF can quickly return an approximate posterior for any new $\mathcal{D}$ within the support of the training distribution. Algorithm 2 summarizes the amortized training loop and the inference procedure for drawing approximate posterior samples via inverse flow sampling.
Algorithm 2 Amortized training and inference for conditional normalizing flows
1:Data source $S$ (pre‑simulated dataset or simulator callable $\sim p(\theta)p(\mathcal{D}\mid\theta)$), base density $q(z)$, conditional flow $f_{\phi}(\cdot;\mathcal{D})$, optimizer Opt, learning rate $\eta$, batch size NN, iterations $T$, inference data $\mathcal{D}^{*}$, inference sample count $M$
2:Trained parameters $\phi$; sampler for approximate posterior $p_{\phi}(\theta\mid\mathcal{D}_{*})$
3:for $t=1$ to $T$ do
4:  $\{(\theta_{i},\mathcal{D}_{i})\}_{i=1}^{N}\leftarrow\text{sample from $S$ (dataset or simulator)}$
5:  for $i=1$ to NN do
6:   $z_{i}\leftarrow f_{\phi}(\theta_{i};\mathcal{D}_{i})$
7:   $\log p_{\phi}(\theta_{i}\mid\mathcal{D}_{i})\leftarrow\log q(z_{i})+\log\Big|\det\Big(\frac{\partial f_{\phi}(\theta_{i};\mathcal{D}_{i})}{\partial\theta}\Big)\Big|$
8:  end for
9:  $\mathcal{L}\leftarrow-\frac{1}{N}\sum_{i=1}^{N}\log p_{\phi}(\theta_{i}\mid\mathcal{D}_{i})$
10:  $\phi\leftarrow\text{Opt}(\phi,\nabla_{\phi}\mathcal{L},\eta)$
11:end for
12:
[…]
14:  for $m=1$ to $M$ do
15:   Sample $z_{m}\sim q(z)$
16:   $\theta_{m}\leftarrow f_{\phi}^{-1}(z_{m};\mathcal{D}_{*})$
17:  end for
18:  return $\{\theta_{m}\}_{m=1}^{M}$
19:end procedure
This amortization not only enables fast inference once training is complete, but also facilitates simulation‑based calibration (SBC) diagnostics by allowing posterior approximations to be generated efficiently across many simulated datasets. At the same time, performance of CNFs is tied to the support of the training distribution: the model can interpolate to new values if they are close to the training support, but performance degrades when extrapolating far beyond it [winkler2019conditional, talts2018validating].
4 Implementation Of The Methods And Their Comparison
In order to render the simulation based approaches computationally feasible, model simulations were conducted using the efficiency of the Julia programming language and the flexible SDE modeling capabilities of SciML and ModelingToolkit.jl [julia_stochdiffeq, ma2021modelingtoolkit]. To approximate the solution of the SDE, we implemented an Euler-Maruyama scheme that constrains trajectories to stay non-negative, effectively clipping potential negative parts and setting them to zero.
For PF, existing software packages were unsuitable as they either lacked Julia support or restricted model specification to simulator-only interfaces. We therefore implemented a custom particle filter in Julia, drawing inspiration from the Python package Particles [chopin_2020]. As an outer sampling scheme, we employed an adaptive Metropolis–Hastings algorithm [adaptive_mh], using the implementation in pyPESTO [pypesto], to account for potential multimodality in the posterior distribution. Missing data were handled by setting the likelihood contribution of the corresponding observation to zero. To mitigate particle degeneracy, we set the number of particles to 200, ensuring that the variance of the likelihood estimate remained below one for parameters around the posterior mode (Supplementary Figure S3). For the outer MH algorithm, we ran 4 Markov chains with a length of 50,000 samples each. We then discarded a burn-in period of the first 25,000 samples per chain. After that, convergence was evaluated using rank‑normalized Gelman–Rubin statistics ($\hat{R}$) and the ESS with autocorrelations up to a maximal lag of 250, both from MCMCDiagnosticTools.jl [turing_jl]. Values of $\hat{R}$ close to 1.0 indicate good convergence and well‑mixed chains [vehtari2021rank]. The ESS assesses and diagnoses poor convergence in the bulk of the distribution due to trends or different locations of the chains. For a given estimand, it is recommended that ESS should exceed $100\cdot\text{nchains}$ [vehtari2021rank]. Together, $\hat{R}$ and ESS provide complementary diagnostics of chain convergence and posterior exploration quality.
For CNF, posterior approximations were obtained via neural posterior estimation with learned summary statistics [radev2020bayesflow], using the BayesFlow software (v1) for amortized Bayesian workflows [radev2023bayesflow]. We trained an independent neural network for each experiment rather than a single globally amortized model. To ensure comparability, all experiments used identical hyperparameters and simulation budgets, determined through a preliminary hyperparameter search. The dimension of the learned summary statistics was set to $2n_{\theta}+2$, where $n_{\theta}$ denotes the number of model parameters. Training was conducted offline with 100,000 simulations for the training dataset and 400 simulations for validation. Models were trained for 100 epochs with a batch size of 32. Summary statistics were extracted using a SequenceNetwork [radev2021outbreakflow] with 64 LSTM units, and the conditional normalizing flow was modeled using an InvertibleNetwork with spline coupling and eight coupling layers, both implemented within the BayesFlow framework. Detailed configurations are provided in Supplementary Information A.5. Handling non‑equidistant time points and missing data required a specialized strategy. We adopted the approach of [wang2024missing], which augments the input with binary indicators marking the presence or absence of each entry. This encoding allows the neural network to account for missing values during inference. For experiments with dense, equidistant data, this additional encoding was not applied. For sparse datasets, the resulting neural networks were trained on the specific missingness pattern present in the simulated data; no amortization across different missingness structures was performed. Empirical cumulative distribution function (ECDF) plots were obtained to assess calibration, and parameter recovery diagnostics were used to evaluate the correspondence between inferred and true parameter values. These diagnostics provide insight into whether the amortized posterior captures uncertainty appropriately and whether point estimates remain consistent with the generative parameters.
For the comparison of both methods we used visual and quantitative measures. The visual comparison using violin and density plots of the marginal posteriors allows comparison of the shape and variability of the posterior approximations. This is complemented by a visual comparison of the model fit. For the evaluations we draw 10,000 parameter vector samples from each approximate posterior distribution and, for each parameter vector, simulate an observable trajectory to create an ensemble. We then plot the central 50%, 90%, and 95% intervals of the ensemble against the data. These are independent posterior predictive samples, obtained by evaluating $p(y\mid\theta)$ for draws $\theta\sim p(\theta\mid\mathcal{D}^{*})$. For a quantitative measure of inference accuracy, we compute the maximum a-posteriori (MAP) estimate in the parameter space via Gaussian kernel density estimation using Silverman’s rule [silverman1998density] as well as the root-mean squared error (RMSE). Specifically, the RMSE was calculated for trajectories simulated from each of the 10,000 posterior samples and then averaged, providing an estimate of the expected RMSE under the posterior distribution. For synthetic experiments, the RMSE was also computed with respect to the true parameters used to generate the data. For the real-data experiment, true parameters are not available. Instead, we report RMSE relative to published parameter estimates from [ethiopian_data], which serve as a reference baseline.
All experiments were conducted on ten cores of an AMD EPYC 7f72 3.2 GHz processor with a total of 200 GB of RAM made available to the process. In order to ensure reusability and reproducibility, we made the code and all artificial data used for the experiments available at Zenodo (https://zenodo.org/records/17779579) and in a GitHub repository (https://github.com/vwiela/Inference-Methods-for-Stochastic-Compartmental-Models.git).
5 Results
We examined inference with CNF and PF for stochastic epidemic models through a sequence of in silico experiments on synthetic data, following the workflow in Figure 1(c). We begin with the standard SIR model (Section 2.1) using a synthetic prior to validate both methods. We then turn to the two-variant SEIR model (Section 2.2) with literature-informed priors [mcaloon2020incubation, fang2020comparisons], designed to probe the influence of parameter non-identifiability (Section A.3). Within the two-variant SEIR setting, we additionally study the effects of data sparsity and timing mismatch and investigate a reparametrization aimed at reducing non-identifiability. For completeness, Table 3 summarizes all settings and variations. Finally, we apply both approaches to longitudinal data from Ethiopia [ethiopian_data] to assess performance under realistic conditions (Section 5.4). Additional configurations and results are provided in the Supplement (Supplementary Tables S4S18).
Table 3: Experiments for method comparison. Overview over the subsections and the corresponding model and data setup used.
Section Description
[…]
Reparametrization of the two-variant SEIR model with synthetic dense datasets.
[…]
Two-variant SEIR model with dataset from a longitudinal cohort study
in Ethiopia [ethiopian_data].
5.1 CNF And PF Provide Accurate Estimates For The SIR Model.
We compared convergence and posterior estimates obtained with CNF and PF on the SIR model, using two distinct simulated dense datasets (details in Supplementary Information A.6). The prior distributions and parameter values are summarized in Supplementary Tables S3 and S4.
Both CNF and PF yielded highly consistent posterior estimates for the SIR model. The marginal and joint posterior distributions overlapped almost perfectly (Figure 2A), and posterior predictive checks showed that epidemic trajectories generated from either method fit the observed data equally well (Figure 2B). MAP estimates from CNF and PF were nearly identical and close to the true parameters, with RMSEs for infection counts and seroprevalence matching those obtained from simulations with the ground truth (Table 4).
Beyond these direct comparisons, the methods differed in sampling quality. PF achieved excellent convergence, with ESS exceeding 2000 and $\mathbf{\hat{R}}$ values close to 1, indicating robust posterior exploration and well-mixed chains (Supplementary Tables S6 and S7). CNF, in contrast, showed deviations in SBC calibration diagnostics (Supplementary Figure S23), reflecting limitations of the amortized posterior. Notably, these calibration issues were not apparent in the parameter recovery plots: recovered parameters aligned closely with the ground truth and error bars remained narrow (Supplementary Figure S27). This highlights that while CNF’s posterior uncertainty is imperfectly calibrated in a formal sense, the practical impact on point estimates and predictive accuracy is minimal.
Figure 2: Evaluation of the posterior approximations for the SIR model
A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade. B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
In summary, CNF and PF both provide robust Bayesian inference for small stochastic compartmental models. Across datasets, both methods produced equivalent posterior distributions and predictive performance. Notably, CNF, despite being an amortized approach, achieved greater computational efficiency in our experiments, roughly 10-times (Supplementary Table S27), while PF excelled in sampling diagnostics. Moreover, since CNF relies purely on simulations it has consistent computational costs across different parameter vectors. In contrast, PF execution times vary greatly, depending on the posterior geometry. Additionally, the evaluation of the likelihood can be computationally expensive in regions where outcomes are highly sensitive to parameter changes or when parameter vectors do not fit the data well.
Table 4: MAP and RMSE for SIR model.
$\mathbf{\beta}$ $\mathbf{\gamma^{-1}}$ $\mathbf{RMSE_{infc}}$ $\mathbf{RMSE_{prev}}$
True Par. 0.1000 20.00 0.0193 0.0230
CNF 0.1026 18.99 0.0168 0.0201
PF 0.1021 19.32 0.0150 0.0121
5.2 Non-Identifiability Issues Lead To Posterior Shape Differences
Previous studies have shown that parameters in common epidemiological models are often weakly determined by commonly available datasets [raimundez2021covid]. To assess how such non-identifiabilities affect inference reliability, we compared PF and CNF on the introduced two‑variant SEIR-model with pronounced non-identifiability, making it a suitable test case for evaluating how posterior estimates differ under CNF and PF.
We first examined posterior distributions and predictive performance using four dense simulated observational time series. Posterior distributions showed markedly lower variance compared to the corresponding priors, indicating strong information gain from the data (Figure 3). The notable exception is $\kappa^{-1}$ under CNF, whose marginal remained close to the prior. This behavior is typical in models where certain parameters are weakly informed by the data (see [wang2021posterior] for a more thorough discussion). The PF posterior approximation showed a higher contraction than that of CNF. Additionally, it was also largely enveloped in the CNF posterior approximation, across parameters (Figure 4A). Despite these width differences, epidemic trajectories generated from both methods fit the data equally well (Figure 4B), and MAP estimates together with RMSE scores were close to the true parameters (Table 5). PF estimates were marginally closer to the ground truth, highlighting the method’s focus on trajectory agreement through resampling.
Figure 3: Marginal posterior approximations and prior distribution. Histograms of posterior approximations and prior distribution for the two-variant SEIR model using a dense dataset.
As a secondary analysis, we assessed calibration and convergence diagnostics. CNF showed good calibration in ECDF plots (Supplementary Figure S20), though parameter recovery was imperfect for several parameters and markedly worse for $\kappa^{-1}$, which showed weak correlation with the ground truth (Supplementary Figure S26). PF chains displayed good convergence, with $\mathbf{\hat{R}}$ being close to one across parameters (Supplementary Table S14). However, ESS were relatively low ($\sim$500–600) for several parameters (Supplementary Table S13), indicating limited exploration of posterior tails.
Figure 4: Evaluation of the posterior approximations for the two-variant SEIR model with a dense dataset.
A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade. B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
Table 5: MAP and RMSE for full SEIR model.
$\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa^{-1}}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I_{0}}$ $\mathbf{RMSE_{infc}}$ $\mathbf{RMSE_{prev}}$
True Par. 17.00 5.00 0.0800 3.00 150.00 500.00 0.0200 0.0245
CNF 15.47 4.93 0.0867 3.29 160.59 446.30 0.0237 0.0237
PF 15.62 5.04 0.0855 3.22 153.02 479.16 0.0228 0.0291
In practice, observations are often non‑equidistant or partially missing. To reflect this, we conducted additional experiments with reduced observation density (Section 2.3). Results for sparse datasets were consistent with those obtained under dense observations. Posterior distributions from CNF and PF showed similar shapes, and epidemic trajectories generated from both methods fit the data equally well (Supplementary Tables S8 to S13), with MAP estimates and RMSE values comparable to the dense setup (Supplementary Table S16). Additionally, CNF remained well calibrated and PF continued to show well‑mixed chains, though ESS values remained low for several parameters, mirroring the diagnostics observed in the dense case (Supplemetary Tables S17 and S18). Together, these results indicate that both methods can accommodate sparse and irregular data without loss of inference quality, though CNF requires careful amortization design to handle heterogeneous observation schedules.
Taken together, these results show that CNF and PF both provide accurate inference for the two‑variant SEIR model, even under pronounced non‑identifiability. CNF yields broader posterior approximations that better capture low‑probability regions, but do not strongly support specific parameter values. PF, in contrast, produces narrower posteriors and more certain point estimates, despite limited exploration of posterior tails due to low ESS. Predictive performance was high for both methods across dense and sparse datasets, with consistent MAP and RMSE values. Sparse observations introduced implementation overhead for CNFs but did not substantially affect inference quality for either method. Additionally, CNF were approximately 10-times faster independent of the used dataset (Supplementary Table S27).
5.3 Reducing Parameter Dependencies Improves Posterior Alignment
Complementing the analysis of the full SEIR2v model, we evaluated the reparametrized version of the SEIR model to explore how reduced parameter dependencies affect inference results for PF and CNF.
Posterior distributions from CNF and PF showed close alignment across parameters, indicating similar exploration of the parameter space (Figure 5A). Both methods yielded comparable MAP estimates near the true values, with minor deviations for $e_{0}$ and $I_{0}$, where PF shifted toward lower values. Epidemic trajectories fit the data equally well (Figure 5B) and RMSE values were similar across CNF, PF, and the true parameters, with PF showing marginally better agreement on the infection count data (Table S21).
Additionally, we evaluated calibration and convergence diagnostics. CNF showed SBC miscalibration in ECDF plots (Figure S22), consistent with the results reported for the identifiable SIR model. PF, in contrast, achieved ESS exceeding 2000 (Supplementary Table S22), indicating well‑mixed chains and thorough posterior exploration.
Figure 5: Evaluation of the posterior approximations for the reparametrized two-variant SEIR model with dense data.
A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade. B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
Table 6: MAP and RMSE for reparametrized SEIR model and dense datasets.
$\mathbf{r0}$ $\mathbf{e0}$ $\mathbf{s0}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$ $\mathbf{RMSE_{infc}}$ $\mathbf{RMSE_{prev}}$
True Par. 1.36 22.00 51.00 150.00 500.00 0.0200 0.0245
CNF 1.34 21.33 50.27 145.46 489.71 0.0216 0.0236
PF 1.34 20.57 51.21 156.94 467.30 0.0319 0.0271
In summary, reducing parameter dependencies improved identifiability and brought CNF and PF posterior shapes into closer agreement. While CNF exhibited minor calibration issues, PF benefited from substantially improved ESS, supporting robust posterior exploration in this setting. However, the reparametrization did not enhance predictive performance and relies on fixing $\kappa^{-1}$, which should be approached with caution in applied inference.
5.4 CNF and PF yield consistent results on real-world datasets
To complement the synthetic experiments, we applied both methods to data from a longitudinal cohort study of the COVID‑19 outbreak in Ethiopia [ethiopian_data]. Our setup differed from the original study in that we used a stochastic rather than deterministic formulation, offering greater flexibility and better reflecting the inherent randomness in infection dynamics. Due to the unavailability of ground truth, we used the published mean of the parameter samples as a reference.
Posterior approximations from CNF and PF shared a similar overall shape, with the PF posterior largely nested within the broader CNF and exhibiting reduced variance, particularly for the rate parameters $\beta$, $\kappa^{-1}$, and $\gamma^{-1}$ (Figure 6A). Trajectory ensembles derived from the inferred posteriors fit the data well, even under the relatively high noise level reported in the original study; one seroprevalence point was not captured by PF within its uncertainty intervals (Figure 6B). Despite this discrepancy, the PF MAP achieved RMSE values that matched or improved upon those reported in the original study. Although the CNF MAP deviated more strongly from both the published parameters and PF, the RMSE values of the respective distributions (Table S24) were comparable and showed improved accuracy for infection count data.
Diagnostics were consistent with patterns observed in controlled experiments: CNF exhibited good calibration alongside imperfect parameter recovery (Figure S21, Figure S25), while PF demonstrated convergence despite low effective sample sizes (Table S25, Table S26).
Figure 6: Evaluation of the posterior approximations for the two-variant SEIR model with real data.
A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade. B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
Table 7: MAP and RMSE for full SEIR model for real dataset.
$\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa^{-1}}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I_{0}}$ $\mathbf{RMSE_{infc}}$ $\mathbf{RMSE_{prev}}$
Pub. Par. 16.70 5.00 0.0800 2.30 184.50 635.00 0.0737 0.0702
CNF 14.80 4.14 0.0882 3.16 183.04 550.16 0.1251 0.0928
PF 15.03 4.67 0.0858 3.36 160.35 752.08 0.0856 0.0606
Taken together, applying both methods to the Ethiopian cohort data yielded results consistent with the synthetic analyses, underscoring the robustness of each approach. The trade‑offs identified in controlled settings—CNF’s broader posterior coverage versus PF’s tighter estimates with lower ESS—carried over to this real‑world application, while predictive performance remained strong for both methods.
6 Discussion
Responses to public health crises, such as outbreaks of infectious diseases, creates a need for informed decision making by realistically modeling the dynamics using stochastic models. Yet, fitting such models to observed data present many difficulties and several methods were developed address this challenge. In this work, we systematically evaluated two state-of-the-art methods for Bayesian parameter inference, Conditional Normalizing Flows (CNF) and pseudo-marginal particle MCMC, referred to as Particle Filters (PF), across a range of stochastic compartmental models. Our analysis encompassed synthetic and real-world data, varied model structure, and different levels of data sparsity, yielding crucial insights into the strengths and limitations of each methodology when applied in epidemiological modeling scenarios. Across all investigated scenarios, both methods produced accurate fits and informative posteriors. However, this study revealed systematic differences of uncertainty in the parameter space due to different posterior geometries and model design choices.
In the context of an identifiable stochastic model, such as the SIR model, both CNF and PF yielded highly consistent posterior distributions and closely matched predictive epidemic trajectories. The marginal and joint posterior estimates and MAP estimates were near the data-generating parameters and the root mean square error (RMSE) values confirmed high predictive accuracy for both methods, despite imperfect calibration of the CNFs. Additionally, CNF exhibited very good recovery plots in this setting. As a exact method, PF showed a good recovery of the true parameters in this setting, that were closely matched by the CNF. Together, this indicates that, once trained, CNF can deliver inference quality on a par with an MCMC based baseline while enabling possible amortization across different datasets and yielding high computational efficiency in the inference phase.
Expanding the analysis, we applied both methods to a more complex stochastic two-variant SEIR model with practical non-identifiabilities and parameter couplings. In this setting, both methods fit the data well and achieved similar predictive power and uncertainty, yet PF produced narrower marginal posteriors that were largely enclosed by the broader CNF posteriors and simultaneously showed low effective sample sizes for several parameters. This pattern is consistent with resampling in PF concentrating particles in high-likelihood basins, i.e., on the fit to the data. This property can limit tail exploration in the parameter space especially when the posterior surface contains ridges or mountain-pass structures. CNF achieved good calibration results in this setting, although its recovery plots indicated that parameter recovery was not perfect. The wider posteriors for CNF contrasts with prior literature where amortized inference often underestimates uncertainty [hermans2022trust]. Despite these differences, both methods captured the inference data accurately and showed strong predictive performance, as evidenced by low and similar RMSEs. Their application in this setting shows that amortized methods may struggle with calibration, yet still provide strong predictive performance and exact MCMC based method estimate the bulk of the posterior well while underestimating its width caused by too less parameter space explorance.
Using sparsely sampled datasets with different missingness patterns did not change the qualitative estimation results of both methods. But they inherently treat sparse datasets differently. While CNF methods can accommodate such dataset variability, amortizing over all relevant combinations of time points may become computationally expensive and risks unreliable extrapolation if not carefully designed. This challenge is specific to setups that aim for broad amortization; alternatively, one can train CNFs for fixed missingness patterns. In contrast, PF methods naturally handle non-equidistant and partially missing data via their sequential simulation and weighting scheme, which decomposes the likelihood over time points and observables without requiring amortization or any other adaptation of the method.
Afterwards we introduced a reparametrization of the two-variant SEIR model that combined or transformed parameters based on their observed dependencies. This adjustment resulted in better alignment between CNF and PF posterior estimates and markedly increased PF mixing. These improvements suggest that the posterior geometry, rather than methods alone, leads to robust inference in the parameter space and model choices should be investigated as thoroughly as tuning of methods. CNF achieved worse calibration than in the full model case but its recovery plots demonstrated good parameter recovery. Both methods showed robustness to practical non-identifiabilities in their quality of fit and predictive power.
Finally, applying both methods to a real dataset from a longitudinal cohort study in Ethiopia confirmed the simulation study findings and yielded broadly similar posterior approximations and epidemic trajectories. Although obtained MAP estimates differed from previously published parameter values, especially for CNFs, both methods captured the data dynamics and produced model fits with RMSEs close to or improving upon the deterministic analysis [ethiopian_data]. By demonstrating their effectiveness in a real-world setting, our analysis underscores the potential of these methods to inform public health strategies and decision-making processes.
We also compared the computational efficiency of the two methods. To ensure a meaningful comparison, both methods were calibrated to produce the same number of samples from the parameter posterior, with CNF’s neural network trained on a CPU. Despite this constraint, CNF is roughly 10 times faster than PF, due to PFs sequential nature and the higher number of simulator calls. The bulk of CNF’s computational time is spent on training, but once trained, inference can be performed in mere seconds. CNF’s amortization allows multiple inferences, providing, in this scenario, a significant advantage over PF, which requires separate computations for each run. To achieve greater speed, CNF can make use of a parallelized simulator or training the neural network on a GPU. Meanwhile, PF could benefit from the use of more advanced filtering algorithms such as guided PFs [snyder_2011, chopin_2020] or using correlated particles [wiqvist_2021] to avoid degeneracy or too high particle numbers. Switching to a more advanced outer MCMC sampler such as a Gibbs sampler [chopin2015] or a Metropolis-Adjusted Langevin algorithm [lowe2023accelerating] and parallelization over particles could speed-up the inference procedure as well. Additionally, the lack of amortization renders assessment of calibration and hyperparameter tuning for PF too computationally expensive.
The scope of this work leads to several limitations. The observation model employed additive Gaussian noise for clarity and comparability; overdispersed counts or serological adjustments may be better represented by negative‑binomial or beta‑binomial formulations in other settings. Our use of an Euler–Maruyama integration with positivity clipping introduces discretization error whose interaction with inference can lead to numerical instabilities near boundaries or in small populations. Algorithmic choices were guided by the ease of their applicability in praxis: We focused on a bootstrap PF within PMMH and a tuned CNF architecture with consistent hyperparameters and fixed simulation budget across our datasets. Alternative PF proposals (e.g., guided/auxiliary filters or tempering within SMC) and alternative SBI estimators (likelihood or ratio estimation) or more exhaustive hyperparameter-tuning could shift efficiency–accuracy trade‑offs. To avoid out‑of‑distribution failure modes, CNFs were trained per experiment rather than globally amortized across all data regimes, which tempers their wall‑clock advantage in this study; broader amortization remains attractive but demands careful conditioning on meta‑data such as time‑grid irregularity and noise scales [radev2023bayesflow, wang2024missing].
In the SBI literature, recent advances focus on workflows that narrow the gap between neural SBI and Monte Carlo by using learned proposals inside sequential Monte Carlo or MCMC [papamakarios2019sequential, hermans2020likelihood]. Variational and amortized SMC approaches train proposal families to reduce weight degeneracy and improve effective sample sizes; conversely, particle methods can provide high‑quality samples to bootstrap or regularize neural estimators. Looking forward, future work in this context of hybrid approaches could focus on a flow trained in the spirit of neural posterior estimation that proposes parameter values to a PMMH kernel or guide latent‑state proposals. These methods could combine the tail coverage and speed of neural estimators with the asymptotic correctness and model flexibility of pseudo‑marginal schemes. Recent generative modeling advancements, i.e., flow matching and diffusion‑style transports, offer further avenues for robust posterior learning with stronger inductive biases and improved stability [lipman2022flow, draxler2024free, arruda2025]. Additionally, the use of adaptive prior schemes where the prior is updated in light of the available data, through approaches such as empirical Bayes, hierarchical modeling, or iterative refinement, may help to mitigate issues of non-identifiability while improving robustness. Such strategies may enhance performance in particularly challenging applications that use real world data where information is both sparse and noisy.
In summary, our comparison suggests that both CNF and PF are viable inference strategies for complex epidemic models, with different focus on parameter space exploration and data fitting. CNF tends to provide a more extensive posterior, whereas PF typically produces tighter distributions. Moreover, CNF’s substantially lower computational cost and amortization capabilities, achieving approximately 10 times faster performance than PF under comparable conditions, makes CNF especially advantageous for large-scale or time-sensitive applications. In contrast, PF maintain exactness (in the pseudo‑marginal sense) under model changes at the cost of potential tail under‑exploration, rendering PF a robust choice across models. The use of sparsely sampled data with different missingness patterns highlights the applicability of both methods and their robustness to different data settings. By clarifying the trade‑offs and giving insights into the performance of these two practical inference methods, this work provides the groundwork for evaluating inference engines ready for uncertainty‑aware epidemic forecasting and real-time support of public-health decisions.
Author Contributions
V.W.: Conceptualization, Application PF, Analysis PF, Visualization, Writing - original draft, Writing - review & editing. N.W.: Conceptualization, Application CNF, Analysis CNF Visualization, Writing - original draft, Writing - review & editing. L.C.: Software PF M.K.: Supervision, Funding acquisition, Writing - review & editing J.H.: Supervision, Project administration, Funding acquisition, Writing - review & editing
Funding
This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research under Germany’s Excellence Strategy (EXC 2151—390873048), by the European Union via ERC grant INTEGRATE (grant no 101126146) to J.H. and by the University of Bonn via the Bonn Center for Mathematical Life Sciences and the Schlegel Professorship of J.H.). This work received further funding from the Initiative and Networking Fund of the Helmholtz Association (grant agreement number KA1-Co-08, Project LOKI-Pandemics).
Appendix A Supplementary Information
A.1 Model Equations for the Stochastic Models
SIR Model
Set of stochastic differential equations describing the time evolution of individuals in the compartments of the two variant SEIR model.
$dS(t)$ $=-\beta\frac{S(t)I(t)}{N(t)}dt-\sqrt{\beta\frac{S(t)I(t)}{N(t)}}dB_{1}(t)$
$dI(t)$ $=\beta\frac{S(t)I(t)}{N(t)}dt-\gamma I(t)dt+\sqrt{\beta\frac{S(t)I(t)}{N(t)}}dB_{1}(t)-\sqrt{\gamma I(t)}dB_{2}(t)$
$dR(t)$ $=\gamma I(t)dt+\sqrt{\gamma I(t)}dB_{2}(t)$
where $B_{j},\,j\in\{1,...,2\}$ are independent Brownian Motions. Parameters with respective prior distributions and bounds are given in Table S3.
Two-Variant SEIR Model
Set of stochastic differential equations describing the time evolution of individuals in the compartments of the two variant SEIR model.
$dS(t)$ $=-\mathrm{\beta}_{\text{wt}}\frac{S(t)\mathrm{I}_{\text{wt}}(t)}{N(t)}dt-\beta_{var}\frac{S(t)(I_{var}(t)+I_{both}(t))}{N(t)}dt$
$\quad-\sqrt{\beta_{wt}\frac{S(t)I_{wt}(t)}{N(t)}}dB_{1}(t)-\sqrt{\beta_{var}\frac{S(t)(I_{var}(t)+I_{both}(t))}{N(t)}}dB_{7}(t)$
$dE_{wt}(t)$ $=\beta_{wt}\frac{S(t)I_{wt}(t)}{N(t)}dt-\kappa_{wt}E_{wt}(t)dt+\sqrt{\beta_{wt}\frac{S(t)I_{wt}(t)}{N(t)}}dB_{1}(t)-\sqrt{\kappa_{wt}E_{wt}(t)}dB_{2}(t)$
$dI_{wt}(t)$ $=\kappa_{wt}E_{wt}(t)dt-\gamma_{wt}I_{wt}(t)dt+\sqrt{\kappa_{wt}E_{wt}(t)}dB_{2}(t)-\sqrt{\gamma_{wt}I_{wt}(t)}dB_{3}(t)$
$dR_{wt}(t)$ $=\gamma_{wt}I_{wt}(t)dt-\beta_{var}\frac{R_{wt}(t)(I_{var}(t)+I_{both}(t))}{N(t)}dt$
$\quad+\sqrt{\gamma_{wt}I_{wt}(t)}dB_{3}(t)-\sqrt{\beta_{var}\frac{R_{wt}(t)(I_{var}(t)+I_{both}(t))}{N(t)}}dB_{4}(t)$
$dE_{var}(t)$ $=\beta_{var}\frac{S(t)(I_{var}(t)+I_{both}(t))}{N(t)}dt-\kappa_{var}E_{var}(t)dt$
$\quad+\sqrt{\beta_{var}\frac{S(t)(I_{var}(t)+I_{both}(t))}{N(t)}}dB_{7}(t)-\sqrt{\kappa_{var}E_{var}(t)}dB_{8}(t)$
$dI_{var}(t)$ $=\kappa_{var}E_{var}(t)dt-\frac{\gamma_{var}}{1.35}I_{var}(t)dt+\sqrt{\kappa_{var}E_{var}(t)}dB_{8}(t)-\sqrt{\frac{\gamma_{var}}{1.35}I_{var}(t)}dB_{9}(t)$
$dR_{var}$ $=\frac{\gamma_{var}}{1.35}I_{var}(t)dt+\sqrt{\frac{\gamma_{var}}{1.35}I_{var}(t)}dB_{9}(t)$
$dE_{both}(t)$ $=\beta_{var}\frac{R_{wt}(t)(I_{var}(t)+I_{both}(t))}{N(t)}dt-\kappa_{var}E_{both}(t)dt$
$\quad+\sqrt{\beta_{var}\frac{R_{wt}(t)(I_{var}(t)+I_{both}(t))}{N(t)}}dB_{4}(t)-\sqrt{\kappa_{var}E_{both}(t)}dB_{5}(t)$
$dI_{both}(t)$ $=\kappa_{var}E_{both}(t)dt-\gamma_{var}I_{both}(t)dt+\sqrt{\kappa_{var}E_{both}(t)}dB_{5}(t)-\sqrt{\gamma_{var}I_{both}(t)}dB_{6}(t)$
$dR_{both}(t)$ $=\gamma_{var}I_{both}(t)dt+\sqrt{\gamma_{var}I_{both}(t)}dB_{6}(t),$
where $B_{j},\,j\in\{1,...,7\}$ are independent Brownian Motions. Parameters with respective prior distributions and bounds are given in Table S10.
A.2 Computation of standard deviation for the noise model
The standard deviations of the noise model introduced in Section 2.3 were derived from binomial count data. This section outlines the rationale and computation procedure in detail.
When conducting Bernoulli trials, such as testing individuals in a population, the sum of positive test is adequately modeled by a binomial distribution, provided the trials are independent and the success probability remains constant. If the number $n$ of such Bernoulli trials conducted is large or the success probability $p$ is neither exceedingly low nor high, the resulting binomial distribution $B(n,p)$ is well approximated by a normal distribution $\mathcal{N}(np,np(1-p))$, thanks to the de Moivre-Laplace Theorem, a special case of the central limit Theorem [supp-walker1985moivre]. This approximation allows us to compute the standard deviation of the measurement noise directly from the count data and the known population size, and to incorporate it into the noise model.
However, this approximation may become inaccurate for small sample sizes or highly skewed probabilities, where alternative methods or exact distributions might be preferable. When generating our synthetic data, we chose $n=500$, which is sufficiently large to justify the Gaussian approximation.
A.3 Practical non-identifiability of the two-variant SEIR model
In our comparison we observed correlated parameter estimates leading to posterior misalignment between the methods. In several parameter combinations, the posterior mass was concentrated along low‑dimensional manifolds in parameter space, indicating practical non‑identifiability. This manifold can be approximated in terms of the following reparameterization into epidemiologically interpretable quantities:
[…]
where $r_{0}$ is the basic reproduction number, $e_{0}$ the mean latent-plus-infectious period, i.e., mean time an individual spends from catching the infection until it stops being infectious, and $s_{0}$ the effective detection factor integrated over the average infectious period, so how many cases are likely to be detected per infectious person, averaged over their infectious period.
A representative example is shown in Figure S2. The substantial overlap between the trajectories shows that combinations of widely different parameters can produce virtually identical epidemic dynamics.
A.4 Tuning of Particle Filter
Particle filters combined with MCMC algorithms are convincing considering their theoretical implications, but in practice tuning the performance of these algorithm can take a lot of work. For a detailed review of efficient implementations of marginal MCMC sampling, the reader may refer to [supp-doucet_2015], a numerical study on tuning PMMH is given in Chapter 16 of [supp-chopin_2020].
A.4.1 Hyperparamters and diagnostics
After deciding on a inner particle filter and an outer MCMC scheme, the only but also most crucial choice is the number of particles $N$. Although any number of particles $N\geq 1$ leaves the target density invariant and theoretically leads to the correct results, the number of particles highly affects the variance of the likelihood estimate and the algorithm’s ability to converge to the target distribution. Hence, choosing $N$ resembles the trade-off between a lower variance in the likelihood estimator, efficiency, and a higher computational cost. Based on the empirical results of experiments in [supp-chopin_2020] and theoretical considerations in [supp-doucet_2015] the number of particles should satisfy the following two criteria
  • The variance of the log-likelihood estimate is less than 1.
  • There exists more than one unique path of particles and their ancestors over time.
Both criteria are satisfied for the SIR and SEIR model with a particle number of 100 (Figure S3, when evaluated around the data generating parameters. To avoid particle degeneracy issues as much as possible in the two-variant SEIR model, we use a higher number of $N=200$ for that model.
In addition, we evaluated the convergence of the chains based on an improvement of the traditional Gelman-Rubin $\hat{R}$ criterium [supp-vehtari2021rank], where a value close to 1.0 is considered good convergence. To assess the mixing properties of the chain, we also monitored the effective samples size with a maximal lag for the autocorrelation of 250. For a given estimand, it is recommended that the ESS should be at least $100\cdot\text{nchains}$ [supp-vehtari2021rank]. Both criteria were evaluated using the last 10,000 samples of the chains.
A.4.2 Resampling in Importance Sampling
An additional choice, that needs to be made, when using the Bootstrap filter with resampling in it, is the resampling scheme. As in importance resampling algorithms in particle filters resampling increases the variance in the past because ancestors are resampled. However, the more diverse particles decrease the variance in the present. Hence, resampling itself can be seen as sacrificing the past to save the present and is especially helpful when low weights are present. If all weights are similar, then resampling increases the variance without helping. Without resampling, the Particle Filter would fall back to a single importance sampling step and be prone to the curse of dimensionality. Conversely, the option of resampling every time introduces a step with a computational complexity of $O(N)$ at every time step and might add additional variance if the weights are all similar. Therefore, in practice, one would not resample in every iteration step, but once the variability of the weights becomes too large. Commonly, one calculates the effective sample size (ESS) to estimate this variability and triggers resampling if it falls a certain threshold $\mathrm{ESS_{min}}$, often $N/2$ is used [supp-chopin_2020].
Let $\{w^{(i)},X^{(i)}\}_{i=1}^{N}$ be the set of weighted samples following the target distribution with density $q(x)$ and $M^{(i)}$ denote the number of offspring associated with $X^{(i)}$. The resampling process is equivalent to sampling $M^{(1:N)}=(M^{(1)},...,M^{(N)})$. Different schemes are proposed in the literature [supp-doucet_particle_tutorial, supp-chopin_2020] to obtain unbiased approximations via resampling, including multinomial, residual and systematize resampling. Among those, the systematic resampling is the best as it is easy to implement, very fast, and outperforms other schemes in terms of lower-variance estimators. This method performs sampling from a multinomial distribution with less variance.
Naively, one samples i.i.d. random variables $\{U_{i}\}_{i=1,...,N}$ from $\mathcal{U}([0,1))$ and then applies the inverse cumulative distribution function to transform them to the desired samples $M^{(i)}=\mathrm{cdf}^{-1}(U_{i})$. A good strategy to reduce the variance of this approach is to replace the i.i.d. values with values covering $[0,1)$ more regularly. To achieve this, one simulates a single uniformly distributed random variable $u\sim\mathcal{U}([0,\frac{1}{N}))$ and sets $U_{i}=u+\frac{i-1}{N}$ for i=1,,Ni=1,...,N. Then we obtain $M^{(i)}$ by
$M^{(i)}=\left|\left\{U_{j}:\sum_{k=1}^{i-1}w^{(k)}\leq\ U_{j}<\sum_{k=1}^{i}w^{(k)}\right\}\right|,$
where we use the convention $\sum_{k=1}^{0}:=0$
For a more thorough comparison of different resampling methods, the reader may refer to [supp-Doucet_SMC, supp-resampling_schemes_comparison, supp-doucet_particle_tutorial].
A.5 Hyperparameters for Amortized Bayesian Inference
Hyperparameters for the Normalizing-Flows workflow were mostly fixed across all experiments. Only the summary dimension was varied between models. For detailed configurations, see Tables S28, S29, S30, and S31.
A.6 Synthetic Datasets
In this work, we used two stochastic compartmental models for a simulation study, where for each model we created several datasets. For the standard SIR model we created two densely sampled (observation time points every 10 days) sets with simulations using the parameters in Table S4. For the two variant SEIR model we used two sets of parameters (Table S20) and for each parameter set generated two densely sampled (observation time points every 7 days) datasets. This results in 4 datasets named ”Dense aba-b”, where $a$ refers to the used parameter set and $b$ to the simulation round from that parameter set.
To mimic realistic data sparsity we drew three integers controlling the number of infection‑count timepoints, the number of seroprevalence timepoints, and the number of timepoints observed by both modalities. Then we randomly selected timepoints from the dense datasets according to these numbers. We repeated this procedure three times for each of the two dense datasets corresponding to the first parameter vector. This results in 6 sparse datasets named ”Sparse abca-b-c, where $a$ indicates the parameter set, $b$ the simulation round, $c$ the subsampling round.
Appendix B Supplementary Tables
Table S1: SIR model description Compartments and their respective description for the SIR model.
Compartment Description
$\mathrm{S}_{\text{}}$ Susceptible; individuals may become infected upon contact with Infectious individuals.
$\mathrm{I}_{\text{}}$ Infectious; individuals infected with the disease.
$\mathrm{R}_{\text{}}$ Recovered; individuals who have recovered from an infection.
Table S2: Parameters and their description for SIR model.
Parameter Description
$\beta$ Transmission risk for an individual susceptible to the disease.
$\gamma^{-1}$ Average infectious time.
Table S3: List of model parameters, bounds and priors for the SIR model
Parameter Lower bound Upper bound Prior
$\beta$ $0.95\cdot\gamma$ $5\cdot\gamma$ $\mathrm{Uniform}(0,1)$
$\gamma^{-1}$ 1 30 $\mathrm{Uniform}(1,30)$
Table S4: Parameters for synthetic datasets of the SIR model.
Dataset $\boldsymbol{\beta}$ $\boldsymbol{\gamma^{-1}}$
SIR 1 0.1 20
SIR 2 0.22 5
Table S5: MAP and RMSE for SIR model.
Dataset $\mathbf{\beta}$ $\mathbf{\gamma^{-1}}$ $\mathbf{RMSE_{infc}}$ $\mathbf{RMSE_{prev}}$
Dense 1 True Par. 0.1000 20.00 0.0193 0.0230
CNF 0.1026 18.99 0.0168 0.0201
PF 0.1021 19.32 0.0150 0.0121
Dense 2 True Par. 0.2200 5.00 0.0081 0.0298
CNF 0.2284 4.90 0.0074 0.0259
PF 0.2331 4.80 0.0090 0.0250
Table S6: Effective sample sizes (ESS) per parameter and data for the SIR model. ESS computed on the last 10,000 samples of the chains resulting from running the PF method on the standard SIR model and using a maximum lag size of 250 for the autocorrelation.
Dataset $\boldsymbol{\beta}$ $\boldsymbol{\gamma^{-1}}$
SIR 1 5189.6 5131.9
SIR 2 2259.2 2256.62
Table S7: $\mathbf{\hat{R}}$ diagnostics for the SIR model.
Model $\boldsymbol{\beta}$ $\boldsymbol{\gamma^{-1}}$
SIR 1 1.001 1.000
SIR 2 1.003 1.003
Table S8: Two-variant SEIR model description Compartments and their respective description for the two-variant SEIR model.
Compartment Description
$\mathrm{S}_{\text{}}$ Susceptible; individuals may become infected upon contact with Infectious individuals.
$\mathrm{E}_{\text{wt}}$ Exposed; individuals with no prior infection and infected with the wild-type but not yet infectious.
$\mathrm{E}_{\text{var}}$ Exposed; individuals with no prior infection and infected with the variant but not yet infectious.
$\mathrm{E}_{\text{wt·var}}$ Exposed; individuals with a prior infection from the wild-type and infected with the variant but not yet infectious.
$\mathrm{I}_{\text{wt}}$ Infectious; individuals with no prior infection, infected with the wild-type and currently infectious.
$\mathrm{I}_{\text{var}}$ Infectious; individuals with no prior infection, infected with the variant and currently infectious.
$\mathrm{I}_{\text{wt·var}}$ Infectious; individuals infected with the variant, infectious and who have previously been infected by the wild-type.
$\mathrm{R}_{\text{wt}}$ Recovered; individuals who have recovered from a wild-type infection, not currently infected.
$\mathrm{R}_{\text{var}}$ Recovered; individuals who have recovered from a variant infection, not currently infected.
$\mathrm{R}_{\text{wt·var}}$ Recovered; individuals recovered from a reinfection with the variant, after having have recovered from a prior wild-type infection.
Table S9: Parameters and their description for the two-variant SEIR model.
Parameter Description
$\beta$ Transmission risk for an individual susceptible to one of the two-variants.
$\kappa^{-1}$ Average latency period length for the wild-type and the variant.
$\gamma^{-1}$ Average infectious time for the wild-type.
s
Scaling factor, measuring the discrepancy between (measured
and reported) newly infected and actually infected.
$\mathrm{I}_{\text{0}}$ Number of individuals infected with the wild-type at start of the simulation.
$\mathrm{t}_{\text{var}}$ Time at which $n_{\textit{var}}$ individuals are introduced into the compartment $\mathrm{I}_{\text{var}}$.
Table S10: List of model parameters, bounds and priors for the two variant SEIR model. Prior distributions were taken from [supp-ethiopian_data].
Parameter Lower bound Upper bound Prior
$\gamma^{-1}$ 0 $\infty$ $\mathrm{Normal}(15.7,6.7)$
$\kappa$ 0 $\infty$ $\mathrm{Lognormal}(1.63,0.5)$
$\beta$ $0.95\cdot\gamma$ $4\cdot\gamma$ $\mathrm{Uniform}(0,1)$
$s$ 0.1 10 $\mathrm{Uniform}(0.1,10)$
$t_{var}$ 120 360 $\mathrm{Uniform}(120,360)$
$I0$ 10 1000 $\mathrm{Uniform}(10,1000)$
Table S11: Parameters for dense synthetic datasets of the reparametrized SEIR model.
Dataset $\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa^{-1}}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$
Dense 1-1 17 5 0.08 3 150 500
Dense 1-2 17 5 0.08 3 150 500
Dense 2-1 11.7 8.4 0.23 1.6 222 560
Dense 2-2 11.7 8.4 0.23 1.6 222 560
Table S12: MAP and RMSE for full SEIR model for dense datasets.
Dataset $\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa^{-1}}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I_{0}}$ $\mathbf{RMSE_{infc}}$ $\mathbf{RMSE_{prev}}$
Dense 1-1 True Par. 17.00 5.00 0.0800 3.00 150.00 500.00 0.0200 0.0245
CNF 15.47 4.93 0.0867 3.29 160.59 446.30 0.0237 0.0237
PF 15.62 5.04 0.0855 3.22 153.02 479.16 0.0228 0.0291
Dense 1-2 True Par. 17.00 5.00 0.0800 3.00 150.00 500.00 0.0214 0.0237
CNF 15.00 4.95 0.0902 3.28 161.67 374.69 0.0188 0.0215
PF 15.04 4.26 0.0890 3.25 165.57 366.60 0.0170 0.0234
Dense 2-1 True Par. 11.70 8.40 0.2300 1.60 222.00 560.00 0.0191 0.0170
CNF 12.39 7.13 0.2162 1.53 223.95 511.84 0.0198 0.0194
PF 11.55 7.90 0.2313 1.61 224.40 508.96 0.0242 0.0171
Dense 2-2 True Par. 11.70 8.40 0.2300 1.60 222.00 560.00 0.0197 0.0206
CNF 15.32 5.06 0.1774 1.26 222.82 541.31 0.0198 0.0210
PF 13.97 6.26 0.1947 1.34 221.29 592.14 0.0208 0.0184
Table S13: Effective sample sizes (ESS) per parameter and model for dense datasets. ESS computed on the last 10,000 samples of the chains resulting from running the PF method on the two-variant SEIR model and using a maximum lag size of 250 for the autocorrelation.
Parameter $\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$
Model
Dense 1-1 501.6 658.8 514.8 528.8 1174.2 1537.4
Dense 1-2 436.2 539.8 401.8 428.6 1493.4 1768.6
Dense 2-1 445.3 459.9 440.9 457.8 1676.1 1842.0
Dense 2-2 1549.2 1633.8 1608.7 1652.8 1954.5 1687.8
Table S14: $\mathbf{\hat{R}}$ diagnostics for the SEIR model with dense data.
Dataset $\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$
Dense 1-1 1.023 1.010 1.022 1.022 1.007 1.005
Dense 1-2 1.025 1.014 1.024 1.021 1.004 1.002
Dense 2-1 1.015 1.012 1.008 1.022 1.007 1.009
Dense 2-2 1.014 1.009 1.008 1.022 1.007 1.004
Table S15: Parameters for dense synthetic datasets of the reparametrized SEIR model. The sparse set Sparse x-y-z is subsampled from the dense set Dense x-y
Dataset $\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa^{-1}}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$
Sparse 1-1-1 17 5 0.08 3 150 500
Sparse 1-1-2 17 5 0.08 3 150 500
Sparse 1-1-3 17 5 0.08 3 150 500
Sparse 1-2-1 17 5 0.08 3 150 500
Sparse 1-2-2 17 5 0.08 3 150 500
Sparse 1-2-3 17 5 0.08 3 150 500
Table S16: MAP and RMSE for full SEIR model for sparse datasets.
Dataset $\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa^{-1}}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I_{0}}$ $\mathbf{RMSE_{infc}}$ $\mathbf{RMSE_{prev}}$
Sparse 1-1-1 True Par. 17.00 5.00 0.0800 3.00 150.00 500.00 0.0229 0.0220
CNF 15.69 4.29 0.0871 3.16 171.04 331.82 0.0302 0.0155
PF 14.65 5.52 0.0933 3.29 165.29 349.35 0.0371 0.0289
Sparse 1-1-2 True Par. 17.00 5.00 0.0800 3.00 150.00 500.00 0.0200 0.0215
CNF 15.95 5.06 0.0841 2.91 149.45 440.37 0.0298 0.0357
PF 14.44 5.12 0.0913 3.12 147.55 443.22 0.0247 0.0195
Sparse 1-1-3 True Par. 17.00 5.00 0.0800 3.00 150.00 500.00 0.0201 0.0101
CNF 14.26 5.27 0.0945 3.65 174.46 395.00 0.0240 0.0320
PF 13.17 5.47 0.1023 3.81 187.08 335.15 0.0292 0.0306
Sparse 1-2-1 True Par. 17.00 5.00 0.0800 3.00 150.00 500.00 0.0229 0.0221
CNF 16.23 4.76 0.0854 3.03 168.65 345.33 0.0281 0.0335
PF 12.82 3.62 0.1016 3.74 180.59 259.94 0.0251 0.0396
Sparse 1-2-2 True Par. 17.00 5.00 0.0800 3.00 150.00 500.00 0.0222 0.0272
CNF 14.74 5.05 0.0913 3.26 153.58 363.09 0.0107 0.0269
PF 15.03 3.96 0.0891 3.23 165.95 290.72 0.0262 0.0183
Sparse 1-2-3 True Par. 17.00 5.00 0.0800 3.00 150.00 500.00 0.0134 0.0229
CNF 16.74 4.48 0.0838 3.10 171.60 355.96 0.0108 0.0223
PF 16.49 4.80 0.0836 3.04 159.82 370.15 0.0186 0.0157
Table S17: Effective sample sizes (ESS) per parameter and model for sparse datasets. ESS computed on the last 10,000 samples of the chains resulting from running the PF method on the two-variant SEIR model and using a maximum lag size of 250 for the autocorrelation.
Parameter $\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$
Model
Sparse 1-1-1 313.9 536.5 302.4 364.4 1550.2 1641.1
Sparse 1-1-2 208.8 938.9 211.3 261.2 907.0 757.0
Sparse 1-1-3 292.1 771.5 287.9 304.3 914.6 1157.6
Sparse 1-2-1 252.0 1121.2 258.3 308.7 463.0 587.4
Sparse 1-2-2 326.5 874.5 303.6 380.3 738.2 1216.5
Sparse 1-2-3 301.1 1234.2 309.9 343.2 760.9 873.2
Table S18: $\mathbf{\hat{R}}$ diagnostics for the SEIR model with sparse data.
Parameter Set $\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$
Sparse 1-1-1 1.064 1.034 1.065 1.052 1.002 1.005
Sparse 1-1-2 1.154 1.003 1.143 1.099 1.008 1.028
Sparse 1-1-3 1.100 1.019 1.100 1.085 1.018 1.007
Sparse 1-2-1 1.112 1.007 1.109 1.076 1.036 1.030
Sparse 1-2-2 1.067 1.016 1.068 1.055 1.024 1.009
Sparse 1-2-3 1.027 1.007 1.028 1.016 1.011 1.007
Table S19: List of model parameters, bounds and priors for the reparametrized two variant SEIR model
Parameter Lower bound Upper bound Prior
$r0$ 0.95 4.0 $\mathrm{Uniform}(0.95,4.0)$
$e0$ 6 30 $\mathrm{Uniform}(6,30)$
$s0$ 1 100 $\mathrm{Uniform}(1,100)$
$t_{\mathrm{var}}$ 120 360 $\mathrm{Uniform}(120,360)$
$I0$ 10 1000 $\mathrm{Uniform}(10,1000)$
Table S20: Parameters for dense synthetic datasets of the reparametrized SEIR model.The datasets are the same as for the full model.
Dataset $\mathbf{r0}$ $\mathbf{e0}$ $\mathbf{s0}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$
Dense 1-1 1.36 22 51 150 500
Dense 1-2 1.36 22 51 150 500
Table S21: MAP and RMSE for reparametrized SEIR model and dense datasets.
Dataset $\mathbf{r0}$ $\mathbf{e0}$ $\mathbf{s0}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$ $\mathbf{RMSE_{infc}}$ $\mathbf{RMSE_{prev}}$
Dense 1-1 True Par. 1.36 22.00 51.00 150.00 500.00 0.0200 0.0245
CNF 1.34 21.33 50.27 145.46 489.71 0.0216 0.0236
PF 1.34 20.57 51.21 156.94 467.30 0.0319 0.0271
Dense 1-2 True Par. 1.36 22.00 51.00 150.00 500.00 0.0214 0.0237
CNF 1.35 20.24 48.11 156.46 375.09 0.0184 0.0258
PF 1.34 19.69 49.17 161.32 382.68 0.0212 0.0236
Table S22: Effective sample sizes (ESS) per parameter and model. ESS computed on the last 10,000 samples of the chains resulting from running the PF method on the reparametrized two-variant SEIR model and using a maximum lag size of 250 for the autocorrelation.
Parameter $\mathbf{r0}$ $\mathbf{e0}$ $\mathbf{s0}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$
Dense 1-1 2409.8 24533.8 2370.5 2276.2 2382.2
Dense 1-2 2675.0 2413.2 2495.5 2537.6 2571.3
Table S23: $\mathbf{\hat{R}}$ diagnostics for the reparametrized SEIR model with dense data.
Dataset $\mathbf{r0}$ $\mathbf{e0}$ $\mathbf{s0}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$
Dense 1-1 1.002 1.001 1.005 1.003 1.002
Dense 1-2 1.002 1.002 1.002 1.002 1.002
Table S24: MAP and RMSE for full SEIR model for real dataset.
Dataset $\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa^{-1}}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I_{0}}$ $\mathbf{RMSE_{infc}}$ $\mathbf{RMSE_{prev}}$
Real data Pub. Par. 16.70 5.00 0.0800 2.30 184.50 635.00 0.0737 0.0702
CNF 14.80 4.14 0.0882 3.16 183.04 550.16 0.1251 0.0928
PF 15.03 4.67 0.0858 3.36 160.35 752.08 0.0856 0.0606
Table S25: Effective sample sizes (ESS) per parameter and model for the real dataset. ESS computed on the last 10,000 samples of the chains resulting from running the PF method on the two-variant SEIR model and using a maximum lag size of 250 for the autocorrelation.
Parameter $\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$
Model
Real data 423.8 1312.0 407.3 959.2 798.6 734.5
Table S26: $\mathbf{\hat{R}}$ diagnostics for the SEIR model with real data.
Parameter Set $\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$
Real 1.034 1.004 1.033 1.007 1.010 1.006
Table S27: Comparison of computation times: Accumulated CPU-time in minutes for both methods applied to a dense dataset. For the PF we ran 4 chains on 4 cores in parallel, the CNF was parallelized over 20 cores.
Datasets PF - normal prior PF - unif. prior CNF - normal prior CNF - unif. prior
SIR 1 1163.33 - 106.72 -
Dense 1-1 2095.98 2104.97 252.96 252.48
Sparse 1-1-1 1798.1 - 132.29 -
Table S28: Hyperparameters Summary Network: SequenceNetwork
Hyperparameter Setting
Summary Dimensions $2\times\texttt{parameter\_size}+2$
LSTM Units 64
Convolutional Layers 2
Table S29: Invertible Network: InvertibleNetwork
Coupling parameter Setting
Coupling Type Spline
Coupling Layers 8
Table S30: Simulation Configuration.
Hyperparameter Setting
Training Mode Offline
Simulations 100,000
Batch Size 32
Validation Simulations 400
Epochs 100
Table S31: Calibration Setup
Hyperparameter Setting
Datasets 2,000
Inference Draws 100 per dataset
Appendix C Supplementary Figures
Figure S1: Posterior Geometry 3D Plot of selected pairwise density plots. The mountain pass geometry on the left makes inference of parameters more challenging compared to the mountain peak geometry on the right.
Figure S2: Illustration of a practical identifiability manifold in the two-variant SEIR model. Orange: simulations from the fixed parameter set (17, 5, 0.08, 3, 150, 500)(17,\,5,\,0.08,\,3,\,150,\,500). Blue: simulations from parameter sets with $\gamma^{-1}$ ranging from $1$ to 20 and other parameters adjusted according to the reparametrization in (A.3) The near-complete overlap of trajectories demonstrates that distinct parameter combinations can produce almost indistinguishable epidemic curves.
Figure S3: Metrics for tuning of the particle number. (a) and (c) visualize the variance of the log-likelihood estimate across different particle numbers for the SIR and SEIR model. (b) and (d) depict the ancestor trees of the particles over time for the SIR and the SEIR model.
Figure S4: Results of the two-variant SEIR model with dense dataset Dense 1-1.
A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade. B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
Figure S5: Results of the two-variant SEIR model with dense dataset Dense 1-2.
A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade. B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
Figure S6: Results of the two-variant SEIR model with dense dataset Dense 2-1.
A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade. B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
Figure S7: Results of the two-variant SEIR model with dense dataset Dense 2-2.
A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade. B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
Figure S8: Results of the two-variant SEIR model with sparse dataset Sparse 1-1-1.
A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade. B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
Figure S9: Results of the two-variant SEIR model with sparse dataset Sparse 1-1-2.
A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade. B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
Figure S10: Results of the two-variant SEIR model with sparse dataset Sparse 1-1-3.
A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade. B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
Figure S11: Results of the two-variant SEIR model with sparse dataset Sparse 1-2-1.
A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade. B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
Figure S12: Results of the two-variant SEIR model with sparse dataset Sparse 1-2-2.
A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade. B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
Figure S13: Results of the two-variant SEIR model with sparse dataset Sparse 1-2-3.
A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade. B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
Figure S14: Results of the two-variant SEIR model with sparse dataset Real data.
A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade. B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
Figure S15: Results of the reparametrized two-variant SEIR model with dense dataset Dense 1-1.
A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade. B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
Figure S16: Results of the reparametrized two-variant SEIR model with dense dataset Dense 1-2.
A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade. B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
Figure S17: Results of the SIR model with dataset SIR 1.
A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade. B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
Figure S18: Results of the SIR model with dataset SIR 2.
A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade. B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
Figure S19: Results of the SIR model with dataset SIR 2.
A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade. B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
Figure S20: ECDF Calibration plots for the full SEIR model and dense datasets.
(a) Dense 1-1 (b) Dense 1-2 (c) Dense 1-3 (d) Dense 2-1 (e) Dense 2-2 (f) Dense 2-3
Figure S21: ECDF Calibration plots for the full SEIR model on sparse and the real datasets.
(a) Sparse 1-1-1 (b) Sparse 1-1-2 (c) Sparse 1-1-3 (d) Sparse 1-2-1 (e) Sparse 1-2-2 (f) Sparse 1-2-3 (g) Real data
Figure S22: ECDF Calibration plots for the reparametrized SEIR model and dense datasets.
(a) Dense 1-1 (b) Dense 1-2
Figure S23: ECDF Calibration plots for the SIR model.
(a) SIR 1 (b) SIR 2
Figure S24: Parameter recovery for the full SEIR model and dense datasets.
(a) Dense 1-1 (b) Dense 1-2 (c) Dense2-1 (d) Dense 2-2
Figure S25: Parameter recovery for for the full SEIR model on sparse and the real datasets.
(a) Sparse 1-1-1 (b) Sparse 1-1-2 (c) Sparse 1-1-3 (d) Sparse 1-2-1 (e) Sparse 1-2-2 (f) Sparse 1-2-3 (g) Real data
Figure S26: Parameter recovery for the reparametrized SEIR model and dense datasets.
(a) Dense 1-1 (b) Dense 1-2
Figure S27: Parameter recovery for for the SIR model.
(a) SIR 1 (b) SIR 2

### v3

@@
-Assessment of Simulation-based Inference Methods for Stochastic Compartmental Models
+Assessment of Simulation-based Inference Methods for Stochastic Compartmental Models in Epidemiological Research
 Abstract
@@
 Such models must be accompanied by methods for estimating parameters in order to generate fast nowcasts and short-term forecasts that can inform public health decisions.
-This paper presents a comparison of two advanced Bayesian inference methods: 1) pseudo-marginal particle Markov chain Monte Carlo, short Particle Filters (PF), and 2) Conditional Normalizing Flows (CNF).
-We investigate their performance on two commonly used compartmental models: a classical Susceptible-Infected-Recovered (SIR) model and a two-variant Susceptible-Exposed-Infected-Recovered (SEIR) model, complemented by an observation model that maps latent trajectories to empirical data.
+This paper presents a comparison of two advanced Bayesian inference methods: 1) pseudo-marginal particle Markov chain Monte Carlo, using an unbiased likelihood estimate obtained by Particle Filter (PF), and 2) Conditional Normalizing Flows (CNF).
+We investigate their performance on three commonly used compartmental models: A classical Susceptible-Infected-Susceptible (SIS), a Susceptible-Infected-Recovered (SIR) model and a two-variant Susceptible-Exposed-Infected-Recovered (SEIR) model, complemented by an observation model that maps latent trajectories to empirical data.
 Addressing the challenges of intractable likelihoods for parameter inference in stochastic settings, our analysis highlights how these likelihood-free methods provide accurate and robust inference capabilities.
 The results of our simulation study further underscore the effectiveness of these approaches in capturing the stochastic dynamics of epidemics, providing prediction capabilities for the control of epidemic outbreaks.
-Results on an Ethiopian cohort study demonstrate operational robustness under real‑world noise and irregular data sampling.
+Results on an Ethiopian cohort study demonstrate operational robustness under real-world noise and irregular data sampling.
 To facilitate reuse and to enable building pipelines that ultimately contribute to better informed decision making in public health, we make code and synthetic datasets publicly available.
@@
 The COVID-19 pandemic underscored the critical importance of epidemic modeling research.
-Rapid, data-driven approaches have demonstrated their capacity to provide public health decision-makers with essential insights for mitigating emerging infectious disease outbreaks and controlling epidemic dynamics [ethiopian_data, doms2018assessing].
-Beyond COVID-19, other outbreaks such as Ebola and MPOX have likewise been designated as “public health emergencies of international concern” by the World Health Organization (WHO) [gpmb2024].
+Rapid, data-driven approaches have demonstrated their capacity to provide public health decision-makers with essential insights for mitigating emerging infectious disease outbreaks and controlling epidemic dynamics [Gudina et al., 2021; Doms et al., 2018].
+Beyond COVID-19, other outbreaks such as Ebola and MPOX have likewise been designated as “public health emergencies of international concern” by the World Health Organization (WHO) [Global Preparedness Monitoring Board, 2025].
 These crises and consequent public health challenges highlight the need for advanced modeling techniques capable of capturing the intricate dynamics of disease spread.
 Epidemiological and medical research revealed that the spread of infectious diseases is shaped by a variety of processes, including transmission dynamics, patient responses to treatment, and pathogen mutation.
-A wide spectrum of mathematical modeling approaches is used to describe the underlying dynamics, spanning from standard ordinary differential equation (ODE) models [McKendrick_Kermack, hethcote2000] to agent-based models [bruch2015agent].
+A wide spectrum of mathematical modeling approaches is used to describe the underlying dynamics, spanning from standard ordinary differential equation (ODE) models [Kermack et al., 1927; Hethcote, 2000] to agent-based models Bruch and Atwell [2015].
 On the one hand, ODE-based population models provide a fast to simulate framework, but fall short to describe the variability of events and the resulting complex, stochastic interactions, due to their deterministic nature.
-Therefore, stochastic models are essential for realistically representing infectious disease dynamics observed in real-world settings, especially in small populations or during the early stages of an outbreak [britton2010stochastic].
-On the other hand, purely stochastic models, such as agent-based and continuous-time Markov chain (CTMC) models [dragatz_book, infect_dis_modelling] provide highly detailed stochastic representations of individual-level processes, but come at the cost of prohibitive computational burden [petzold_2013].
-Mathematical models based on stochastic differential equations – so called stochastic (meta-)population models – can solve the trade-off between fast model simulations and keeping the stochasticity in the model [dragatz_paper].
+Therefore, stochastic models are essential for realistically representing infectious disease dynamics observed in real-world settings, especially in small populations or during the early stages of an outbreak [Britton, 2010].
+On the other hand, purely stochastic models, such as agent-based and continuous-time Markov chain (CTMC) models [Dragatz, 2010; Allen, 2017] provide highly detailed stochastic representations of individual-level processes, but come at the cost of prohibitive computational burden [Gillespie et al., 2013].
+Mathematical models based on stochastic differential equations – so called stochastic (meta-)population models – can solve the trade-off between fast model simulations and keeping the stochasticity in the model [Dragatz, 2006].
 In this study, we consider compartmental models described by stochastic differential equations (SDE).
-In compartmental models individuals are arranged in a finite number of mutually disjoint compartments and interaction happens through transfer from one compartment to another which is described by transition equations [Brauer2008].
+In compartmental models individuals are arranged in a finite number of mutually disjoint compartments and interaction happens through transfer from one compartment to another which is described by transition equations [Brauer, 2008].
 Their description in terms of SDEs can be seen as a mesoscopic view on the model dynamics.
-The SDEs are derived as a diffusion approximation to the microscopic, individual-based, discrete-state CTMC description [dragatz_paper, expansion_vanKampen, fearnhead2014inference] and in the large population limit resemble the macroscopic, population-based, (deterministic) ODE models [kurtz1970solutions].
-For a more detailed discussion on stochastic epidemic modeling see [infect_dis_modelling, dragatz_paper, AnderssonAndBritton] and references contained therein.
+The SDEs are derived as a diffusion approximation to the microscopic, individual-based, discrete-state CTMC description [Dragatz, 2006; van Kampen, 1976; Fearnhead et al., 2014] and in the large population limit resemble the macroscopic, population-based, (deterministic) ODE models [Kurtz, 1970].
+For a more detailed discussion on stochastic epidemic modeling see Allen [2017]; Dragatz [2006]; Andersson and Britton [2001] and references contained therein.
 A key challenge in mathematical-epidemiological modeling is determining the unknown model parameters that best explain the observed data.
@@
 Many methods have been developed to address the issue of intractable likelihoods and posterior distributions.
-Early attempts used data augmentation or employed Markov chain Monte Carlo (MCMC) methods [dragatz_book, gibson1998estimating, roberts1999bayesian].
-Later approaches sought to reduce computational cost through surrogate models such as Gaussian processes [swallow2022challenges] or linear noise approximations of the underlying Markov jump process [fintzi2022linear].
-Hybrid strategies combining data augmentation with MCMC have also been proposed [mckinley2014simulation, el2016bayesian].
-Sequential Monte Carlo methods were introduced as a flexible approach for analyzing stochastic state space models and performing Bayesian filtering tasks [doucet_sequential, chopin2013].
-These methods, also known as particle filters (PF), have been used to generate parameter proposals of an MCMC algorithm, leading to pseudo-marginal methods for Bayesian inference on model parameters [andrieu_roberts, andrieu_2010, chopin_2020], a popular method for exact Bayesian inference in stochastic models [chopin_2020, wiqvist_2021].
+Early attempts used data augmentation or employed Markov chain Monte Carlo (MCMC) methods [Dragatz, 2010; Gibson and Renshaw, 1998; O’Neill and Roberts, 1999].
+Later approaches sought to reduce computational cost through surrogate models such as Gaussian processes [Swallow et al., 2022] or linear noise approximations of the underlying Markov jump process [Fintzi et al., 2022].
+Hybrid strategies combining data augmentation with MCMC have also been proposed [McKinley et al., 2014; El Maroufy et al., 2016].
+Sequential Monte Carlo methods were introduced as a flexible approach for analyzing stochastic state space models and performing Bayesian filtering tasks [Doucet et al., 2000; Chopin et al., 2012].
+These methods, also known as particle filters (PF), have been used to produce an unbiased likelihood estimate for an MCMC algorithm, leading to pseudo-marginal methods for Bayesian inference on model parameters [Andrieu and Roberts, 2009; Andrieu et al., 2010; Chopin and Papaspiliopoulos, 2020], a popular method for exact Bayesian inference in stochastic models [Chopin and Papaspiliopoulos, 2020; Wiqvist et al., 2021].
 Instead of determining model parameters via sampling based on evaluations of the likelihood or an approximation to it, it is possible to conduct the inference procedure solely based on simulations.
-In the widely used technique of Approximate Bayesian Computing (ABC), simulated and observed data is compared based on summary statistics [rubin1984, minter_2019].
+In the widely used technique of Approximate Bayesian Computing (ABC), simulated and observed data is compared based on summary statistics [Rubin, 1984; Minter and Retkute, 2019].
 However, manually selected summary statistics can be uninformative and their selection poses a challenge on its own.
-Additionally, as these methods require massive numbers of simulations, more recent machine learning approaches have been adapted for simulation-based inference with neural networks [cranmer2020frontier], namely neural posterior estimation methods.
-These promising techniques try to circumvent summary statistics and directly approximate the distribution of the parameters from samples generated by the simulator.
-Therein, conditional normalizing flows (CNF) are an advanced method that uses invertible neural networks to generate samples from the posterior distribution of the parameters [radev2020bayesflow].
-PF and CNFs show great potential for Bayesian inference on stochastic models and are increasingly recognized within the epidemiological modeling community for providing probabilistic estimates of disease characteristics to effectively monitor pandemic outbreaks [radev2021outbreakflow, storvik2023].
+Additionally, as these methods require massive numbers of simulations, more recent machine learning approaches have been adapted for simulation-based inference with neural networks [Cranmer et al., 2020].
+These promising techniques often circumvent hand-crafted summary statistics and directly approximate the distribution of the parameters from samples generated by the simulator.
+Therein, conditional normalizing flows (CNF) are an advanced method that uses invertible neural networks to generate samples from the posterior distribution of the parameters [Papamakarios and Murray, 2016; Papamakarios et al., 2021].
+PF and CNFs show great potential for Bayesian inference on stochastic models and are increasingly recognized within the epidemiological modeling community for providing probabilistic estimates of disease characteristics to effectively monitor pandemic outbreaks [Radev et al., 2021; Storvik et al., 2023].
 However, their adoption has been limited by the lack of a comprehensive comparison assessing their relative performance.
 This work offers a praxis-driven comparison between both inference methods, assessing strengths and weaknesses of the particular approaches, as well as their alignment.
-In the first part, we described the two stochastic compartmental models, the well-known SIR model [McKendrick_Kermack] and a two-variant SEIR model, which was used to analyze the spread of COVID-19 in Ethiopia [ethiopian_data].
+In the first part, we described the three stochastic compartmental models, the well-known SIS and SIR models [Kermack et al., 1927] and a two-variant SEIR model, which was used to analyze the spread of COVID-19 in Ethiopia [Gudina et al., 2021].
 Together with the observational model bridging compartment trajectories with empirical data they serve as the foundation for our study.
 Moreover, we outline Bayesian inference for stochastic differential equation-based models, and explain the chosen techniques of CNF and PF together with the implementation of our comparison workflow.
-In the second part of the work, we outline the results from comparing the two inference methods on synthetically generated data for the two chosen compartmental models and additionally using real-data for the two-variant SEIR model.
-Both methods provide robust and reliable inference results on the stochastic version of the SIR model with synthetically generated data, validating their implementation at hand.
+In the second part of the work, we outline the results from comparing the two inference methods on synthetically generated data for the chosen compartmental models and additionally using real-data for the two-variant SEIR model.
+For the two smaller compartmental models, SIS and SIR, we compared the estimated posterior distributions with a reference posterior obtained by Hamiltonian Monte Carlo (HMC) sampling on a discretized version of the stochastic model.
+Both methods provide robust and reliable inference results on the stochastic versions of the SIS and SIR model with synthetically generated data, validating their implementation at hand.
 Analyzing the agreement of the estimated marginal posteriors and model fits based on different synthetic datasets for the two-variant SEIR model, including (partially) missing data, shows that both methods yield good fits to the data at hand.
@@
 Applying both methods to an reparametrized version with parameter dimension reduction can overcome these issues and improves posterior alignment.
-Inference based on the data from [ethiopian_data] shows the applicability of both methods to real-world data.
+Inference based on the data from Gudina et al.
+[2021] shows the applicability of both methods to real-world data.
 A discussion of the results, their limitations and further research directions follows in Section 6.
 2 Stochastic Compartmental Models
-Compartmental models are a mathematical framework to describe how individuals in a population move between different states and are fundamental for modeling infectious diseases [Brauer2008].
-Such a model is characterized by the vector of compartment sizes $X\in\mathbb{R}^{d_{X}}$ and the transitions of individuals between compartments.
-A transition from compartment $i$ to compartment $j$ can be described in terms of the stochiometric coefficient $s_{i,j}\in\mathbb{R}^{d_{x}}$, the vector of changes to the compartment sizes, and the rate at which the transition happens
-$a_{i,j}^{N}(X)=cX_{i}\sum_{k\in I_{i,j}}\frac{X_{k}}{N},$
-where $c$ denotes a positive parameter, $I_{i,j}$ is an index set containing all indices $k\neq i$ of compartments involved in that transition, and $N$ gives the total number of individuals in the population.
-Let $\mathcal{T}$ denote the set of all tuples $(i,j)$, such that a transition from $X_{i}$ to $X_{j}$ exists.
-Then the time evolution of the compartment sizes can be modeled using a multivariate stochastic differential equation [dragatz_paper]:
+Compartmental models are a mathematical framework to describe how individuals in a population move between different states and are fundamental for modeling infectious diseases [Brauer, 2008].
+Such a model partitions a population into $d_{X}$ mutually exclusive compartments and is characterized by the vector of compartment sizes $X\in\mathbb{R}^{d_{X}}$ together with the transitions of individuals between these compartments.
+Let $\mathcal{T}$ denote the set of admissible transitions.
+Each transition $(i,j)\in\mathcal{T}$ is characterized by a stoichiometric vector $s_{i,j}\in\mathbb{Z}^{d_{X}}$, which specifies the change in the compartment counts when the transition occurs.
+For the movement of one individual from compartment $i$ to compartment $j$, $s_{i,j}=e_{j}-e_{i}$, where $e_{i}$ denotes the $i$-th unit vector.
+The transition intensity is described by a propensity function $a_{i,j}^{N}:\mathcal{S}_{N}\to\mathbb{R}_{+}$, defined through
+$\Pr\bigl(X^{N}(t+h)=x+s_{i,j}\mid X^{N}(t)=x\bigr)=a_{i,j}^{N}(x)\,h+o(h),\qquad h\downarrow 0.$
+For infection-driven transitions, a common choice is a mass-action propensity of the form
+$a_{i,j}^{N}(X)=c_{i,j}X_{i}\sum_{k\in I_{i,j}}\frac{X_{k}}{N},$
+where $c_{i,j}>0$ is a rate parameter and $I_{i,j}$ contains the compartments contributing to the force of infection.
+Other transitions, such as recovery, are typically modeled by linear propensities, for example $a_{i,j}^{N}(x)=c_{i,j}x_{i}$.
+The resulting stochastic compartmental model is a continuous-time Markov jump process.
+In the large-population regime, it can be approximated by the multivariate stochastic differential equation (SDE) [Dragatz, 2006]:
 $dX(t)$ $=\mu(X(t))dt+\sigma(X(t))d\mathbf{B}(t)$
-$=\sum_{(i,j)\in\mathcal{T}}a_{i,j}(X(t))s_{i,j}dt+\sum_{(i,j)\in\mathcal{T}}\sqrt{a_{i,j}(X(t)))}s_{i,j}s_{i,j}^{T}d\mathbf{B}(t)$
+$=\sum_{(i,j)\in\mathcal{T}}a_{i,j}(X(t))s_{i,j}dt+\sum_{(i,j)\in\mathcal{T}}\sqrt{a_{i,j}(X(t))}s_{i,j}s_{i,j}^{T}d\mathbf{B}(t)$
 where $\mathbf{B}$ denotes a vector of independent Brownian motions.
-The representation of compartmental models in terms of multivariate SDEs can capture the randomness of disease spread in their diffusion part $\sigma(X(t))$ and retain a drift component $\mu(X(t))$ for representing the deterministic trend of the population [AnderssonAndBritton, dragatz_book].
+The representation of compartmental models in terms of multivariate SDEs can capture the randomness of disease spread in their diffusion part $\sigma(X(t))$ and retain a drift component $\mu(X(t))$ for representing the deterministic trend of the population [Andersson and Britton, 2001; Dragatz, 2010].
 This makes them a favorable choice for modeling disease outbreaks and dynamics.
-In this work, we consider two examples of SDE-based compartmental models, a simple standard Susceptible-Infected-Recovered (SIR) model [McKendrick_Kermack] and an extended multi-variant Susceptible-Exposed-Infected-Recovered (SEIR) model [hethcote2000] considering two variants of a single pathogen.
+In this work, we consider two examples of SDE-based compartmental models, a simple standard Susceptible-Infected-Recovered (SIR) model [Kermack et al., 1927] and an extended two-variant Susceptible-Exposed-Infected-Recovered (SEIR) model [Hethcote, 2000] considering two variants of a single pathogen.
 Figure 1: Schematic representations of the models and methods.
-(a) Graph of compartments and possible transitions with corresponding rate parameters for the SIR model.
-(b) Graph of compartments and possible transitions with corresponding rate parameters for the two‑variant SEIR model.
-(c) Workflow for the assessment of the Bayesian inference.
+(a) Graph of compartments and possible transitions with corresponding rate parameters for the Susceptible - Infected - Susceptible (SIS) model.
+(b) Graph of compartments and possible transitions with corresponding rate parameters for the Susceptible - Infected - Recovered (SIR) model.
+(c) Graph of compartments and possible transitions with corresponding rate parameters for the Susceptible-Exposed-Infected-Recovered (SEIR) model with virus variants wild-type (wt) and variant (var).
+(d) Workflow for the assessment of the Bayesian inference.
+2.1 SIS Model
+The simplest compartmental model is the SIS-model [Kermack et al., 1927] consisting of one Susceptible compartment $S$ and one Infectious compartment $I$, where people can transition between them (Figure 1(a)).
+This model can be used for infectious diseases that do not confer any long-lasting immunity, such as the common cold or influenza.
+Following simple mass-action incidence individuals get infected with a rate $\beta SI$ and leave the infective class at rate $\gamma I$ (Table 1).
+Table 1: Description of the SIS model. Table including all events, such as initialization and transition of individuals, with their respective rate and model parameters.
+Event Description Rate Parameter
+Initialization Set initial number of infected individuals - $\mathrm{I}_{0}$
+Infection $\mathrm{S}_{\text{}}\rightarrow\mathrm{I}_{\text{}}$ $\beta\frac{\mathrm{S}(t)\mathrm{I}(t)}{\mathrm{N}(t)}$ $\beta$
+Recovery $\mathrm{I}_{\text{}}\rightarrow\mathrm{S}_{\text{}}$ $\gamma\mathrm{I}(t)$ $\gamma^{-1}$
 2.2 SIR Model
-The standard SIR model (Figure 1(a)) considers three compartments: $S$ (Susceptible), $I$ (Infectious) and $R$ (Recovered).
+The standard SIR model (Figure 1(b)) considers three compartments: $S$ (Susceptible), $I$ (Infectious) and $R$ (Recovered).
 The model describes the number of individuals in these compartments given the transition parameters $\beta$ (the transmission rate) and $\gamma$ (the recovery rate).
-Due to its simplicity, the solution of the SIR model and its properties can be analyzed analytically [harko2014exact] and it often serves as an initial model for the modeling of unknown diseases [daley1999epidemic].
+Due to its simplicity, the solution of the deterministic SIR model and its properties can be analysed analytically [Harko et al., 2014] and it often serves as an initial model for the modeling of unknown diseases [Daley and Gani, 1999].
 For consistency across models, in the following we compare recovery in terms of the mean infectious period, $\gamma^{-1}$, rather than $\gamma$.
@@
 2.3 Two-Variant SEIR Model
-Many extensions of the SIR model have been developed to account for the heterogeneity of various infectious diseases [keeling2008], for example the Susceptible-Exposed-Infectious-Recovered model, in which the Exposed compartment covers the latent period between being infected and becoming infectious.
-Recently, an extension to the ODE-based SEIR model was introduced to capture the early COVID-19 dynamics in Ethiopia [ethiopian_data].
-In addition to a classical SEIR model, this model distinguishes between two virus variants, namely the wild-type and a novel variant, leading to a total of 10 compartments (Figure 1(b)).
-The novel variant is more potent in the sense that it shows a prolonged infectious period and that immunity from the wild-type variant does not protect against reinfection with the novel variant.
+Many extensions of the SIR model have been developed to account for the heterogeneity of various infectious diseases [Keeling and Rohani, 2008], for example the Susceptible-Exposed-Infectious-Recovered model, in which the Exposed compartment covers the latent period between being infected and becoming infectious.
+Recently, an extension to the ODE-based SEIR model was introduced to capture the early COVID-19 dynamics in Ethiopia [Gudina et al., 2021].
+In addition to a classical SEIR model, this model distinguishes between two virus variants, namely the wild-type and a novel variant, leading to a total of 10 compartments (Figure 1(c)).
+The novel variant is assumed to have a 35% longer infectious period, following the original formulation of the two‑variant SEIR model in Gudina et al.
+[2021].
+We adopt this fixed factor to remain consistent with the published model on which our comparison is based.
+Additionally, immunity from the wild-type variant does not protect against reinfection with the novel variant.
 However, an infection with the novel variant type confers immunity against subsequent infections with both the wild-type and the novel variant.
 The dynamics of the compartment sizes are described by initialization, variant entering and transitions between the compartments (Table 3).
-The model exhibits non-identifiability issues in the sense that different parameter combinations, lying along certain curves, can generate nearly indistinguishable epidemic trajectories (see Supplementary Information A.3).
-Therefore, we also consider a reparametrized version of the model with a reduced number of model parameters to be estimated.
-We fix $\kappa^{-1}$ and use the following redefined parameters
-$r_{0}=\beta\cdot\gamma^{-1},\quad e_{0}=\kappa^{-1}+\gamma^{-1},\quad s_{0}=s\cdot\gamma^{-1},$
+The model exhibits non-identifiability issues in the sense that different parameter combinations, lying along certain curves, can generate nearly indistinguishable epidemic trajectories (see Supplementary Information S1.3.6.2).
+Therefore, in addition to this model, we also consider a reparametrized version of the model with a reduced number of model parameters to be estimated.
+For this, we fix $\kappa^{-1}=5$, since $\kappa$ seems to participate in the non-identifiable parameter combinations and use the following redefined parameters
+$r_{0}=\beta\cdot\gamma^{-1},\quad e_{0}=\kappa^{-1}+\gamma^{-1},\quad s_{0}=s\cdot\gamma^{-1},$ (1)
 while keeping $t_{\mathrm{var}}$ and $I_{0}$ as well as the observational and noise models unchanged.
@@
 The course of an epidemic is usually only partially observed and data is often only available in aggregated form on population level.
-For both models introduced above, we consider an observation model reflecting two often measured quantities to monitor infectious diseases.
+For all models introduced above, we consider an observation model reflecting one to two often measured quantities to monitor infectious diseases.
+For the SIS model, the observable is
+$y(t)=\left(\begin{array}[]{c}\frac{\mathrm{I}_{\text{}}(t)}{\mathrm{N}_{\text{}}(t)}\end{array}\right).$ (3)
 For the SIR model, the observables are
@@
 where the first entry is the ratio of currently infected people describing the count data from testing, and the second entry describes seroprevalence, the ratio of individuals with antibodies present due to a prior infection.
-Because the observation process may have several limitations, an additional scaling parameter $s$ is used to capture that only a fraction of cases is reported [ethiopian_data].
+Because the observation process may have several limitations, an additional scaling parameter $s$ is used to capture that only a fraction of cases is reported [Gudina et al., 2021].
 To reflect measurement uncertainty, synthetic realizations of the observables are perturbed with noise. In this study, we leverage additive Gaussian noise. That means we have
@@
 where we truncate $\overline{y}_{1}(t_{i})$ and $\overline{y}_{2}(t_{i})$ at a lower bound of 0 to avoid unrealistic negative observations.
-As in the dataset given by [ethiopian_data], the standard deviations $\sigma_{i,1},\sigma_{i,2}$ were computed directly using a normal approximation to the binomially distributed testing count data.
+As in the dataset given by Gudina et al.
+[2021], the standard deviations $\sigma_{i,1},\sigma_{i,2}$ were computed directly using a normal approximation to the binomially distributed testing count data.
 Alternative probability distributions may be more appropriate in other contexts.
-Further details on the computations of measurement noise are available in Supplementary Information A.2.
-From each of the two stochastic compartmental models we generated synthetical data emulating the time course of a disease outbreak.
+From each of the stochastic compartmental models we generated synthetic data emulating the time course of a disease outbreak.
 We generated dense datasets with observations every seven or ten days.
+We used two types of parameter regimes:
+  1. 1.
+    hand‑selected values chosen to represent plausible epidemiological regimes, and
+  2. 2.
+    parameters sampled directly from the corresponding priors as an unbiased baseline
+Prior draws that generated trajectories with collapsed variance were discarded and resampled to maintain numerical stability of the particle filter.
 Additionally, to create realistic, heterogeneous observation schedules, we randomly subsampled the dense time series into sparse datasets reducing the number of time points from 57 to fewer than 15 per observable.
 We also introduced partial missingness in the sparse datasets, where only one of the two observables was available at certain time points.
-Details of the data generation and different datasets used are outlined in Supplementary Information A.6.
+Details on the computation of measurement noise and data generation are provided in the Models section of the Supplementary Information S1 (Section S1.3).
 3 Bayesian Inference for Stochastic Compartmental Models
-To infer the unknown model parameters $\mathbf{\theta}$, which best explain the data $\mathcal{D}^{*}=\{\overline{y}_{t_{i}}\mid i\in\{1,...,n\}\}$}, we apply Bayesian inference techniques.
+To infer the unknown model parameters $\mathbf{\theta}$ given data $\mathcal{D}^{*}=\{\overline{y}_{t_{i}}\mid i\in\{1,...,n\}\}$}, we apply Bayesian inference techniques.
 For this, we denote the space of unobserved states of the diffusion process by $\mathcal{X}$.
@@
 The posterior represents the updated belief about the parameters after taking the observed data into account via the likelihood, effectively integrating both previous knowledge and the new information provided by the data.
-While the full posterior distribution is typically of primary interest, we additionally consider the mode (called maximum a-posteriori) of the posterior distribution as the best point estimate [bernardo1994bayesian].
+While the full posterior distribution is typically of primary interest, we additionally consider the mode (called maximum a-posteriori) of the posterior distribution as the best point estimate [Bernardo et al., 1994].
 One challenge in applying Bayesian inference to stochastic compartmental models is dealing with an intractable posterior distribution, because evaluating the likelihood
@@
 requires marginalizing over stochastic realizations of the unobserved states at measurement time points.
-In most cases, evaluating the density of hidden states analytically is impossible [Doucet_SMC], which prohibits evaluating the likelihood.
+In most cases, evaluating the density of hidden states analytically is impossible [Doucet et al., 2001], which prohibits evaluating the likelihood.
 Even in the case of a tractable likelihood, normalization of the posterior is done via the marginal likelihood, which can remain intractable.
-High‑dimensional parameter spaces further make discretization or numerical integration prohibitively expensive [wilkinson2010, hasenauer2013thesis].
+High‑dimensional parameter spaces further make discretization or numerical integration prohibitively expensive [Wilkinson, 2011; Hasenauer, 2013].
 Approximation techniques that estimate the posterior distribution without directly computing the complicated likelihood function enable robust parameter estimation, even in settings where traditional methods may fail due to mathematical or computational complexity.
-In this section, we will introduce two approaches (particle filters [golightly_particle_2011] and conditional normalizing flows [cranmer2020frontier]) that circumvent the computation of the likelihood function.
+In this section, we will introduce two approaches (particle filters (PF) [Golightly and Wilkinson, 2011] and one particular class of conditional density estimators, conditional normalizing flows (CNF) [Papamakarios and Murray, 2016; Cranmer et al., 2020]) that circumvent the computation of the likelihood function.
 PF approximate the likelihood with an unbiased estimator obtained by sequential importance sampling, and conditional normalizing flows learn a series of invertible mappings to transform a Gaussian density into the desired density.
 3.1 Particle Markov Chain Monte Carlo Methods
-Particle Markov chain Monte Carlo (PMCMC) methods are based on the idea of using an unbiased estimator of the likelihood function, obtained by a particle filter, inside a Markov chain Monte Carlo (MCMC) method [golightly_particle_2011].
+Particle Markov chain Monte Carlo (PMCMC) methods are based on the idea of using an unbiased estimator of the likelihood function, obtained by a particle filter, inside a Markov chain Monte Carlo (MCMC) method [Golightly and Wilkinson, 2011].
 Particle filters are sequential Monte Carlo (SMC) techniques based on importance sampling with resampling used for handling non-linear and non-Gaussian time-series models.
@@
 And third, they are resampled using auxiliary variables $A_{t_{j}}^{1:N}$ to discard low-weight particles.
+In practice, the resampling is not done in every iteration, but only if the variability of the weights becomes too large, as measured by the effective sample size (ESS) (for more details see Supplementary Information S1.1.2.2).
 The resampled particles are then again propagated forward to the next time point.
 Hence, for each time $t_{i}$, the particle filter constructs a system of $N$ weighted particles $\{X_{t_{i}}^{(k)},\,w_{t_{i}}^{(k)}\}_{k=1}^{N}$ that approximates the filtering distribution $p(X_{t_{i}}\,\mid\,y_{t_{0}:t_{i}},\theta)$.
-Mathematically, this is closely linked to the framework of Feynman-Kac models [moral2004feynman, andrieu_2010, chopin_2020].
+Mathematically, this is closely linked to the framework of Feynman-Kac models [Moral, 2004; Andrieu et al., 2010; Chopin and Papaspiliopoulos, 2020].
 For the task of estimating the likelihood in stochastic compartmental models, we use a bootstrap filter Algorithm 1.
@@
 7:  if $\mathrm{ESS(W}_{t_{i-1}}^{1:N})<\mathrm{ESS_{min}}$ then
-8:   Sample $A^{1:N}_{t_{i}}\sim\text{resample}(W_{t_{i}}^{1:N})$    $\triangleright$ See Supplementary Information A.4.2
+8:   Sample $A^{1:N}_{t_{i}}\sim\text{resample}(W_{t_{i-1}}^{1:N})$    $\triangleright$ See Supplementary Information S1.1.2.2
 9:   $\hat{w}_{t_{i-1}}^{(n)}\leftarrow 1$
@@
 17:end for
-Although the name filtering refers to the task of estimating the latent state $X_{t_{i}}$ online given data $y_{t_{0}:t_{i}}$, a PF yields, as a by-product, a scheme to compute an unbiased estimator of the likelihood contributions ${\{p(y_{t_{0}:t_{i}}\,\mid\,\theta)\}}_{i\geq 0}$, crucial for the task of parameter inference [chopin_2020].
+Although the name filtering refers to the task of estimating the latent state $X_{t_{i}}$ online given data $y_{t_{0}:t_{i}}$, a PF yields, as a by-product, a scheme to compute an unbiased estimator of the likelihood contributions ${\{p(y_{t_{0}:t_{i}}\,\mid\,\theta)\}}_{i\geq 0}$, crucial for the task of parameter inference [Chopin and Papaspiliopoulos, 2020].
 This estimator is obtained by computing the normalizing constants
@@
 The estimate $L_{t_{M}}^{N}(\theta)$ can then be used inside a pseudo-marginal MCMC algorithm to sample from the posterior $p(\theta\mid y_{t_{0}:t_{M}})$.
-In this work, we use the likelihood estimate obtained by the bootstrap filter within the acceptance probability of a Metropolis-Hasting (MH) algorithm as the outer MCMC scheme, making it a member of the class of Pseudo-Marginal Metropolis-Hasting methods (PMMH) [andrieu_roberts, andrieu_2010, chopin_2020].
-Using an unbiased estimator $\hat{p}(y|z,\theta)$ of the $p(y|\theta)$ that is obtained by the use of auxiliary random variables $z\sim r(\cdot)$ within the acceptance of a MH algorithm then corresponds to simulating a Markov chain $\{\theta^{i},z^{i}\}_{i\geq 0}$ targeting the joint density
+In this work, we use the likelihood estimate obtained by the bootstrap filter within the acceptance probability of a Metropolis-Hasting (MH) algorithm as the outer MCMC scheme, making it a member of the class of Pseudo-Marginal Metropolis-Hasting methods (PMMH) [Andrieu and Roberts, 2009; Andrieu et al., 2010; Chopin and Papaspiliopoulos, 2020].
+The full PMMH-algorithm is provided in the Supplementary Information S1.1.1.
+Using an unbiased estimator $\hat{p}(y|z,\theta)$ of the likelihood $p(y|\theta)$ that is obtained by the use of auxiliary random variables $z\sim r(\cdot)$ within the acceptance of a MH algorithm then corresponds to simulating a Markov chain $\{\theta^{i},z^{i}\}_{i\geq 0}$ targeting the joint density
 $\bar{p}(\theta,z)\propto p(\theta)r(z)\hat{p}(y\mid\theta,z).$
 For the bootstrap filter, we can, for instance, think of the $A_{t_{j}}^{1:N}$ used in the resampling as being the auxiliary random variables.
-The joint density admits the correct posterior density of the model $p(\theta\mid y)$ as a marginal and therefore the algorithm samples the parameter vectors $\theta$ from the true posterior [andrieu_roberts, andrieu_2010, doucet_2015].
+The joint density admits the correct posterior density of the model $p(\theta\mid y)$ as a marginal and therefore the algorithm samples the parameter vectors $\theta$ from the true posterior [Andrieu and Roberts, 2009; Andrieu et al., 2010; Doucet et al., 2015].
 With respect to an extended distribution, PMMH algorithms are standard MCMC samplers.
 One can use any unbiased estimator within any MCMC algorithm to produce a PMCMC algorithm.
-Different filtering and sampling algorithms impose different practical and implementational challenges and we focus here on the most general applicable version [doucet_2015, chopin_2020].
+Different filtering and sampling algorithms impose different practical and implementational challenges and we focus here on the most general applicable version [Doucet et al., 2015; Chopin and Papaspiliopoulos, 2020].
 In the subsequent sections PF method always refers to a PMMH algorithm consisting of a bootstrap filter inside an adaptive Metropolis-Hastings algorithm.
 3.2 Neural Posterior Estimation And Conditional Normalizing Flows
-Neural posterior estimation (NPE) is a simulation-based inference (SBI) method which frames posterior construction as a (conditional) density-estimation problem.
+Neural posterior estimation (NPE) is a simulation-based inference (SBI) method which frames posterior construction as a (conditional) density-estimation problem [Papamakarios and Murray, 2016].
 Given a simulator that produces samples $(\mathcal{D},\theta)$ by first drawing $\theta\sim p(\theta)$ and then $\mathcal{D}\sim p(\mathcal{D}\mid\theta)$, NPE fits a conditional neural network $p_{\phi}(\theta\mid\mathcal{D})$ to approximate the true posterior $p(\theta\mid\mathcal{D})$.
-Typically the data $\mathcal{D}$ are preprocessed by a (learned) summary statistic [radev2020bayesflow], which serves as the input to the conditional density estimator.
+Typically the data $\mathcal{D}$ are preprocessed by a (learned) summary statistic [Radev et al., 2020], which serves as the input to the conditional density estimator.
 For clarity, this preprocessing step is omitted from the notation below.
-In this work, we focus on Conditional Normalizing Flows (CNFs) [rezende2015variational].
+In this work, we focus on Conditional Normalizing Flows (CNFs) [Rezende and Mohamed, 2015].
+Normalizing flows have since been substantially developed and systematized [Papamakarios et al., 2021; Boelts et al., 2025; Radev et al., 2023].
 In general, normalizing flows transform a simple base distribution (e.g., a standard Gaussian) into a complex target distribution via a sequence of invertible, differentiable mappings.
@@
 CNFs extend this idea by allowing the transformation to additionally depend on datasets $\mathcal{D}$, generated by first drawing $\theta\sim p(\theta)$ and then $\mathcal{D}\sim p(\mathcal{D}\mid\theta)$.
-Rather than learning a single transformation $f$, CNFs learn a family of conditional flows $\{f_{\mathcal{D}}\}$, with each $f_{\mathcal{D}}$ mapping the base distribution $q$ to the posterior $p(\theta\mid\mathcal{D})$ [winkler2019conditional].
-This conditioning enables CNFs to capture how the posterior changes with different observations, making them particularly well-suited for amortized inference tasks, where a model learns to perform inference efficiently across multiple datasets rather than solving each case from scratch [kobyzev2020normalizing, rezende2015variational].
-Instead of hand-designing each $f_{\mathcal{D}}$, we parameterize the family using a neural network $f_{\phi}(\cdot,\,\mathcal{D})$ whose parameters $\phi$ are shared across all $\mathcal{D}$, but whose behavior is modulated by $\mathcal{D}$ itself [ardizzone2019guided].
-We train the neural network by minimizing the Kullback-Leibler divergence between the true and learned posteriors for all $\mathcal{D}$ in the training distribution [weilbach2020structured, siahkoohi2020conditional, radev2023bayesflow].
+Rather than learning a single transformation $f$, CNFs learn a family of conditional flows $\{f_{\mathcal{D}}\}$, with each $f_{\mathcal{D}}$ mapping the base distribution $q$ to the posterior $p(\theta\mid\mathcal{D})$ [Winkler et al., 2019].
+This conditioning enables CNFs to capture how the posterior changes with different observations, making them particularly well-suited for amortized inference tasks, where a model learns to perform inference efficiently across multiple datasets rather than solving each case from scratch [Kobyzev et al., 2020; Rezende and Mohamed, 2015].
+Instead of hand-designing each $f_{\mathcal{D}}$, we parameterize the family using a neural network $f_{\phi}(\cdot,\,\mathcal{D})$ whose parameters $\phi$ are shared across all $\mathcal{D}$, but whose behavior is modulated by $\mathcal{D}$ itself [Ardizzone et al., 2019].
+We train the neural network by minimizing the Kullback-Leibler divergence between the true and learned posteriors for all $\mathcal{D}$ in the training distribution [Weilbach et al., 2020; Siahkoohi et al., 2020; Radev et al., 2023].
 This means, our optimization objective is finding $\hat{\phi}$
@@
 Algorithm 2 Amortized training and inference for conditional normalizing flows
-1:Data source $S$ (pre‑simulated dataset or simulator callable $\sim p(\theta)p(\mathcal{D}\mid\theta)$), base density $q(z)$, conditional flow $f_{\phi}(\cdot;\mathcal{D})$, optimizer Opt, learning rate $\eta$, batch size NN, iterations $T$, inference data $\mathcal{D}^{*}$, inference sample count $M$
-2:Trained parameters $\phi$; sampler for approximate posterior $p_{\phi}(\theta\mid\mathcal{D}_{*})$
-3:for $t=1$ to $T$ do
-4:  $\{(\theta_{i},\mathcal{D}_{i})\}_{i=1}^{N}\leftarrow\text{sample from $S$ (dataset or simulator)}$
-5:  for $i=1$ to NN do
-6:   $z_{i}\leftarrow f_{\phi}(\theta_{i};\mathcal{D}_{i})$
-7:   $\log p_{\phi}(\theta_{i}\mid\mathcal{D}_{i})\leftarrow\log q(z_{i})+\log\Big|\det\Big(\frac{\partial f_{\phi}(\theta_{i};\mathcal{D}_{i})}{\partial\theta}\Big)\Big|$
-8:  end for
-9:  $\mathcal{L}\leftarrow-\frac{1}{N}\sum_{i=1}^{N}\log p_{\phi}(\theta_{i}\mid\mathcal{D}_{i})$
-10:  $\phi\leftarrow\text{Opt}(\phi,\nabla_{\phi}\mathcal{L},\eta)$
-11:end for
-12:
+1:Data source $S$ (pre‑simulated dataset or simulator callable $\sim p(\theta)p(\mathcal{D}\mid\theta)$), base density $q(z)$, conditional flow $f_{\phi}(\cdot;\mathcal{D})$, optimizer Opt, learning rate $\eta$, batch size NN, inference data $\mathcal{D}^{*}$, inference sample count $M$
+2:Trained parameters $\phi$; sampler for approximate posterior $p_{\phi}(\theta\mid\mathcal{D}_{*})$ repeat
 […]
-14:  for $m=1$ to $M$ do
-15:   Sample $z_{m}\sim q(z)$
-16:   $\theta_{m}\leftarrow f_{\phi}^{-1}(z_{m};\mathcal{D}_{*})$
-17:  end for
-18:  return $\{\theta_{m}\}_{m=1}^{M}$
-19:end procedure
+4:for $i=1$ to NN do
+5:  $z_{i}\leftarrow f_{\phi}(\theta_{i};\mathcal{D}_{i})$
+6:  $\log p_{\phi}(\theta_{i}\mid\mathcal{D}_{i})\leftarrow\log q(z_{i})+\log\Big|\det\Big(\frac{\partial f_{\phi}(\theta_{i};\mathcal{D}_{i})}{\partial\theta}\Big)\Big|$
+7:end for
+[…]
+[…]
+10:
+[…]
+12:  for $m=1$ to $M$ do
+13:   Sample $z_{m}\sim q(z)$
+14:   $\theta_{m}\leftarrow f_{\phi}^{-1}(z_{m};\mathcal{D}_{*})$
+15:  end for
+16:  return $\{\theta_{m}\}_{m=1}^{M}$
+17:end procedure
 This amortization not only enables fast inference once training is complete, but also facilitates simulation‑based calibration (SBC) diagnostics by allowing posterior approximations to be generated efficiently across many simulated datasets.
-At the same time, performance of CNFs is tied to the support of the training distribution: the model can interpolate to new values if they are close to the training support, but performance degrades when extrapolating far beyond it [winkler2019conditional, talts2018validating].
+At the same time, performance of CNFs is tied to the support of the training distribution: the model can interpolate to new values if they are close to the training support, but performance degrades when extrapolating far beyond it [Winkler et al., 2019; Talts et al., 2018].
+In practice, this amortized structure is precisely what makes SBC feasible at scale, since once trained, the CNF can be evaluated repeatedly on equally structured datasets without retraining.
 4 Implementation Of The Methods And Their Comparison
-In order to render the simulation based approaches computationally feasible, model simulations were conducted using the efficiency of the Julia programming language and the flexible SDE modeling capabilities of SciML and ModelingToolkit.jl [julia_stochdiffeq, ma2021modelingtoolkit].
+In order to render the simulation based approaches computationally feasible, model simulations were conducted using the efficiency of the Julia programming language and the flexible SDE modeling capabilities of SciML and ModelingToolkit.jl [Rackauckas and Nie, 2017; Ma et al., 2021].
 To approximate the solution of the SDE, we implemented an Euler-Maruyama scheme that constrains trajectories to stay non-negative, effectively clipping potential negative parts and setting them to zero.
 For PF, existing software packages were unsuitable as they either lacked Julia support or restricted model specification to simulator-only interfaces.
-We therefore implemented a custom particle filter in Julia, drawing inspiration from the Python package Particles [chopin_2020].
-As an outer sampling scheme, we employed an adaptive Metropolis–Hastings algorithm [adaptive_mh], using the implementation in pyPESTO [pypesto], to account for potential multimodality in the posterior distribution.
+We therefore implemented a custom particle filter in Julia, drawing inspiration from the Python package Particles [Chopin and Papaspiliopoulos, 2020].
+As an outer sampling scheme, we employed an adaptive Metropolis–Hastings algorithm [Haario et al., 2001], using the implementation in pyPESTO [Schälte et al., 2023], to account for potential multimodality in the posterior distribution.
 Missing data were handled by setting the likelihood contribution of the corresponding observation to zero.
-To mitigate particle degeneracy, we set the number of particles to 200, ensuring that the variance of the likelihood estimate remained below one for parameters around the posterior mode (Supplementary Figure S3).
+To mitigate particle degeneracy, we set the number of particles to 200, ensuring that the variance of the likelihood estimate remained below one for parameters around the posterior mode (Supplementary Fig.
+S1.1).
 For the outer MH algorithm, we ran 4 Markov chains with a length of 50,000 samples each.
 We then discarded a burn-in period of the first 25,000 samples per chain.
-After that, convergence was evaluated using rank‑normalized Gelman–Rubin statistics ($\hat{R}$) and the ESS with autocorrelations up to a maximal lag of 250, both from MCMCDiagnosticTools.jl [turing_jl].
-Values of $\hat{R}$ close to 1.0 indicate good convergence and well‑mixed chains [vehtari2021rank].
+After that, convergence was evaluated using rank‑normalized Gelman–Rubin statistics ($\hat{R}$) and the ESS with autocorrelations up to a maximal lag of 250, both from MCMCDiagnosticTools.jl [Fjelde et al., 2025].
+Values of $\hat{R}$ close to 1.0 indicate good convergence and well‑mixed chains [Vehtari et al., 2021].
 The ESS assesses and diagnoses poor convergence in the bulk of the distribution due to trends or different locations of the chains.
-For a given estimand, it is recommended that ESS should exceed $100\cdot\text{nchains}$ [vehtari2021rank].
+For a given estimand, it is recommended that ESS should exceed $100\cdot\text{nchains}$ [Vehtari et al., 2021].
 Together, $\hat{R}$ and ESS provide complementary diagnostics of chain convergence and posterior exploration quality.
-For CNF, posterior approximations were obtained via neural posterior estimation with learned summary statistics [radev2020bayesflow], using the BayesFlow software (v1) for amortized Bayesian workflows [radev2023bayesflow].
+For CNF, posterior approximations were obtained via neural posterior estimation with learned summary statistics [Radev et al., 2020], using the BayesFlow software (v1) for amortized Bayesian workflows [Radev et al., 2023].
 We trained an independent neural network for each experiment rather than a single globally amortized model.
-To ensure comparability, all experiments used identical hyperparameters and simulation budgets, determined through a preliminary hyperparameter search.
+This preserves amortization over latent trajectories and parameter values within each experiment, while avoiding the need to amortize over continuously varying noise levels across experiments, a setting in which global amortization would be inefficient and unnecessary.
+The resulting per‑experiment amortization still enables efficient SBC diagnostics and fast posterior evaluation once training is complete.
+Rather than fine‑tuning each method for every individual experiment, which would confound performance differences with tuning effort, we selected CNF hyperparameters through a principled search procedure that we repeated to assess stability.
+We verified that training was well‑behaved by inspecting training and validation loss curves, which consistently stabilized well before 100 epochs.
 The dimension of the learned summary statistics was set to $2n_{\theta}+2$, where $n_{\theta}$ denotes the number of model parameters.
@@
 Models were trained for 100 epochs with a batch size of 32.
-Summary statistics were extracted using a SequenceNetwork [radev2021outbreakflow] with 64 LSTM units, and the conditional normalizing flow was modeled using an InvertibleNetwork with spline coupling and eight coupling layers, both implemented within the BayesFlow framework.
-Detailed configurations are provided in Supplementary Information A.5.
+Summary statistics were extracted using a SequenceNetwork [Radev et al., 2021] with 64 LSTM units, and the conditional normalizing flow was modeled using an InvertibleNetwork with spline coupling and eight coupling layers, both implemented within the BayesFlow framework.
+Detailed configurations are provided in Supplementary Information S1.1.3.
 Handling non‑equidistant time points and missing data required a specialized strategy.
-We adopted the approach of [wang2024missing], which augments the input with binary indicators marking the presence or absence of each entry.
-This encoding allows the neural network to account for missing values during inference.
+We adopted the approach of Wang et al.
+[2024], which augments the input with binary indicators marking the presence or absence of each entry.
+This encoding allows the neural network to account for missing values during inference and provides a consistent representation when combining observation functions recorded at different time points.
+In our setup, only the dimensional consistency of the two observation functions is strictly required; however, we included the binary mask deliberately to facilitate rapid amortization for practitioners who may wish to extend our framework.
+For the present experiments, this amortization capability was not used: for sparse datasets, the neural networks were trained solely on the specific missingness pattern present in the simulated data to ensure a fair comparison between PF and CNF and to avoid confounding method differences with amortization artifacts.
 For experiments with dense, equidistant data, this additional encoding was not applied.
-For sparse datasets, the resulting neural networks were trained on the specific missingness pattern present in the simulated data; no amortization across different missingness structures was performed.
 Empirical cumulative distribution function (ECDF) plots were obtained to assess calibration, and parameter recovery diagnostics were used to evaluate the correspondence between inferred and true parameter values.
@@
 For the comparison of both methods we used visual and quantitative measures.
-The visual comparison using violin and density plots of the marginal posteriors allows comparison of the shape and variability of the posterior approximations.
+The visual comparison using density plots of the marginal posteriors allows comparison of the shape and variability of the posterior approximations.
 This is complemented by a visual comparison of the model fit.
@@
 These are independent posterior predictive samples, obtained by evaluating $p(y\mid\theta)$ for draws $\theta\sim p(\theta\mid\mathcal{D}^{*})$.
-For a quantitative measure of inference accuracy, we compute the maximum a-posteriori (MAP) estimate in the parameter space via Gaussian kernel density estimation using Silverman’s rule [silverman1998density] as well as the root-mean squared error (RMSE).
-Specifically, the RMSE was calculated for trajectories simulated from each of the 10,000 posterior samples and then averaged, providing an estimate of the expected RMSE under the posterior distribution.
-For synthetic experiments, the RMSE was also computed with respect to the true parameters used to generate the data.
+As an additional qualitative indicator, we report the maximum a‑posteriori (MAP) estimate computed via Gaussian kernel density estimation using Silverman’s rule [Silverman, 1998], which provides a consistent definition across methods.
+For a quantitative measure of inference accuracy, we first assess marginal coverage by reporting how frequently the true generative parameters fall within the 95% marginal posterior intervals across datasets.
+We also quantify posterior differences using the 1‑Wasserstein distance in log‑parameter space.
+All Wasserstein distances are obtained by repeatedly subsampling 1,000 draws per posterior, and we report in‑distribution values to characterize the Monte Carlo error of this procedure.
+Heuristically, the 1‑Wasserstein distance in log‑parameter space measures the average multiplicative shift required to transform one posterior into another, providing an interpretable notion of scale‑adjusted discrepancy.
+Concretely, we compute these distances in three settings:
+  • Baseline comparisons: when a baseline is available, we measure the Wasserstein distance between each method and the baseline and compare it to the corresponding in‑distribution reference.
+  • Model‑parameterization comparison: for the full and reparametrized SEIR2v models, we evaluate how the Wasserstein distance changes when keeping the dataset fixed, again reporting the in‑distribution values as a baseline.
+  • Within‑method variability: for one representative parameter set, we assess the variability of each method by computing Wasserstein distances between multiple PF runs or multiple CNF trainings.
+Finally, we assess predictive accuracy via the energy score of the posterior predictive distributions.
+For synthetic experiments, the predictive energy score was also computed with respect to the true parameters used to generate the data.
 For the real-data experiment, true parameters are not available.
-Instead, we report RMSE relative to published parameter estimates from [ethiopian_data], which serve as a reference baseline.
+Instead, we report the energy score relative to published parameter estimates from [Gudina et al., 2021], which serve as a reference baseline.
 All experiments were conducted on ten cores of an AMD EPYC 7f72 3.2 GHz processor with a total of 200 GB of RAM made available to the process.
@@
 5 Results
-We examined inference with CNF and PF for stochastic epidemic models through a sequence of in silico experiments on synthetic data, following the workflow in Figure 1(c).
-We begin with the standard SIR model (Section 2.1) using a synthetic prior to validate both methods.
-We then turn to the two-variant SEIR model (Section 2.2) with literature-informed priors [mcaloon2020incubation, fang2020comparisons], designed to probe the influence of parameter non-identifiability (Section A.3).
+We examined inference with CNF and PF for stochastic epidemic models through a sequence of in-silico experiments on synthetic data, following the workflow in Figure 1(d).
+To assess the performance of both inference methods against a baseline posterior distribution, we use the standard SIS (Section 2.1) and SIR (Section 2.2) model with a synthetic prior.
+The baseline posterior was computed on a finely discretized version of the SDE using a state-of-the-art Hamiltonian Monte Carlo sampler [Hoffman et al., 2014](Details in Supplementary Information S1.1.4).
+We then turn to the complex two-variant SEIR model (Section 2.3) with literature-informed priors [McAloon et al., 2020; Fang et al., 2020], designed to probe the influence of parameter non-identifiability.
 Within the two-variant SEIR setting, we additionally study the effects of data sparsity and timing mismatch and investigate a reparametrization aimed at reducing non-identifiability.
 For completeness, Table 4 summarizes all settings and variations.
-Finally, we apply both approaches to longitudinal data from Ethiopia [ethiopian_data] to assess performance under realistic conditions (Section 5.4).
-Additional configurations and results are provided in the Supplement (Supplementary Tables S4S18).
+Finally, we apply both approaches to longitudinal data from Ethiopia [Gudina et al., 2021] to assess performance under realistic conditions.
+Evaluation metrics are described in Supplementary Information S1.2 and synthetic dataset generation together with model setups are described in detail in Supplementary Information S1.3.
 Table 4: Experiments for method comparison. Overview over the subsections and the corresponding model and data setup used.
@@
 Reparametrization of the two-variant SEIR model with synthetic dense datasets.
-[…]
+$5Subsection~\ref${subsec:results_ethiopia}
 Two-variant SEIR model with dataset from a longitudinal cohort study
-in Ethiopia [ethiopian_data].
+in Ethiopia [Gudina et al., 2021].
+5.1 CNF and PF Agree With A Reference Posterior For The SIS Model
+We evaluated convergence and posterior estimates for the SIS model using CNF, PF, and a Hamiltonian Monte Carlo (HMC) baseline applied to a finely discretized version of the SDE.
+Across 10 synthetic dense datasets, all methods achieved excellent marginal coverage and produced highly consistent Wasserstein distances.
+Two datasets (sis‑2 and sis‑4) exhibited unusually large distances and were excluded from the aggregate analysis but even for these outliers, visual posterior differences remained minor.
+After exclusion, the in‑distribution Monte Carlo baseline (HMC–HMC) was small, with values ranging from 0.056 to 0.119.
+Between‑method distances were only moderately larger: CNF–PF ranged from 0.063 to 0.156, CNF–HMC ranged from 0.065 to 0.176 and PF–HMC ranged from 0.058 to 0.124.
+Interpreted on the log‑parameter scale, these values correspond to average multiplicative shifts of roughly $6\%-13\%$ for the baseline and $6\%-17\%$, $7\%-19\%$, and $6\%-13\%$ for the between‑method comparisons, respectively.
+These results indicate that systematic differences between inference algorithms are small, consistent with the strong visual overlap in marginal and joint posteriors (Figure 2A).
+Posterior predictive simulations from all three methods reproduced the observed epidemic trajectories with comparable accuracy (Figure 2B), as evidenced by consistent energy scores computed from the posterior predictive distributions of the methods and the ground truth (Table 5).
+In terms of sampling behavior, PF showed excellent convergence, with ESS values above 2000 and $\mathbf{\hat{R}}$ near 1.
+CNF showed good calibration in simulation‑based calibration diagnostics, with near‑uniform rank histograms and accurate parameter recovery.
+Overall, both PF and CNF matched the predictive and inferential performance of the HMC reference.
+Full diagnostics, including ESS, $\mathbf{\hat{R}}$, Wasserstein distance results, calibration plots, and parameter-recovery analyses are provided in Supplementary Information S2.
+Figure 2: Evaluation of the posterior approximations for the SIS model
+A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade.
+B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
+Table 5: Predictive Energy Score for the SIS model.
+sis-1 sis-2 sis-3 sis-4 sis-5
+True $0.034$ $0.051$ $0.028$ $0.019$ $0.024$
+CNF $0.035$ $0.046$ $0.025$ $0.018$ $0.024$
+PF $0.034$ $0.046$ $0.025$ $0.018$ $0.023$
+HMC $0.035$ $0.046$ $0.025$ $0.018$ $0.023$
+sis-6 sis-7 sis-8 sis-9 sis-10
+True $0.041$ $0.025$ $0.045$ $0.042$ $0.041$
+CNF $0.040$ $0.027$ $0.044$ $0.035$ $0.040$
+PF $0.037$ $0.022$ $0.042$ $0.031$ $0.040$
+HMC $0.037$ $0.022$ $0.042$ $0.031$ $0.039$
 5.2 CNF And PF Provide Accurate Estimates For The SIR Model.
-We compared convergence and posterior estimates obtained with CNF and PF on the SIR model, using two distinct simulated dense datasets (details in Supplementary Information A.6).
-The prior distributions and parameter values are summarized in Supplementary Tables S3 and S4.
-Both CNF and PF yielded highly consistent posterior estimates for the SIR model.
-The marginal and joint posterior distributions overlapped almost perfectly (Figure 2A), and posterior predictive checks showed that epidemic trajectories generated from either method fit the observed data equally well (Figure 2B).
-MAP estimates from CNF and PF were nearly identical and close to the true parameters, with RMSEs for infection counts and seroprevalence matching those obtained from simulations with the ground truth (Table 4).
+We compared convergence and posterior estimates obtained with CNF and PF on the SIR model, using twelve distinct synthetic dense datasets.
+As a reference, we additionally computed a reference posterior using Hamiltonian Monte Carlo (HMC) applied to a finely discretized version of the underlying SDE.
+The in‑distribution Monte Carlo baseline (HMC–HMC) ranged from 0.034 to 0.096.
+Between‑method distances were modestly larger: CNF–PF ranged from 0.063 to 0.165, CNF–HMC from 0.076 to 0.183, and PF–HMC from 0.048 to 0.125.
+Interpreted on the log‑parameter scale, these values correspond to average multiplicative shifts of approximately 3–10% for the baseline and 6–18%, 8–20%, and 5–13% for the between‑method comparisons, respectively.
+Posterior predictive checks showed that epidemic trajectories generated from all three methods fit the observed data equally well (Figure 3B), as evidenced by the close agreement in predictive energy scores (Table 6).
+Across all datasets, PF and CNF achieved 83.33% marginal coverage and HMC achieved 87.5%, and the three posterior approximations showed high agreement overall.
+Marginal and joint posterior distributions overlapped closely (Figure 3A) with CNF exhibiting a small systematic offset to HMC and PF.This offset is consistent with the mild posterior miscalibration discussed below.
+The Wasserstein distances reflect this pattern.
+The in‑distribution Monte Carlo baseline (HMC–HMC) ranged from 0.034 to 0.096.
+Between‑method distances were modestly larger: CNF–PF ranged from 0.063 to 0.165, CNF–HMC from 0.076 to 0.183 and PF–HMC ranged from 0.048 to 0.125.
+Interpreted on the log‑parameter scale, these values correspond to average multiplicative shifts of approximately 3–10% for the baseline and 6–18%, 8–20%, and 5–13% for the between‑method comparisons, respectively.
+Detailed results are provided in Supplementary Fig.
+S3.13.
+Posterior predictive checks showed that epidemic trajectories generated from all three methods fit the observed data equally well (Figure 3B), as evidenced by the close agreement in predictive energy scores (Table 6).
 Beyond these direct comparisons, the methods differed in sampling quality.
-PF achieved excellent convergence, with ESS exceeding 2000 and $\mathbf{\hat{R}}$ values close to 1, indicating robust posterior exploration and well-mixed chains (Supplementary Tables S6 and S7).
-CNF, in contrast, showed deviations in SBC calibration diagnostics (Supplementary Figure S23), reflecting limitations of the amortized posterior.
-Notably, these calibration issues were not apparent in the parameter recovery plots: recovered parameters aligned closely with the ground truth and error bars remained narrow (Supplementary Figure S27).
+PF achieved excellent convergence, with ESS exceeding 2000 and $\mathbf{\hat{R}}$ values close to 1, indicating robust posterior exploration and well-mixed chains (Supplementary Tables S3.2 and S3.3).
+CNF, in contrast, showed deviations in SBC diagnostics (Supplementary Fig.
+S3.14), reflecting limitations of the amortized posterior.
+Notably, these calibration issues were not apparent in the parameter recovery plots: recovered parameters aligned closely with the ground truth and error bars remained narrow (Supplementary Fig.
+S3.16).
+The miscalibration geometry also matches the observed offset of CNF relative to PF and HMC; for example, an upward‑sloping SBC histogram for $\beta$ corresponds to a left-shift in the $\beta$ marginal.
 This highlights that while CNF’s posterior uncertainty is imperfectly calibrated in a formal sense, the practical impact on point estimates and predictive accuracy is minimal.
@@
 Across datasets, both methods produced equivalent posterior distributions and predictive performance.
-Notably, CNF, despite being an amortized approach, achieved greater computational efficiency in our experiments, roughly 10-times (Supplementary Table S27), while PF excelled in sampling diagnostics.
+Notably, CNF, despite being an amortized approach, achieved greater computational efficiency in our experiments (roughly 10-fold) while PF excelled in sampling diagnostics.
 Moreover, since CNF relies purely on simulations it has consistent computational costs across different parameter vectors.
@@
 Additionally, the evaluation of the likelihood can be computationally expensive in regions where outcomes are highly sensitive to parameter changes or when parameter vectors do not fit the data well.
-Table 4: MAP and RMSE for SIR model.
-$\mathbf{\beta}$ $\mathbf{\gamma^{-1}}$ $\mathbf{RMSE_{infc}}$ $\mathbf{RMSE_{prev}}$
-True Par. 0.1000 20.00 0.0193 0.0230
-CNF 0.1026 18.99 0.0168 0.0201
-PF 0.1021 19.32 0.0150 0.0121
+Full diagnostics, including ESS, $\mathbf{\hat{R}}$, Wasserstein distance results, calibration plots, and parameter-recovery analyses are provided in Supplementary Information S3.
+Table 6: Predictive Energy Score for the SIR model.
+sir-1 sir-2 sir-3 sir-4 sir-5 sir-6
+True $0.046$ $0.057$ $0.032$ $0.020$ $0.052$ $0.021$
+CNF $0.046$ $0.040$ $0.033$ $0.033$ $0.052$ $0.018$
+PF $0.045$ $0.039$ $0.033$ $0.025$ $0.054$ $0.020$
+HMC $0.045$ $0.039$ $0.032$ $0.024$ $0.054$ $0.021$
+sir-7 sir-8 sir-9 sir-10 sir-11 sir-12
+True $0.050$ $0.054$ $0.056$ $0.071$ $0.061$ $0.073$
+CNF $0.048$ $0.053$ $0.057$ $0.069$ $0.061$ $0.066$
+PF $0.047$ $0.052$ $0.057$ $0.070$ $0.062$ $0.060$
+HMC $0.047$ $0.053$ $0.057$ $0.069$ $0.061$ $0.060$
 5.3 Non-Identifiability Issues Lead To Posterior Shape Differences
-Previous studies have shown that parameters in common epidemiological models are often weakly determined by commonly available datasets [raimundez2021covid].
-To assess how such non-identifiabilities affect inference reliability, we compared PF and CNF on the introduced two‑variant SEIR-model with pronounced non-identifiability, making it a suitable test case for evaluating how posterior estimates differ under CNF and PF.
-We first examined posterior distributions and predictive performance using four dense simulated observational time series.
+Previous studies have shown that parameters in common epidemiological models are often weakly determined by commonly available datasets Raimúndez et al.
+[2021].
+To assess how such non-identifiabilities affect inference reliability, we compared PF and CNF on the introduced two‑variant SEIR-model with pronounced non-identifiability for specific parameter combinations.
+We first examined posterior distributions and predictive performance using twelve dense simulated observational time series.
 Posterior distributions showed markedly lower variance compared to the corresponding priors, indicating strong information gain from the data (Figure 4).
-The notable exception is $\kappa^{-1}$ under CNF, whose marginal remained close to the prior.
-This behavior is typical in models where certain parameters are weakly informed by the data (see [wang2021posterior] for a more thorough discussion).
-The PF posterior approximation showed a higher contraction than that of CNF.
-Additionally, it was also largely enveloped in the CNF posterior approximation, across parameters (Figure 4A).
-Despite these width differences, epidemic trajectories generated from both methods fit the data equally well (Figure 4B), and MAP estimates together with RMSE scores were close to the true parameters (Table 5).
-PF estimates were marginally closer to the ground truth, highlighting the method’s focus on trajectory agreement through resampling.
+The notable exception is $\kappa^{-1}$ under CNF, whose marginal remained close to the prior for some experiments, consistent with weak identifiability Wang et al.
+[2021].
+PF posteriors were generally more concentrated than those of CNF and were largely enveloped by the CNF approximation across parameters (Figure 5A).
+Despite these width differences, both methods generally include the true parameters in their posteriors, with CNF achieving 95.2% marginal coverage and PF achieving 91.7%.
+Epidemic trajectories generated from both methods fit the data equally well (Figure 5B) with comparable energy scores (Table 7).
+PF produced marginally lower energy scores, highlighting the method’s focus on trajectory agreement through resampling.
 Figure 4: Marginal posterior approximations and prior distribution. Histograms of posterior approximations and prior distribution for the two-variant SEIR model using a dense dataset.
-As a secondary analysis, we assessed calibration and convergence diagnostics.
-CNF showed good calibration in ECDF plots (Supplementary Figure S20), though parameter recovery was imperfect for several parameters and markedly worse for $\kappa^{-1}$, which showed weak correlation with the ground truth (Supplementary Figure S26).
-PF chains displayed good convergence, with $\mathbf{\hat{R}}$ being close to one across parameters (Supplementary Table S14).
-However, ESS were relatively low ($\sim$500–600) for several parameters (Supplementary Table S13), indicating limited exploration of posterior tails.
+As an additional diagnostic, Wasserstein distances showed that variability across repeated CNF runs (CNF-CNF) was of the same order as the differences between CNF and PF (CNF-PF) indicating that the two methods agree up to the intrinsic stochasticity of the CNF estimator.
+CNF showed good calibration in ECDF plots, though parameter recovery was imperfect for several parameters and markedly worse for $\kappa^{-1}$, which showed weak correlation with the ground truth.
+PF chains displayed good convergence, with $\mathbf{\hat{R}}$ being close to one across parameters.
+However, ESS were relatively low (<1000<1000) for several parameter sets, indicating limited exploration of posterior tails.
 Figure 5: Evaluation of the posterior approximations for the two-variant SEIR model with a dense dataset.
@@
 B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
-Table 5: MAP and RMSE for full SEIR model.
-$\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa^{-1}}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I_{0}}$ $\mathbf{RMSE_{infc}}$ $\mathbf{RMSE_{prev}}$
-True Par. 17.00 5.00 0.0800 3.00 150.00 500.00 0.0200 0.0245
-CNF 15.47 4.93 0.0867 3.29 160.59 446.30 0.0237 0.0237
-PF 15.62 5.04 0.0855 3.22 153.02 479.16 0.0228 0.0291
+Table 7: Predictive Energy Score for the full SEIR2V model with dense data.
+d-1-1 d-1-2 d-2-1 d-2-2 d-3 d-4 d-5
+True $0.113$ $0.115$ $0.090$ $0.112$ $0.059$ $0.126$ $0.089$
+CNF $0.108$ $0.106$ $0.090$ $0.098$ $0.058$ $0.124$ $0.060$
+PF $0.107$ $0.104$ $0.088$ $0.098$ $0.057$ $0.123$ $0.059$
+d-6 d-7 d-8 d-9 d-10 d-11 d-12
+True $0.118$ $0.100$ $0.101$ $0.108$ $0.105$ $0.099$ $0.120$
+CNF $0.116$ $0.109$ $0.087$ $0.102$ $0.103$ $0.088$ $0.118$
+PF $0.115$ $0.114$ $0.088$ $0.101$ $0.097$ $0.086$ $0.117$
 In practice, observations are often non‑equidistant or partially missing.
@@
 Results for sparse datasets were consistent with those obtained under dense observations.
-Posterior distributions from CNF and PF showed similar shapes, and epidemic trajectories generated from both methods fit the data equally well (Supplementary Tables S8 to S13), with MAP estimates and RMSE values comparable to the dense setup (Supplementary Table S16).
-Additionally, CNF remained well calibrated and PF continued to show well‑mixed chains, though ESS values remained low for several parameters, mirroring the diagnostics observed in the dense case (Supplemetary Tables S17 and S18).
+Posterior distributions from CNF and PF showed similar shapes, and epidemic trajectories generated from both methods fit the data equally well, with energy scores of comparable magnitude (Table 8).
+Additionally, CNF remained well calibrated and PF continued to show well‑mixed chains, though ESS values remained low for several parameters, mirroring the diagnostics observed in the dense case.
 Together, these results indicate that both methods can accommodate sparse and irregular data without loss of inference quality, though CNF requires careful amortization design to handle heterogeneous observation schedules.
+Table 8: Predictive Energy Score for the full SEIR2V model with sparse data.
+s-1-1-1 s-1-1-2 s-1-1-3 s-1-2-1 s-1-2-2 s-1-2-3 s-3 s-4
+True $0.042$ $0.039$ $0.039$ $0.041$ $0.053$ $0.038$ $0.018$ $0.056$
+CNF $0.039$ $0.040$ $0.035$ $0.041$ $0.036$ $0.033$ $0.015$ $0.038$
+PF $0.037$ $0.038$ $0.032$ $0.033$ $0.035$ $0.033$ $0.015$ $0.038$
+s-5 s-6 s-7 s-8 s-9 s-10 s-11 s-12
+True $0.030$ $0.037$ $0.029$ $0.035$ $0.052$ $0.036$ $0.036$ $0.042$
+CNF $0.017$ $0.032$ $0.030$ $0.029$ $0.041$ $0.029$ $0.033$ $0.037$
+PF $0.015$ $0.032$ $0.030$ $0.029$ $0.038$ $0.024$ $0.030$ $0.036$
 Taken together, these results show that CNF and PF both provide accurate inference for the two‑variant SEIR model, even under pronounced non‑identifiability.
@@
 PF, in contrast, produces narrower posteriors and more certain point estimates, despite limited exploration of posterior tails due to low ESS.
-Predictive performance was high for both methods across dense and sparse datasets, with consistent MAP and RMSE values.
+Predictive performance was high for both methods across dense and sparse datasets, with consistent MAP and RMSE values predictive energy scores.
 Sparse observations introduced implementation overhead for CNFs but did not substantially affect inference quality for either method.
-Additionally, CNF were approximately 10-times faster independent of the used dataset (Supplementary Table S27).
+Additionally, CNF were approximately 10-times faster independent of the used dataset (Supplementary Table S1.22).
+Full diagnostics, including calibration plots, convergence statistics, and parameter‑recovery analyses, are provided in Supplementary Information S4 (dense datasets) and S5 (sparse datasets).
 5.4 Reducing Parameter Dependencies Improves Posterior Alignment
-Complementing the analysis of the full SEIR2v model, we evaluated the reparametrized version of the SEIR model to explore how reduced parameter dependencies affect inference results for PF and CNF.
-Posterior distributions from CNF and PF showed close alignment across parameters, indicating similar exploration of the parameter space (Figure 5A).
-Both methods yielded comparable MAP estimates near the true values, with minor deviations for $e_{0}$ and $I_{0}$, where PF shifted toward lower values.
-Epidemic trajectories fit the data equally well (Figure 5B) and RMSE values were similar across CNF, PF, and the true parameters, with PF showing marginally better agreement on the infection count data (Table S21).
+Complementing the analysis of the full SEIR2V model, we evaluated the reparametrized version of this model to explore how reduced parameter dependencies affect inference results for PF and CNF.To this end, we analysed twelve synthetic datasets in both the full and the reparametrized settings.
+In the full SEIR2V model, two datasets (r-2 and r-5) did not yield converged Markov chains under our PMMH workflow.
+In the reparametrized setting, posterior distributions from CNF and PF showed close alignment across parameters, indicating similar exploration of the parameter space (Figure 6A).
+Unlike in the full SEIR2V model, the posterior approximations exhibited no signs of long‑tailed behaviour.
+This is reflected in the 1‑Wasserstein distances in log‑parameter space, which were numerically smaller under reparametrization and decreased more strongly than the within‑method baselines (CNF–CNF and PF–PF).
+Moreover, CNF–PF distances were of the same order as the variability observed across independent runs of each method, indicating markedly improved cross‑method alignment.
+Epidemic trajectories fit the data equally well (Figure 6B) with energy scores between CNF, PF, and the true parameters being close to each other in both model settings (Table 9, Table 10).
 Additionally, we evaluated calibration and convergence diagnostics.
-CNF showed SBC miscalibration in ECDF plots (Figure S22), consistent with the results reported for the identifiable SIR model.
-PF, in contrast, achieved ESS exceeding 2000 (Supplementary Table S22), indicating well‑mixed chains and thorough posterior exploration.
+While results in the full model setting were consistent with the ones presented above, CNF showed SBC miscalibration in ECDF plots in the reparametrized setting(Supplementary Fig.
+S6.14), aligning with the results reported for the identifiable SIR model and hinting at a potential offset in the posterior.
+PF, in contrast, achieved ESS exceeding 2000 (Supplementary Table S6.2), indicating well‑mixed chains and thorough posterior exploration.
 Figure 6: Evaluation of the posterior approximations for the reparametrized two-variant SEIR model with dense data.
@@
 B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
-Table 6: MAP and RMSE for reparametrized SEIR model and dense datasets.
-$\mathbf{r0}$ $\mathbf{e0}$ $\mathbf{s0}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$ $\mathbf{RMSE_{infc}}$ $\mathbf{RMSE_{prev}}$
-True Par. 1.36 22.00 51.00 150.00 500.00 0.0200 0.0245
-CNF 1.34 21.33 50.27 145.46 489.71 0.0216 0.0236
-PF 1.34 20.57 51.21 156.94 467.30 0.0319 0.0271
+Table 9: Predictive Energy Score for the reparametrized SEIR2V model with dense data.
+d-1-1 d-1-2 r-1 r-2 r-3 r-4
+True Param. $0.1133$ $0.1153$ $0.0973$ $0.1386$ $0.0986$ $0.0997$
+CNF 0.1098 $0.1066$ $0.0832$ $0.1176$ $0.1003$ $0.1035$
+PF 0.1074 $0.1043$ $0.0799$ $0.1163$ $0.0935$ $0.1052$
+r-5 r-6 r-7 r-8 r-9 r-10
+True Param. $0.1455$ $0.1206$ $0.1463$ $0.1226$ $0.0913$ $0.0628$
+CNF 0.1264 $0.1166$ $0.1418$ $0.0980$ $0.0818$ 0.058 030.058\,03
+PF 0.1277 $0.1059$ $0.1384$ $0.0991$ $0.0792$ $0.0587$
+Table 10: Predictive Energy Score for the full SEIR2V model with dense data from the reparametrized setting.
+d-1-1 d-1-2 r-1 r-2 r-3 r-4
+True Param. $0.1132$ $0.1148$ 0.097 540.097\,54 $0.1388$ 0.098 260.098\,26 0.099 690.099\,69
+CNF $0.108$ $0.1067$ 0.092 690.092\,69 $0.1227$ $0.1004$ $0.1063$
+PF $0.1062$ $0.1035$ 0.079 610.079\,61 0.094 380.094\,38 $0.1005$
+r-5 r-6 r-7 r-8 r-9 r-10
+True Param. $0.1456$ $0.1203$ $0.146$ $0.124$ 0.091 420.091\,42 0.063 120.063\,12
+CNF $0.1327$ $0.1127$ $0.1445$ 0.098 980.098\,98 0.080 960.080\,96 0.057 730.057\,73
+PF $0.1059$ $0.1382$ 0.098 920.098\,92 $0.0796$ 0.057 370.057\,37
 In summary, reducing parameter dependencies improved identifiability and brought CNF and PF posterior shapes into closer agreement.
@@
 However, the reparametrization did not enhance predictive performance and relies on fixing $\kappa^{-1}$, which should be approached with caution in applied inference.
+Full diagnostics, including Wasserstein distance analyses, calibration plots, convergence statistics, and parameter‑recovery results, are provided in Supplementary Information 6 (reparametrized model) and Supplementary Information 7 (full model with datasets for the reparametrized model).
 5.5 CNF and PF yield consistent results on real-world datasets
-To complement the synthetic experiments, we applied both methods to data from a longitudinal cohort study of the COVID‑19 outbreak in Ethiopia [ethiopian_data].
+To complement the synthetic experiments, we applied both methods to data from a longitudinal cohort study of the COVID‑19 outbreak in Ethiopia [Gudina et al., 2021].
 Our setup differed from the original study in that we used a stochastic rather than deterministic formulation, offering greater flexibility and better reflecting the inherent randomness in infection dynamics.
@@
 Posterior approximations from CNF and PF shared a similar overall shape, with the PF posterior largely nested within the broader CNF and exhibiting reduced variance, particularly for the rate parameters $\beta$, $\kappa^{-1}$, and $\gamma^{-1}$ (Figure 7A).
-Trajectory ensembles derived from the inferred posteriors fit the data well, even under the relatively high noise level reported in the original study; one seroprevalence point was not captured by PF within its uncertainty intervals (Figure 6B).
-Despite this discrepancy, the PF MAP achieved RMSE values that matched or improved upon those reported in the original study.
-Although the CNF MAP deviated more strongly from both the published parameters and PF, the RMSE values of the respective distributions (Table S24) were comparable and showed improved accuracy for infection count data.
-Diagnostics were consistent with patterns observed in controlled experiments: CNF exhibited good calibration alongside imperfect parameter recovery (Figure S21, Figure S25), while PF demonstrated convergence despite low effective sample sizes (Table S25, Table S26).
+Trajectory ensembles derived from the inferred posteriors fit the data well, even under the relatively high noise level reported in the original study; one seroprevalence point was not captured by either method within their uncertainty intervals (Figure 7B).
+Predictive energy scores between CNF and PF improved on the energy score calculated from the the originally published parameter set, indicating that both inference methods provide a better probabilistic fit to the observed data than the originally published parameter set (Table 11).
+Diagnostics were consistent with patterns observed in controlled experiments: CNF exhibited good calibration alongside imperfect parameter recovery, while PF demonstrated convergence despite low effective sample sizes.
 Figure 7: Evaluation of the posterior approximations for the two-variant SEIR model with real data.
-A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade.
+A Posterior approximations based on 10,000 samples, publication parameters and their reported standard deviation are indicated by bold black lines and joint MAP estimates appear in a darker shade.
 B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
-Table 7: MAP and RMSE for full SEIR model for real dataset.
-$\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa^{-1}}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I_{0}}$ $\mathbf{RMSE_{infc}}$ $\mathbf{RMSE_{prev}}$
-Pub. Par. 16.70 5.00 0.0800 2.30 184.50 635.00 0.0737 0.0702
-CNF 14.80 4.14 0.0882 3.16 183.04 550.16 0.1251 0.0928
-PF 15.03 4.67 0.0858 3.36 160.35 752.08 0.0856 0.0606
+Table 11: Predictive Energy Score for the full SEIR2V model with real-world data.
+Ethiopia
+Pub. 0.131
+CNF 0.107
+PF 0.096
 Taken together, applying both methods to the Ethiopian cohort data yielded results consistent with the synthetic analyses, underscoring the robustness of each approach.
 The trade‑offs identified in controlled settings—CNF’s broader posterior coverage versus PF’s tighter estimates with lower ESS—carried over to this real‑world application, while predictive performance remained strong for both methods.
+Full diagnostics, including calibration plots, convergence statistics, and parameter‑recovery results, are provided in Supplementary Information S8.
 6 Discussion
 Responses to public health crises, such as outbreaks of infectious diseases, creates a need for informed decision making by realistically modeling the dynamics using stochastic models.
-Yet, fitting such models to observed data present many difficulties and several methods were developed address this challenge.
+Yet, fitting such models to observed data present many difficulties and several methods were developed to address this challenge.
 In this work, we systematically evaluated two state-of-the-art methods for Bayesian parameter inference, Conditional Normalizing Flows (CNF) and pseudo-marginal particle MCMC, referred to as Particle Filters (PF), across a range of stochastic compartmental models.
 Our analysis encompassed synthetic and real-world data, varied model structure, and different levels of data sparsity, yielding crucial insights into the strengths and limitations of each methodology when applied in epidemiological modeling scenarios.
+Where feasible, we calculated a reference posterior using Hamiltonian Monte Carlo (HMC) on a finely discretized version of the stochastic models.
 Across all investigated scenarios, both methods produced accurate fits and informative posteriors.
 However, this study revealed systematic differences of uncertainty in the parameter space due to different posterior geometries and model design choices.
-In the context of an identifiable stochastic model, such as the SIR model, both CNF and PF yielded highly consistent posterior distributions and closely matched predictive epidemic trajectories.
-The marginal and joint posterior estimates and MAP estimates were near the data-generating parameters and the root mean square error (RMSE) values confirmed high predictive accuracy for both methods, despite imperfect calibration of the CNFs.
-Additionally, CNF exhibited very good recovery plots in this setting.
-As a exact method, PF showed a good recovery of the true parameters in this setting, that were closely matched by the CNF.
-Together, this indicates that, once trained, CNF can deliver inference quality on a par with an MCMC based baseline while enabling possible amortization across different datasets and yielding high computational efficiency in the inference phase.
+In the context of small stochastic models, such as the SIR model and the SIS model, both CNF and PF yielded highly consistent posterior distributions and closely matched predictive epidemic trajectories.
+The marginal and joint posterior estimates and MAP estimates were near the data-generating parameters and the predictive energy scores confirmed high predictive accuracy for both methods.
+CNFs exhibited miscalibration for the SIR model, which resulted in small offsets relative to PF and HMC.
+In contrast to that, CNF exhibited very good recovery plots in these settings.
+As a exact method, PF showed a good recovery of the true parameters and a high agreement with the reference posterior distributions, that were closely matched by the CNF.
+Together, this indicates that, once trained, CNF can deliver inference quality on a par with an exact MCMC based baseline while enabling possible amortization across different datasets and yielding high computational efficiency in the inference phase but miscalibration needs to be handled and analysed carefully.
 Expanding the analysis, we applied both methods to a more complex stochastic two-variant SEIR model with practical non-identifiabilities and parameter couplings.
@@
 This property can limit tail exploration in the parameter space especially when the posterior surface contains ridges or mountain-pass structures.
+Additionally, PF based methods are, as most MCMC methods, highly dependent on the choice of the initial parameter value.
+Uninformative initial parameter values lead to degenerate likelihood estimate and therefore stuck Markov chains.
 CNF achieved good calibration results in this setting, although its recovery plots indicated that parameter recovery was not perfect.
-The wider posteriors for CNF contrasts with prior literature where amortized inference often underestimates uncertainty [hermans2022trust].
-Despite these differences, both methods captured the inference data accurately and showed strong predictive performance, as evidenced by low and similar RMSEs.
+The wider posteriors for CNF contrasts with prior literature where amortized inference often underestimates uncertainty [Hermans et al., 2022].
+Despite these differences, both methods captured the inference data accurately and showed strong predictive performance, as evidenced by low and similar predictive energy scores.
 Their application in this setting shows that amortized methods may struggle with calibration, yet still provide strong predictive performance and exact MCMC based method estimate the bulk of the posterior well while underestimating its width caused by too less parameter space explorance.
 Using sparsely sampled datasets with different missingness patterns did not change the qualitative estimation results of both methods.
-But they inherently treat sparse datasets differently.
+But the two methods inherently treat sparse datasets differently.
 While CNF methods can accommodate such dataset variability, amortizing over all relevant combinations of time points may become computationally expensive and risks unreliable extrapolation if not carefully designed.
@@
 Finally, applying both methods to a real dataset from a longitudinal cohort study in Ethiopia confirmed the simulation study findings and yielded broadly similar posterior approximations and epidemic trajectories.
-Although obtained MAP estimates differed from previously published parameter values, especially for CNFs, both methods captured the data dynamics and produced model fits with RMSEs close to or improving upon the deterministic analysis [ethiopian_data].
+Although obtained MAP estimates differed from previously published parameter values, especially for CNFs, both methods captured the data dynamics and produced model fits with predictive energy scores close to or improving upon the deterministic analysis [Gudina et al., 2021].
 By demonstrating their effectiveness in a real-world setting, our analysis underscores the potential of these methods to inform public health strategies and decision-making processes.
@@
 To achieve greater speed, CNF can make use of a parallelized simulator or training the neural network on a GPU.
-Meanwhile, PF could benefit from the use of more advanced filtering algorithms such as guided PFs [snyder_2011, chopin_2020] or using correlated particles [wiqvist_2021] to avoid degeneracy or too high particle numbers.
-Switching to a more advanced outer MCMC sampler such as a Gibbs sampler [chopin2015] or a Metropolis-Adjusted Langevin algorithm [lowe2023accelerating] and parallelization over particles could speed-up the inference procedure as well.
+Meanwhile, PF could benefit from the use of more advanced filtering algorithms such as guided PFs [Snyder, 2011; Chopin and Papaspiliopoulos, 2020] or using correlated particles [Wiqvist et al., 2021] to avoid degeneracy or too high particle numbers.
+Switching to a more advanced outer MCMC sampler such as a Gibbs sampler [Chopin and Singh, 2015] or a Metropolis-Adjusted Langevin algorithm [Lowe, 2023] and parallelization over particles could speed-up the inference procedure as well.
 Additionally, the lack of amortization renders assessment of calibration and hyperparameter tuning for PF too computationally expensive.
@@
 Our use of an Euler–Maruyama integration with positivity clipping introduces discretization error whose interaction with inference can lead to numerical instabilities near boundaries or in small populations.
-Algorithmic choices were guided by the ease of their applicability in praxis: We focused on a bootstrap PF within PMMH and a tuned CNF architecture with consistent hyperparameters and fixed simulation budget across our datasets.
+Algorithmic choices were guided by the ease of their applicability in praxis: We focused on a bootstrap PF within PMMH and a tuned CNF architecture with consistent hyperparameters and fixed simulation budget across our datasets and experimental setups.
 Alternative PF proposals (e.g., guided/auxiliary filters or tempering within SMC) and alternative SBI estimators (likelihood or ratio estimation) or more exhaustive hyperparameter-tuning could shift efficiency–accuracy trade‑offs.
-To avoid out‑of‑distribution failure modes, CNFs were trained per experiment rather than globally amortized across all data regimes, which tempers their wall‑clock advantage in this study; broader amortization remains attractive but demands careful conditioning on meta‑data such as time‑grid irregularity and noise scales [radev2023bayesflow, wang2024missing].
-In the SBI literature, recent advances focus on workflows that narrow the gap between neural SBI and Monte Carlo by using learned proposals inside sequential Monte Carlo or MCMC [papamakarios2019sequential, hermans2020likelihood].
+To avoid out‑of‑distribution failure modes, CNFs were trained per experiment rather than globally amortized across all data regimes, which tempers their wall‑clock advantage in this study; broader amortization remains attractive but demands careful conditioning on meta‑data such as time‑grid irregularity and noise scales [Radev et al., 2023; Wang et al., 2024].
+In the SBI literature, recent advances focus on workflows that narrow the gap between neural SBI and Monte Carlo by using learned proposals inside sequential Monte Carlo or MCMC [Papamakarios et al., 2019; Hermans et al., 2020].
 Variational and amortized SMC approaches train proposal families to reduce weight degeneracy and improve effective sample sizes; conversely, particle methods can provide high‑quality samples to bootstrap or regularize neural estimators.
@@
 These methods could combine the tail coverage and speed of neural estimators with the asymptotic correctness and model flexibility of pseudo‑marginal schemes.
-Recent generative modeling advancements, i.e., flow matching and diffusion‑style transports, offer further avenues for robust posterior learning with stronger inductive biases and improved stability [lipman2022flow, draxler2024free, arruda2025].
+Recent generative modeling advancements, i.e., flow matching and diffusion‑style transports, offer further avenues for robust posterior learning with stronger inductive biases and improved stability [Lipman et al., 2022; Draxler et al., 2024; Arruda et al., 2025].
 Additionally, the use of adaptive prior schemes where the prior is updated in light of the available data, through approaches such as empirical Bayes, hierarchical modeling, or iterative refinement, may help to mitigate issues of non-identifiability while improving robustness.
@@
 By clarifying the trade‑offs and giving insights into the performance of these two practical inference methods, this work provides the groundwork for evaluating inference engines ready for uncertainty‑aware epidemic forecasting and real-time support of public-health decisions.
+Supporting Information
+S1 Text: Supporting Information
+S2 Results: Supplementary Results SIS-Model
+S3 Results: Supplementary Results SIR-Model
+S4 Results: Supplementary Results SEIR-Model with Dense Data
+S5 Results: Supplementary Results SEIR-Model with Sparse Data
+S6 Results: Supplementary Results Reparametrized SEIR-Model
+S7 Results: Supplementary Results Comparison Full SEIR-Model with Reparametrized Model
+S8 Results: Supplementary Results SEIR-Model with Real Data
 Author Contributions
@@
 This work received further funding from the Initiative and Networking Fund of the Helmholtz Association (grant agreement number KA1-Co-08, Project LOKI-Pandemics).
-Appendix A Supplementary Information
-A.1 Model Equations for the Stochastic Models
-SIR Model
-Set of stochastic differential equations describing the time evolution of individuals in the compartments of the two variant SEIR model.
-$dS(t)$ $=-\beta\frac{S(t)I(t)}{N(t)}dt-\sqrt{\beta\frac{S(t)I(t)}{N(t)}}dB_{1}(t)$
-$dI(t)$ $=\beta\frac{S(t)I(t)}{N(t)}dt-\gamma I(t)dt+\sqrt{\beta\frac{S(t)I(t)}{N(t)}}dB_{1}(t)-\sqrt{\gamma I(t)}dB_{2}(t)$
-$dR(t)$ $=\gamma I(t)dt+\sqrt{\gamma I(t)}dB_{2}(t)$
-where $B_{j},\,j\in\{1,...,2\}$ are independent Brownian Motions. Parameters with respective prior distributions and bounds are given in Table S3.
-Two-Variant SEIR Model
-Set of stochastic differential equations describing the time evolution of individuals in the compartments of the two variant SEIR model.
-$dS(t)$ $=-\mathrm{\beta}_{\text{wt}}\frac{S(t)\mathrm{I}_{\text{wt}}(t)}{N(t)}dt-\beta_{var}\frac{S(t)(I_{var}(t)+I_{both}(t))}{N(t)}dt$
-$\quad-\sqrt{\beta_{wt}\frac{S(t)I_{wt}(t)}{N(t)}}dB_{1}(t)-\sqrt{\beta_{var}\frac{S(t)(I_{var}(t)+I_{both}(t))}{N(t)}}dB_{7}(t)$
-$dE_{wt}(t)$ $=\beta_{wt}\frac{S(t)I_{wt}(t)}{N(t)}dt-\kappa_{wt}E_{wt}(t)dt+\sqrt{\beta_{wt}\frac{S(t)I_{wt}(t)}{N(t)}}dB_{1}(t)-\sqrt{\kappa_{wt}E_{wt}(t)}dB_{2}(t)$
-$dI_{wt}(t)$ $=\kappa_{wt}E_{wt}(t)dt-\gamma_{wt}I_{wt}(t)dt+\sqrt{\kappa_{wt}E_{wt}(t)}dB_{2}(t)-\sqrt{\gamma_{wt}I_{wt}(t)}dB_{3}(t)$
-$dR_{wt}(t)$ $=\gamma_{wt}I_{wt}(t)dt-\beta_{var}\frac{R_{wt}(t)(I_{var}(t)+I_{both}(t))}{N(t)}dt$
-$\quad+\sqrt{\gamma_{wt}I_{wt}(t)}dB_{3}(t)-\sqrt{\beta_{var}\frac{R_{wt}(t)(I_{var}(t)+I_{both}(t))}{N(t)}}dB_{4}(t)$
-$dE_{var}(t)$ $=\beta_{var}\frac{S(t)(I_{var}(t)+I_{both}(t))}{N(t)}dt-\kappa_{var}E_{var}(t)dt$
-$\quad+\sqrt{\beta_{var}\frac{S(t)(I_{var}(t)+I_{both}(t))}{N(t)}}dB_{7}(t)-\sqrt{\kappa_{var}E_{var}(t)}dB_{8}(t)$
-$dI_{var}(t)$ $=\kappa_{var}E_{var}(t)dt-\frac{\gamma_{var}}{1.35}I_{var}(t)dt+\sqrt{\kappa_{var}E_{var}(t)}dB_{8}(t)-\sqrt{\frac{\gamma_{var}}{1.35}I_{var}(t)}dB_{9}(t)$
-$dR_{var}$ $=\frac{\gamma_{var}}{1.35}I_{var}(t)dt+\sqrt{\frac{\gamma_{var}}{1.35}I_{var}(t)}dB_{9}(t)$
-$dE_{both}(t)$ $=\beta_{var}\frac{R_{wt}(t)(I_{var}(t)+I_{both}(t))}{N(t)}dt-\kappa_{var}E_{both}(t)dt$
-$\quad+\sqrt{\beta_{var}\frac{R_{wt}(t)(I_{var}(t)+I_{both}(t))}{N(t)}}dB_{4}(t)-\sqrt{\kappa_{var}E_{both}(t)}dB_{5}(t)$
-$dI_{both}(t)$ $=\kappa_{var}E_{both}(t)dt-\gamma_{var}I_{both}(t)dt+\sqrt{\kappa_{var}E_{both}(t)}dB_{5}(t)-\sqrt{\gamma_{var}I_{both}(t)}dB_{6}(t)$
-$dR_{both}(t)$ $=\gamma_{var}I_{both}(t)dt+\sqrt{\gamma_{var}I_{both}(t)}dB_{6}(t),$
-where $B_{j},\,j\in\{1,...,7\}$ are independent Brownian Motions. Parameters with respective prior distributions and bounds are given in Table S10.
-A.2 Computation of standard deviation for the noise model
-The standard deviations of the noise model introduced in Section 2.3 were derived from binomial count data. This section outlines the rationale and computation procedure in detail.
-When conducting Bernoulli trials, such as testing individuals in a population, the sum of positive test is adequately modeled by a binomial distribution, provided the trials are independent and the success probability remains constant.
-If the number $n$ of such Bernoulli trials conducted is large or the success probability $p$ is neither exceedingly low nor high, the resulting binomial distribution $B(n,p)$ is well approximated by a normal distribution $\mathcal{N}(np,np(1-p))$, thanks to the de Moivre-Laplace Theorem, a special case of the central limit Theorem [supp-walker1985moivre].
-This approximation allows us to compute the standard deviation of the measurement noise directly from the count data and the known population size, and to incorporate it into the noise model.
-However, this approximation may become inaccurate for small sample sizes or highly skewed probabilities, where alternative methods or exact distributions might be preferable.
-When generating our synthetic data, we chose $n=500$, which is sufficiently large to justify the Gaussian approximation.
-A.3 Practical non-identifiability of the two-variant SEIR model
-In our comparison we observed correlated parameter estimates leading to posterior misalignment between the methods.
-In several parameter combinations, the posterior mass was concentrated along low‑dimensional manifolds in parameter space, indicating practical non‑identifiability.
-This manifold can be approximated in terms of the following reparameterization into epidemiologically interpretable quantities:
-[…]
-where $r_{0}$ is the basic reproduction number, $e_{0}$ the mean latent-plus-infectious period, i.e., mean time an individual spends from catching the infection until it stops being infectious, and $s_{0}$ the effective detection factor integrated over the average infectious period, so how many cases are likely to be detected per infectious person, averaged over their infectious period.
-A representative example is shown in Figure S2.
-The substantial overlap between the trajectories shows that combinations of widely different parameters can produce virtually identical epidemic dynamics.
-A.4 Tuning of Particle Filter
-Particle filters combined with MCMC algorithms are convincing considering their theoretical implications, but in practice tuning the performance of these algorithm can take a lot of work.
-For a detailed review of efficient implementations of marginal MCMC sampling, the reader may refer to [supp-doucet_2015], a numerical study on tuning PMMH is given in Chapter 16 of [supp-chopin_2020].
-A.4.1 Hyperparamters and diagnostics
-After deciding on a inner particle filter and an outer MCMC scheme, the only but also most crucial choice is the number of particles $N$.
-Although any number of particles $N\geq 1$ leaves the target density invariant and theoretically leads to the correct results, the number of particles highly affects the variance of the likelihood estimate and the algorithm’s ability to converge to the target distribution.
-Hence, choosing $N$ resembles the trade-off between a lower variance in the likelihood estimator, efficiency, and a higher computational cost.
-Based on the empirical results of experiments in [supp-chopin_2020] and theoretical considerations in [supp-doucet_2015] the number of particles should satisfy the following two criteria
-  • The variance of the log-likelihood estimate is less than 1.
-  • There exists more than one unique path of particles and their ancestors over time.
-Both criteria are satisfied for the SIR and SEIR model with a particle number of 100 (Figure S3, when evaluated around the data generating parameters.
-To avoid particle degeneracy issues as much as possible in the two-variant SEIR model, we use a higher number of $N=200$ for that model.
-In addition, we evaluated the convergence of the chains based on an improvement of the traditional Gelman-Rubin $\hat{R}$ criterium [supp-vehtari2021rank], where a value close to 1.0 is considered good convergence.
-To assess the mixing properties of the chain, we also monitored the effective samples size with a maximal lag for the autocorrelation of 250.
-For a given estimand, it is recommended that the ESS should be at least $100\cdot\text{nchains}$ [supp-vehtari2021rank].
-Both criteria were evaluated using the last 10,000 samples of the chains.
-A.4.2 Resampling in Importance Sampling
-An additional choice, that needs to be made, when using the Bootstrap filter with resampling in it, is the resampling scheme.
-As in importance resampling algorithms in particle filters resampling increases the variance in the past because ancestors are resampled.
-However, the more diverse particles decrease the variance in the present.
-Hence, resampling itself can be seen as sacrificing the past to save the present and is especially helpful when low weights are present.
-If all weights are similar, then resampling increases the variance without helping.
-Without resampling, the Particle Filter would fall back to a single importance sampling step and be prone to the curse of dimensionality.
-Conversely, the option of resampling every time introduces a step with a computational complexity of $O(N)$ at every time step and might add additional variance if the weights are all similar.
-Therefore, in practice, one would not resample in every iteration step, but once the variability of the weights becomes too large.
-Commonly, one calculates the effective sample size (ESS) to estimate this variability and triggers resampling if it falls a certain threshold $\mathrm{ESS_{min}}$, often $N/2$ is used [supp-chopin_2020].
-Let $\{w^{(i)},X^{(i)}\}_{i=1}^{N}$ be the set of weighted samples following the target distribution with density $q(x)$ and $M^{(i)}$ denote the number of offspring associated with $X^{(i)}$.
-The resampling process is equivalent to sampling $M^{(1:N)}=(M^{(1)},...,M^{(N)})$.
-Different schemes are proposed in the literature [supp-doucet_particle_tutorial, supp-chopin_2020] to obtain unbiased approximations via resampling, including multinomial, residual and systematize resampling.
-Among those, the systematic resampling is the best as it is easy to implement, very fast, and outperforms other schemes in terms of lower-variance estimators.
-This method performs sampling from a multinomial distribution with less variance.
-Naively, one samples i.i.d.
-random variables $\{U_{i}\}_{i=1,...,N}$ from $\mathcal{U}([0,1))$ and then applies the inverse cumulative distribution function to transform them to the desired samples $M^{(i)}=\mathrm{cdf}^{-1}(U_{i})$.
-A good strategy to reduce the variance of this approach is to replace the i.i.d.
-values with values covering $[0,1)$ more regularly.
-To achieve this, one simulates a single uniformly distributed random variable $u\sim\mathcal{U}([0,\frac{1}{N}))$ and sets $U_{i}=u+\frac{i-1}{N}$ for i=1,,Ni=1,...,N.
-Then we obtain $M^{(i)}$ by
-$M^{(i)}=\left|\left\{U_{j}:\sum_{k=1}^{i-1}w^{(k)}\leq\ U_{j}<\sum_{k=1}^{i}w^{(k)}\right\}\right|,$
-where we use the convention $\sum_{k=1}^{0}:=0$
-For a more thorough comparison of different resampling methods, the reader may refer to [supp-Doucet_SMC, supp-resampling_schemes_comparison, supp-doucet_particle_tutorial].
-A.5 Hyperparameters for Amortized Bayesian Inference
-Hyperparameters for the Normalizing-Flows workflow were mostly fixed across all experiments.
-Only the summary dimension was varied between models.
-For detailed configurations, see Tables S28, S29, S30, and S31.
-A.6 Synthetic Datasets
-In this work, we used two stochastic compartmental models for a simulation study, where for each model we created several datasets.
-For the standard SIR model we created two densely sampled (observation time points every 10 days) sets with simulations using the parameters in Table S4.
-For the two variant SEIR model we used two sets of parameters (Table S20) and for each parameter set generated two densely sampled (observation time points every 7 days) datasets.
-This results in 4 datasets named ”Dense aba-b”, where $a$ refers to the used parameter set and $b$ to the simulation round from that parameter set.
-To mimic realistic data sparsity we drew three integers controlling the number of infection‑count timepoints, the number of seroprevalence timepoints, and the number of timepoints observed by both modalities.
-Then we randomly selected timepoints from the dense datasets according to these numbers.
-We repeated this procedure three times for each of the two dense datasets corresponding to the first parameter vector.
-This results in 6 sparse datasets named ”Sparse abca-b-c, where $a$ indicates the parameter set, $b$ the simulation round, $c$ the subsampling round.
-Appendix B Supplementary Tables
-Table S1: SIR model description Compartments and their respective description for the SIR model.
-Compartment Description
-$\mathrm{S}_{\text{}}$ Susceptible; individuals may become infected upon contact with Infectious individuals.
-$\mathrm{I}_{\text{}}$ Infectious; individuals infected with the disease.
-$\mathrm{R}_{\text{}}$ Recovered; individuals who have recovered from an infection.
-Table S2: Parameters and their description for SIR model.
-Parameter Description
-$\beta$ Transmission risk for an individual susceptible to the disease.
-$\gamma^{-1}$ Average infectious time.
-Table S3: List of model parameters, bounds and priors for the SIR model
-Parameter Lower bound Upper bound Prior
-$\beta$ $0.95\cdot\gamma$ $5\cdot\gamma$ $\mathrm{Uniform}(0,1)$
-$\gamma^{-1}$ 1 30 $\mathrm{Uniform}(1,30)$
-Table S4: Parameters for synthetic datasets of the SIR model.
-Dataset $\boldsymbol{\beta}$ $\boldsymbol{\gamma^{-1}}$
-SIR 1 0.1 20
-SIR 2 0.22 5
-Table S5: MAP and RMSE for SIR model.
-Dataset $\mathbf{\beta}$ $\mathbf{\gamma^{-1}}$ $\mathbf{RMSE_{infc}}$ $\mathbf{RMSE_{prev}}$
-Dense 1 True Par. 0.1000 20.00 0.0193 0.0230
-CNF 0.1026 18.99 0.0168 0.0201
-PF 0.1021 19.32 0.0150 0.0121
-Dense 2 True Par. 0.2200 5.00 0.0081 0.0298
-CNF 0.2284 4.90 0.0074 0.0259
-PF 0.2331 4.80 0.0090 0.0250
-Table S6: Effective sample sizes (ESS) per parameter and data for the SIR model.
-ESS computed on the last 10,000 samples of the chains resulting from running the PF method on the standard SIR model and using a maximum lag size of 250 for the autocorrelation.
-Dataset $\boldsymbol{\beta}$ $\boldsymbol{\gamma^{-1}}$
-SIR 1 5189.6 5131.9
-SIR 2 2259.2 2256.62
-Table S7: $\mathbf{\hat{R}}$ diagnostics for the SIR model.
-Model $\boldsymbol{\beta}$ $\boldsymbol{\gamma^{-1}}$
-SIR 1 1.001 1.000
-SIR 2 1.003 1.003
-Table S8: Two-variant SEIR model description Compartments and their respective description for the two-variant SEIR model.
-Compartment Description
-$\mathrm{S}_{\text{}}$ Susceptible; individuals may become infected upon contact with Infectious individuals.
-$\mathrm{E}_{\text{wt}}$ Exposed; individuals with no prior infection and infected with the wild-type but not yet infectious.
-$\mathrm{E}_{\text{var}}$ Exposed; individuals with no prior infection and infected with the variant but not yet infectious.
-$\mathrm{E}_{\text{wt·var}}$ Exposed; individuals with a prior infection from the wild-type and infected with the variant but not yet infectious.
-$\mathrm{I}_{\text{wt}}$ Infectious; individuals with no prior infection, infected with the wild-type and currently infectious.
-$\mathrm{I}_{\text{var}}$ Infectious; individuals with no prior infection, infected with the variant and currently infectious.
-$\mathrm{I}_{\text{wt·var}}$ Infectious; individuals infected with the variant, infectious and who have previously been infected by the wild-type.
-$\mathrm{R}_{\text{wt}}$ Recovered; individuals who have recovered from a wild-type infection, not currently infected.
-$\mathrm{R}_{\text{var}}$ Recovered; individuals who have recovered from a variant infection, not currently infected.
-$\mathrm{R}_{\text{wt·var}}$ Recovered; individuals recovered from a reinfection with the variant, after having have recovered from a prior wild-type infection.
-Table S9: Parameters and their description for the two-variant SEIR model.
-Parameter Description
-$\beta$ Transmission risk for an individual susceptible to one of the two-variants.
-$\kappa^{-1}$ Average latency period length for the wild-type and the variant.
-$\gamma^{-1}$ Average infectious time for the wild-type.
-s
-Scaling factor, measuring the discrepancy between (measured
-and reported) newly infected and actually infected.
-$\mathrm{I}_{\text{0}}$ Number of individuals infected with the wild-type at start of the simulation.
-$\mathrm{t}_{\text{var}}$ Time at which $n_{\textit{var}}$ individuals are introduced into the compartment $\mathrm{I}_{\text{var}}$.
-Table S10: List of model parameters, bounds and priors for the two variant SEIR model. Prior distributions were taken from [supp-ethiopian_data].
-Parameter Lower bound Upper bound Prior
-$\gamma^{-1}$ 0 $\infty$ $\mathrm{Normal}(15.7,6.7)$
-$\kappa$ 0 $\infty$ $\mathrm{Lognormal}(1.63,0.5)$
-$\beta$ $0.95\cdot\gamma$ $4\cdot\gamma$ $\mathrm{Uniform}(0,1)$
-$s$ 0.1 10 $\mathrm{Uniform}(0.1,10)$
-$t_{var}$ 120 360 $\mathrm{Uniform}(120,360)$
-$I0$ 10 1000 $\mathrm{Uniform}(10,1000)$
-Table S11: Parameters for dense synthetic datasets of the reparametrized SEIR model.
-Dataset $\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa^{-1}}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$
-Dense 1-1 17 5 0.08 3 150 500
-Dense 1-2 17 5 0.08 3 150 500
-Dense 2-1 11.7 8.4 0.23 1.6 222 560
-Dense 2-2 11.7 8.4 0.23 1.6 222 560
-Table S12: MAP and RMSE for full SEIR model for dense datasets.
-Dataset $\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa^{-1}}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I_{0}}$ $\mathbf{RMSE_{infc}}$ $\mathbf{RMSE_{prev}}$
-Dense 1-1 True Par. 17.00 5.00 0.0800 3.00 150.00 500.00 0.0200 0.0245
-CNF 15.47 4.93 0.0867 3.29 160.59 446.30 0.0237 0.0237
-PF 15.62 5.04 0.0855 3.22 153.02 479.16 0.0228 0.0291
-Dense 1-2 True Par. 17.00 5.00 0.0800 3.00 150.00 500.00 0.0214 0.0237
-CNF 15.00 4.95 0.0902 3.28 161.67 374.69 0.0188 0.0215
-PF 15.04 4.26 0.0890 3.25 165.57 366.60 0.0170 0.0234
-Dense 2-1 True Par. 11.70 8.40 0.2300 1.60 222.00 560.00 0.0191 0.0170
-CNF 12.39 7.13 0.2162 1.53 223.95 511.84 0.0198 0.0194
-PF 11.55 7.90 0.2313 1.61 224.40 508.96 0.0242 0.0171
-Dense 2-2 True Par. 11.70 8.40 0.2300 1.60 222.00 560.00 0.0197 0.0206
-CNF 15.32 5.06 0.1774 1.26 222.82 541.31 0.0198 0.0210
-PF 13.97 6.26 0.1947 1.34 221.29 592.14 0.0208 0.0184
-Table S13: Effective sample sizes (ESS) per parameter and model for dense datasets.
-ESS computed on the last 10,000 samples of the chains resulting from running the PF method on the two-variant SEIR model and using a maximum lag size of 250 for the autocorrelation.
-Parameter $\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$
-Model
-Dense 1-1 501.6 658.8 514.8 528.8 1174.2 1537.4
-Dense 1-2 436.2 539.8 401.8 428.6 1493.4 1768.6
-Dense 2-1 445.3 459.9 440.9 457.8 1676.1 1842.0
-Dense 2-2 1549.2 1633.8 1608.7 1652.8 1954.5 1687.8
-Table S14: $\mathbf{\hat{R}}$ diagnostics for the SEIR model with dense data.
-Dataset $\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$
-Dense 1-1 1.023 1.010 1.022 1.022 1.007 1.005
-Dense 1-2 1.025 1.014 1.024 1.021 1.004 1.002
-Dense 2-1 1.015 1.012 1.008 1.022 1.007 1.009
-Dense 2-2 1.014 1.009 1.008 1.022 1.007 1.004
-Table S15: Parameters for dense synthetic datasets of the reparametrized SEIR model. The sparse set Sparse x-y-z is subsampled from the dense set Dense x-y
-Dataset $\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa^{-1}}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$
-Sparse 1-1-1 17 5 0.08 3 150 500
-Sparse 1-1-2 17 5 0.08 3 150 500
-Sparse 1-1-3 17 5 0.08 3 150 500
-Sparse 1-2-1 17 5 0.08 3 150 500
-Sparse 1-2-2 17 5 0.08 3 150 500
-Sparse 1-2-3 17 5 0.08 3 150 500
-Table S16: MAP and RMSE for full SEIR model for sparse datasets.
-Dataset $\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa^{-1}}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I_{0}}$ $\mathbf{RMSE_{infc}}$ $\mathbf{RMSE_{prev}}$
-Sparse 1-1-1 True Par. 17.00 5.00 0.0800 3.00 150.00 500.00 0.0229 0.0220
-CNF 15.69 4.29 0.0871 3.16 171.04 331.82 0.0302 0.0155
-PF 14.65 5.52 0.0933 3.29 165.29 349.35 0.0371 0.0289
-Sparse 1-1-2 True Par. 17.00 5.00 0.0800 3.00 150.00 500.00 0.0200 0.0215
-CNF 15.95 5.06 0.0841 2.91 149.45 440.37 0.0298 0.0357
-PF 14.44 5.12 0.0913 3.12 147.55 443.22 0.0247 0.0195
-Sparse 1-1-3 True Par. 17.00 5.00 0.0800 3.00 150.00 500.00 0.0201 0.0101
-CNF 14.26 5.27 0.0945 3.65 174.46 395.00 0.0240 0.0320
-PF 13.17 5.47 0.1023 3.81 187.08 335.15 0.0292 0.0306
-Sparse 1-2-1 True Par. 17.00 5.00 0.0800 3.00 150.00 500.00 0.0229 0.0221
-CNF 16.23 4.76 0.0854 3.03 168.65 345.33 0.0281 0.0335
-PF 12.82 3.62 0.1016 3.74 180.59 259.94 0.0251 0.0396
-Sparse 1-2-2 True Par. 17.00 5.00 0.0800 3.00 150.00 500.00 0.0222 0.0272
-CNF 14.74 5.05 0.0913 3.26 153.58 363.09 0.0107 0.0269
-PF 15.03 3.96 0.0891 3.23 165.95 290.72 0.0262 0.0183
-Sparse 1-2-3 True Par. 17.00 5.00 0.0800 3.00 150.00 500.00 0.0134 0.0229
-CNF 16.74 4.48 0.0838 3.10 171.60 355.96 0.0108 0.0223
-PF 16.49 4.80 0.0836 3.04 159.82 370.15 0.0186 0.0157
-Table S17: Effective sample sizes (ESS) per parameter and model for sparse datasets.
-ESS computed on the last 10,000 samples of the chains resulting from running the PF method on the two-variant SEIR model and using a maximum lag size of 250 for the autocorrelation.
-Parameter $\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$
-Model
-Sparse 1-1-1 313.9 536.5 302.4 364.4 1550.2 1641.1
-Sparse 1-1-2 208.8 938.9 211.3 261.2 907.0 757.0
-Sparse 1-1-3 292.1 771.5 287.9 304.3 914.6 1157.6
-Sparse 1-2-1 252.0 1121.2 258.3 308.7 463.0 587.4
-Sparse 1-2-2 326.5 874.5 303.6 380.3 738.2 1216.5
-Sparse 1-2-3 301.1 1234.2 309.9 343.2 760.9 873.2
-Table S18: $\mathbf{\hat{R}}$ diagnostics for the SEIR model with sparse data.
-Parameter Set $\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$
-Sparse 1-1-1 1.064 1.034 1.065 1.052 1.002 1.005
-Sparse 1-1-2 1.154 1.003 1.143 1.099 1.008 1.028
-Sparse 1-1-3 1.100 1.019 1.100 1.085 1.018 1.007
-Sparse 1-2-1 1.112 1.007 1.109 1.076 1.036 1.030
-Sparse 1-2-2 1.067 1.016 1.068 1.055 1.024 1.009
-Sparse 1-2-3 1.027 1.007 1.028 1.016 1.011 1.007
-Table S19: List of model parameters, bounds and priors for the reparametrized two variant SEIR model
-Parameter Lower bound Upper bound Prior
-$r0$ 0.95 4.0 $\mathrm{Uniform}(0.95,4.0)$
-$e0$ 6 30 $\mathrm{Uniform}(6,30)$
-$s0$ 1 100 $\mathrm{Uniform}(1,100)$
-$t_{\mathrm{var}}$ 120 360 $\mathrm{Uniform}(120,360)$
-$I0$ 10 1000 $\mathrm{Uniform}(10,1000)$
-Table S20: Parameters for dense synthetic datasets of the reparametrized SEIR model.The datasets are the same as for the full model.
-Dataset $\mathbf{r0}$ $\mathbf{e0}$ $\mathbf{s0}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$
-Dense 1-1 1.36 22 51 150 500
-Dense 1-2 1.36 22 51 150 500
-Table S21: MAP and RMSE for reparametrized SEIR model and dense datasets.
-Dataset $\mathbf{r0}$ $\mathbf{e0}$ $\mathbf{s0}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$ $\mathbf{RMSE_{infc}}$ $\mathbf{RMSE_{prev}}$
-Dense 1-1 True Par. 1.36 22.00 51.00 150.00 500.00 0.0200 0.0245
-CNF 1.34 21.33 50.27 145.46 489.71 0.0216 0.0236
-PF 1.34 20.57 51.21 156.94 467.30 0.0319 0.0271
-Dense 1-2 True Par. 1.36 22.00 51.00 150.00 500.00 0.0214 0.0237
-CNF 1.35 20.24 48.11 156.46 375.09 0.0184 0.0258
-PF 1.34 19.69 49.17 161.32 382.68 0.0212 0.0236
-Table S22: Effective sample sizes (ESS) per parameter and model.
-ESS computed on the last 10,000 samples of the chains resulting from running the PF method on the reparametrized two-variant SEIR model and using a maximum lag size of 250 for the autocorrelation.
-Parameter $\mathbf{r0}$ $\mathbf{e0}$ $\mathbf{s0}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$
-Dense 1-1 2409.8 24533.8 2370.5 2276.2 2382.2
-Dense 1-2 2675.0 2413.2 2495.5 2537.6 2571.3
-Table S23: $\mathbf{\hat{R}}$ diagnostics for the reparametrized SEIR model with dense data.
-Dataset $\mathbf{r0}$ $\mathbf{e0}$ $\mathbf{s0}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$
-Dense 1-1 1.002 1.001 1.005 1.003 1.002
-Dense 1-2 1.002 1.002 1.002 1.002 1.002
-Table S24: MAP and RMSE for full SEIR model for real dataset.
-Dataset $\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa^{-1}}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I_{0}}$ $\mathbf{RMSE_{infc}}$ $\mathbf{RMSE_{prev}}$
-Real data Pub. Par. 16.70 5.00 0.0800 2.30 184.50 635.00 0.0737 0.0702
-CNF 14.80 4.14 0.0882 3.16 183.04 550.16 0.1251 0.0928
-PF 15.03 4.67 0.0858 3.36 160.35 752.08 0.0856 0.0606
-Table S25: Effective sample sizes (ESS) per parameter and model for the real dataset.
-ESS computed on the last 10,000 samples of the chains resulting from running the PF method on the two-variant SEIR model and using a maximum lag size of 250 for the autocorrelation.
-Parameter $\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$
-Model
-Real data 423.8 1312.0 407.3 959.2 798.6 734.5
-Table S26: $\mathbf{\hat{R}}$ diagnostics for the SEIR model with real data.
-Parameter Set $\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I0}$
-Real 1.034 1.004 1.033 1.007 1.010 1.006
-Table S27: Comparison of computation times: Accumulated CPU-time in minutes for both methods applied to a dense dataset.
-For the PF we ran 4 chains on 4 cores in parallel, the CNF was parallelized over 20 cores.
-Datasets PF - normal prior PF - unif. prior CNF - normal prior CNF - unif. prior
-SIR 1 1163.33 - 106.72 -
-Dense 1-1 2095.98 2104.97 252.96 252.48
-Sparse 1-1-1 1798.1 - 132.29 -
-Table S28: Hyperparameters Summary Network: SequenceNetwork
-Hyperparameter Setting
-Summary Dimensions $2\times\texttt{parameter\_size}+2$
-LSTM Units 64
-Convolutional Layers 2
-Table S29: Invertible Network: InvertibleNetwork
-Coupling parameter Setting
-Coupling Type Spline
-Coupling Layers 8
-Table S30: Simulation Configuration.
-Hyperparameter Setting
-Training Mode Offline
-Simulations 100,000
-Batch Size 32
-Validation Simulations 400
-Epochs 100
-Table S31: Calibration Setup
-Hyperparameter Setting
-Datasets 2,000
-Inference Draws 100 per dataset
-Appendix C Supplementary Figures
-Figure S1: Posterior Geometry 3D Plot of selected pairwise density plots.
-The mountain pass geometry on the left makes inference of parameters more challenging compared to the mountain peak geometry on the right.
-Figure S2: Illustration of a practical identifiability manifold in the two-variant SEIR model.
-Orange: simulations from the fixed parameter set (17, 5, 0.08, 3, 150, 500)(17,\,5,\,0.08,\,3,\,150,\,500).
-Blue: simulations from parameter sets with $\gamma^{-1}$ ranging from $1$ to 20 and other parameters adjusted according to the reparametrization in (A.3) The near-complete overlap of trajectories demonstrates that distinct parameter combinations can produce almost indistinguishable epidemic curves.
-Figure S3: Metrics for tuning of the particle number.
-(a) and (c) visualize the variance of the log-likelihood estimate across different particle numbers for the SIR and SEIR model.
-(b) and (d) depict the ancestor trees of the particles over time for the SIR and the SEIR model.
-Figure S4: Results of the two-variant SEIR model with dense dataset Dense 1-1.
-A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade.
-B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
-Figure S5: Results of the two-variant SEIR model with dense dataset Dense 1-2.
-A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade.
-B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
-Figure S6: Results of the two-variant SEIR model with dense dataset Dense 2-1.
-A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade.
-B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
-Figure S7: Results of the two-variant SEIR model with dense dataset Dense 2-2.
-A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade.
-B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
-Figure S8: Results of the two-variant SEIR model with sparse dataset Sparse 1-1-1.
-A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade.
-B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
-Figure S9: Results of the two-variant SEIR model with sparse dataset Sparse 1-1-2.
-A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade.
-B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
-Figure S10: Results of the two-variant SEIR model with sparse dataset Sparse 1-1-3.
-A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade.
-B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
-Figure S11: Results of the two-variant SEIR model with sparse dataset Sparse 1-2-1.
-A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade.
-B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
-Figure S12: Results of the two-variant SEIR model with sparse dataset Sparse 1-2-2.
-A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade.
-B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
-Figure S13: Results of the two-variant SEIR model with sparse dataset Sparse 1-2-3.
-A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade.
-B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
-Figure S14: Results of the two-variant SEIR model with sparse dataset Real data.
-A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade.
-B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
-Figure S15: Results of the reparametrized two-variant SEIR model with dense dataset Dense 1-1.
-A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade.
-B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
-Figure S16: Results of the reparametrized two-variant SEIR model with dense dataset Dense 1-2.
-A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade.
-B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
-Figure S17: Results of the SIR model with dataset SIR 1.
-A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade.
-B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
-Figure S18: Results of the SIR model with dataset SIR 2.
-A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade.
-B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
-Figure S19: Results of the SIR model with dataset SIR 2.
-A Posterior approximations based on 10,000 samples, true parameters are indicated by bold black lines and joint MAP estimates appear in a darker shade.
-B Model fit with percentiles ($50\%$, $90\%$, $95\%$) based on simulating the 10,000 samples.
-Figure S20: ECDF Calibration plots for the full SEIR model and dense datasets.
-(a) Dense 1-1 (b) Dense 1-2 (c) Dense 1-3 (d) Dense 2-1 (e) Dense 2-2 (f) Dense 2-3
-Figure S21: ECDF Calibration plots for the full SEIR model on sparse and the real datasets.
-(a) Sparse 1-1-1 (b) Sparse 1-1-2 (c) Sparse 1-1-3 (d) Sparse 1-2-1 (e) Sparse 1-2-2 (f) Sparse 1-2-3 (g) Real data
-Figure S22: ECDF Calibration plots for the reparametrized SEIR model and dense datasets.
-(a) Dense 1-1 (b) Dense 1-2
-Figure S23: ECDF Calibration plots for the SIR model.
-(a) SIR 1 (b) SIR 2
-Figure S24: Parameter recovery for the full SEIR model and dense datasets.
-(a) Dense 1-1 (b) Dense 1-2 (c) Dense2-1 (d) Dense 2-2
-Figure S25: Parameter recovery for for the full SEIR model on sparse and the real datasets.
-(a) Sparse 1-1-1 (b) Sparse 1-1-2 (c) Sparse 1-1-3 (d) Sparse 1-2-1 (e) Sparse 1-2-2 (f) Sparse 1-2-3 (g) Real data
-Figure S26: Parameter recovery for the reparametrized SEIR model and dense datasets.
-(a) Dense 1-1 (b) Dense 1-2
-Figure S27: Parameter recovery for for the SIR model.
-(a) SIR 1 (b) SIR 2
+There was no additional external funding received for this study.